A Bibliography of Publications about ISO Standards for Programming Languages

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

01 November 2024 Version 2.45

Title word cross-reference

1 [II94, ISO94d, ISOxxl]. 1.7 [ISO08a]. 10 [ISOxxn]. 1003.4 [GL91]. 10118-3 [ISO04b]. 10303 [ISOxxf, ISOxxa]. 10303-11 [ISO94a]. 10303-12 [ISO97s, ISOxxf]. 10303-23 [ISOxxa]. 10514-1 [ISO96d]. 10514-2 [ISO98d]. 10514-3 [ISO98e]. 10646-1 [ISO93d]. 10743 [ISOxxh]. 10858 [ISOxxi]. 10918-1 [ISO94c]. 10918-2 [ISO95c]. 10918-3 [ISO97g, ISO99d]. 10918-4 [ISO99e]. 10967 [ISO01a, ISO06b, ISO06a]. 10967-1 [ISO94d]. 10967-2 [ISO01a, ISO06b]. 10967-3 [ISO06a]. 11730 [Ame94]. 11730-1994 [Ame94]. 11756 [ISOxxj]. 1178 [Ins91]. 1178-1990 [Ins91]. 12 [ISOxxf].

12088-4 [ISO95d]. 13211-1 [ISO95g]. 13211-2 [ISO00c]. 13249-3 [ISO99g]. 13281-2 [ISO00a]. 13719-1 [ISO98f]. 13719-2 [ISO98g]. 13719-3 [ISO98h]. 13719-4 [ISO98i]. 13751 [ISOxxk]. 13817-1 [ISO96f]. 14496-1 [ISO99h]. 14496-2 [ISO99i]. 14496-3 [ISO99j]. 14496-5 [ISO99k, ISO00e]. 14496-6 [ISO99l]. 14515 [ISOxxl]. 14515-1 [ISOxxl]. 14772-1 [ISO98l]. 1538-1972 [ISO72]. 1539-1 [Ame97, ISO97k, ISO97l, ISO04a, ISO04c, ISO10a]. 1539-2 [ISO94h, ISO00f]. 1539-3 [ISO99p]. 15445 [ISOxxm]. 15889 [ISOxxc]. 19005-1 [ISO05a]. 1976 [ISO79b]. 1978 [ISO78a, ISO78b]. 1979 [ISO79b]. 1981 [ISO81]. 1983 [Eur85]. 1985 [ISO85d, ISO85g, ISO85f, ISO85b]. 1985/Amd [ISO85c]. 1986 [ISO86e].

1986-10-15 [ISO86b]. 1986/Amd [ISO88j]. 1986/Cor [ISO96b, ISO99c]. 1988 [ISO88j, ISO88g, ISO88m, ISO88n]. 1989 [ISO89c, ISO89b]. 1989.2 [ISO21]. 1990 [Ins91, ISO90a, ISO90g, ISO90c, ISO90h]. 1990/AMD [ISO95r, ISO94p, ISO95p]. 1990/Cor [ISO93j, ISO93k, ISO94o, ISO94q]. 1991 [ISO91b, ISO91d]. 1991/Amd [ISO91i, ISO94r, ISO94b, ISO98w, ISO98y]. 1991/Cor [ISO92h, ISO93i, ISO94m, ISO94s, ISO95n, ISO95o]. 1992 [ISO92h, ISO92c, ISO92e, ISO92a]. 1992/Cor [ISO96w, ISO99-41]. 1993 [Int93, ISO93i, ISO93j, ISO93k, ISO93c, ISO93e, ISO93l, ISO93b, ISO93m]. 1993/A1 [ISO95r]. 1993/Amd [ISO93f]. 1994 [Ame94, ISO94a, ISO94r, ISO94m, ISO94o, ISO94q, ISO94s, ISO94f, ISO94g, ISO94h, ISO94l, ISO94i]. 1994-12 [ISO94d]. 1995 [Ame95, ISO95p, ISO95n, ISO95o, ISO95e, ISO95d, ISO95b, ISO95g, ISO95k, ISO95l, ISO95q]. 1995/Amd [ISO07c]. 1995/Cor [ISO01b]. 1996 [ISO96b, ISO96w, ISO96c, ISO96d, ISO96e, ISO96f, ISO96g, ISO96h, ISO96i, ISO96j, ISO96k, ISO96l, ISO96m, ISO96n, ISO96o, ISO96p, ISO96q, ISO96r, ISO96s, ISO96t, ISO96y]. 1997 [Ame97, ISO97e, ISO97k, ISO97f, ISO97i, ISO97h, ISO97j, ISO97o, ISO97l, ISO97b, ISO97c, ISO97s]. 1997/Amd [ISO99d]. 1998 [ISO98v, ISO98d, ISO98e, ISO98f, ISO98g, ISO98h, ISO98i, ISO98j, ISO98k, ISO98l, ISO98n, ISO98p, ISO98r, ISO98u, ISO98q, ISO98s, ISO98t, ISO98m, ISO98z, ISO98-27, ISO98-28, ISO98-29, ISO01c, ISO01d]. 1999 [Ame99, ISO99c, ISO99-41, ISO99a, ISO99f, ISO99-32, ISO99-33, ISO99-34, ISO99-35, ISO99-36, ISO99-37, ISO99-31, ISO99m, ISO99n, ISO99o, ISO99p, ISO99r, ISO99-29, ISO99-42, ISO99-43, ISO07a]. 1999/Cor [ISO01e].

2 [ISO96d, ISO98d, ISO98e, ISO01a, ISO06b, ISO09b, ISO19]. 2.0.1 [ISO05c]. 2000 [Int00, ISO00f, ISO00c]. 2000-type [ISO78b]. 2001 [ISO01e]. 2003 [ISO03c, ISO03b, ISO03a, ISO05b]. 2005 [ISO05a]. 2006 [ISO06c, ISO06d, ISO06f, ISO06g]. 2007 [ISO17b]. 2009 [ISO09a, ISO09d]. 2010 [ISO10b]. 2012 [ISO12a]. 2014 [ISO14b]. 2017 [ISO17a]. 2019 [ISO19]. 2020 [ISO20c]. 2023 [ISO23]. 2024 [ISO24]. 23 [ISOxxa]. 23272 [ISO05d]. 233-2 [ISO93c]. 233-3 [ISO99a]. 23360 [ISO05c]. 2382-1 [ISO93g]. 2382-10 [ISO79a]. 2382-12 [ISO88d]. 2382-13 [ISO96u]. 2382-14 [ISO97m]. 2382-15 [ISO85d, ISO99u]. 2382-16 [ISO96v]. 2382-17 [ISO99v]. 2382-18 [ISO99w]. 2382-19 [ISO89a]. 2382-2 [ISO76]. 2382-20 [ISO90d]. 2382-21 [ISO85e]. 2382-22 [ISO86c]. 2382-23 [ISO94j]. 2382-24 [ISO95h]. 2382-25 [ISO92d]. 2382-26 [ISO93h]. 2382-27 [ISO94k]. 2382-28 [ISO95i]. 2382-29 [ISO99x]. 2382-3 [ISO87c]. 2382-31 [ISO97n]. 2382-32 [ISO99y]. 2382-34 [ISO99z]. 2382-4 [ISO99-27]. 2382-5 [ISO99-28]. 2382-6 [ISO87d]. 2382-7 [ISO00g]. 2382-8 [ISO98o]. 2382-9 [ISO95j]. 24731-1 [ISO07d]. 2709 [ISO81]. 2709-1981 [ISO81].

3 [ISO06a]. 3.0 [ISO20a]. 3166-1 [ISO97d]. 3166-2 [ISO98b]. 3166-3 [ISO99b]. 32000-1 [ISO08a]. 32000-2 [ISO20a]. 3592 [ISOxxd]. 3D [ISO88i, ISO91d].

4 [II94]. 4343 [ISOxxe].

61131-3 [Int93]. 61131-8 [Int00]. 646 [ISO83b]. 690 [ISO87a].

7-Bit [ISO83b, ISO91c]. 7185 [Eur85]. 7185-1982 [ISO82]. 7185-1983 [Eur85]. 7816-7 [ISO99-29]. 7942 [ISO85a]. 7942-1 [ISO94l]. 7942-2 [ISO97o]. 7942-3 [ISO99-30]. 7942-4 [ISO98p].

8 [ISO98l]. 8-bit [ISO98r, ISO98u, ISO98q, ISO98s, ISO98t, ISO99-32, ISO99-33, ISO99-34, ISO99-35, ISO99-36, ISO99-37, ISO99-31, ISO01c, ISO01d, ISO03c]. 80000 [ISO09b]. 80000-2 [ISO19, ISO09b]. 8485 [ISO83a]. 8651-1 [ISO88f]. 8651-2 [ISO88g]. 8651-3 [ISO88h]. 8651-4 [ISO95k]. 8652 [Ame95, II94]. 8652-1995 [Ame95]. 8806-4 [ISO91d]. 8859-1 [ISO98r]. 8859-10 [ISO98q]. 8859-11 [ISO01c]. 8859-13 [ISO98s]. 8859-14 [ISO98t]. 8859-15 [ISO99-31]. 8859-16 [ISO01d]. 8859-2 [ISO99-32]. 8859-3 [ISO99-33]. 8859-4 [ISO98u]. 8859-5 [ISO99-34]. 8859-6 [ISO99-35]. 8859-7 [ISO03c]. 8859-8 [ISO99-36]. 8859-9 [ISO99-37].

9075 [ISOxxn]. 9075-1 [ISO99-38, ISO11b]. 9075-10 [ISO08b, ISOxxn]. 9075-13 [ISO08c]. 9075-14 [ISO11c]. 9075-2 [ISO99-39, ISO11e, ISO11d]. 9075-3 [ISO95m, ISO08d]. 9075-4 [ISO99-40, ISO11f]. 9075-5 [ISOxxg]. 9075-9 [ISO08e]. 95 [ABM^{+}97]. 9541 [SBSG89]. 9541-1 [ISO91f, ISO92h, ISO94m, ISO95n, ISO98w]. 9541-2 [ISO91g, ISO93i, ISO95o]. 9541-3 [ISO94n]. 9573-11 [ISO92i]. 9573-13 [ISO91j]. 9579-2 [ISO98x]. 9579-3 [ISO96x]. 9592-1 [ISO97p]. 9592-2 [ISO97q]. 9592-3 [ISO97r]. 9593-1 [ISO90e, ISO93j, ISO94o, ISO95p]. 9593-3 [ISO90f, ISO93k, ISO94p, ISO94q]. 9593-4 [ISO91i, ISO91h, ISO94r, ISO94s, ISO98y]. 9638-3 [ISO94t]. 9899 [Ame99]. 9899-1999 [Ame99]. 9945-1 [ISO96y].

= [ISO86b, ISO88l, ISO88b, ISO88c].

A-1 [ISO05a]. A1 [ISO95r]. abbreviation [ISO97e]. abstract [ISO96h, ISO96j, ISO96l, ISO96n, ISO98f]. acceptability [ISO88n]. Access [Ano93, ISO96x, ISO98x, ISOxxa, ISO00h]. Ada [ISO93f, ISO95f, ISO99m, ISO99t, ISO00i, ISO90f, ISO94e, Ame95, ISO88h, II94, ISO94f, ISO94t, ISO95e, ISO95l, ISO96z, ISO98h, ISO98j, ISO98k, ISO99o, ISO01b, ISO07c, ISO12b]. address [ISO98w]. addressing [ISO93b]. advances [ISO98w]. ALGOL [ISO72]. allow [ISO99d]. alphabet [ISO98r, ISO98u, ISO98q, ISO98s, ISO98t, ISO99-32, ISO99-33, ISO99-34, ISO99-35, ISO99-36, ISO99-37, ISO99-31, ISO01c, ISO01d, ISO03c]. AMD [ISO95r, ISO85c, ISO88j, ISO91i, ISO93f, ISO94p, ISO94r, ISO94b, ISO95p, ISO98w, ISO98y, ISO99d, ISO07c]. Amendment [ISO95r, ISO07c, ISO85c]. amendments [ISO91i, ISO98y]. American [Sch90]. Analog [ISO89a]. Annotated [Sch90]. ANSI [Ame94, Ame95, Ame97, Ame99, ABM^{+}97, ISO79b, ISO85b, Sch90]. ANSI/ISO/IEC [Ame94, Ame95, Ame97, Ame99]. API [ISO96h, ISO96i, ISO96j, ISO96k, ISO96l, ISO96m, ISO96n, ISO96o, ISO96p, ISO96q, ISO96r, ISO96s, ISO96y, ISO99m, ISO99r]. APL [ISO83a, ISO89b, ISOxxk]. Application [ISO92i, ISO96y, ISO99m, ISO99g, Ass87a, Ass87b, Ass89, Int00, Smi87, ISO95d, ISO96h, ISO96i, ISO96j, ISO96k, ISO96l, ISO96m, ISO96n, ISO96o, ISO96p, ISO96q, ISO96r, ISO96s]. Applications [ISO06e]. APPn [ISO99e]. Approach [DDHR94]. April [IEE94]. Arabic [ISO84, ISO93c, ISO99a, ISO99-35]. Architecture [App91, ISO93d, ATL93, ISO98w, Mol91, FFK92, ISO91f, ISO97a]. Archive [ISO97q, ISO97r, ISO98p]. area [ISO92d]. Arithmetic [ISO76, Kar05, Rei93, ISO94d, ISOxxi, ISO01a, ISO06b, ISO06a, ISO09d, ISO20c, ISO11h]. Artificial [ISO95i, ISO97n, ISO99x, ISO99z]. ASCII [ISO97c]. ASIS [ISO99o]. assessment [ISO99t]. audio [ISO99k, ISO99l, ISO99h, ISO99i, ISO99j, ISO00e, ISO99j]. audio-visual [ISO99k, ISO99l, ISO99h, ISO99i, ISO99j, ISO00e]. Audit [ISO99-30]. Authorities [ISO99e]. Automated [LL93a]. automation [ISO93m, ISO94a, ISO97a, ISO97s, ISOxxf, ISOxxa, ISO00a, ISO94k]. availability [Sch88, ISO97m]. Available [Ass87a, Hil95].

b [Int00]. balisage [ISO86b]. Baltimore [ITC93]. Base [Ame97, ISO96d, ISO97k, ISO04a, ISO04c, ISO05c, ISO10a, ISO96f, ISO97l]. based [ISO96k, ISO96o, NKN91, ISO96i, ISO96m, ISO97f, MW92]. Basic [ISO93d, ISO95i, ISO98j, ISO91b, ISO85f]. be [ISO09b]. Beginning [SP87]. between [ISO85e]. Bibliographic [ISO81, ISO87a]. binding [ISO98g, ISO98h, ISO98i, ISOxxa, ISO96l, ISO96m, ISO96n, ISO96o, ISO96r, ISO96s, ISO99m, ISO93f]. Bindings [ISOxxg, ISO08b, ISO88f, ISO88g, ISO88h, Int88, ISO90e, ISO90f, ISO91d, ISO91h, ISO93l, ISO94t, ISO95d, ISO95k, ISOxxn]. birth [Adl92]. Bit [ISO83b, ISO91c, ISO98r, ISO98u, ISO98q, ISO98s, ISO98t, ISO99-32, ISO99-33, ISO99-34, ISO99-35, ISO99-36, ISO99-37, ISO99-31, ISO01c, ISO01d, ISO03c]. BNF [ISO96t]. Book [ISO92b]. Bounds [ISO07d]. Bounds-checking [ISO07d]. BS [Bri03b]. bureaucratique [ISO88b]. bureautiques [ISO86b]. byte [ISO98r, ISO98u, ISO98q, ISO98s, ISO98t, ISO99-32, ISO99-33, ISO99-34, ISO99-35, ISO99-36, ISO99-37, ISO99-31, ISO01c, ISO01d, ISO03c].

C [ISO95r, ISO07d, ISO09c, ISO11a, ISO14a, ISO17c, Ame99, Bri03a, Bri03b, ISO90g, ISO91d, ISO91h, ISO95d, ISO95k, ISO96l, ISO96m, ISO96n, ISO96o, ISO96r, ISO96s, ISO96y, ISO98g, ISO98m, ISO99-42, ISOxxa, ISO03a, ISO09d, ISO10b, ISO11g, ISO18, ISO24, Sch90]. C# [ISO03b, ISO06c]. c1987 [Ass87a]. Calculators [ISO86c]. Call [ISO95m, ISO08d]. Call-Level [ISO95m, ISO08d]. Calling [ISO96g]. CAPS [AW93, Eng93]. Card [ISO99-29]. cards [ISO99-29]. CCITT [ISO98v]. CD [ISO88l, ISO21]. CD-ROM [ISO88l]. Celtic [ISO98t]. Center [ITC93]. Central [ISO92i]. CGI [ISO94t]. Character [BB93, Gay95, ISO83b, ISO93d, ISO94i, SP87, ISO91c, ISO92e, ISO94b, ISO94h, ISO98r, ISO98u, ISO98q, ISO98s, ISO98t, ISO99-32, ISO99-33, ISO99-34, ISO99-35, ISO99-36, ISO99-37, ISO99-31, ISO00f, ISO01c, ISO01d, ISO03c, Kim92]. characters [ISO84, ISO93c, ISO99a, ISO84]. charts [ISO97i]. checking [ISO07d]. CHILL [ISO98v]. circuit [ISO99-29]. clarification [ISO85c]. clauses [ISO90h]. clear [ISO97r]. clear-text [ISO97r]. CLI [ISO95m, ISO06d, ISO06f, ISO06g, ISO08d]. CMM [Pau95]. COBOL [ISO85c, ISO85b, ISO02, ISO14b, ISO21, ISO23]. Code [BB93, ISO86d, ISO95q, ISO99b, ISO94i, ISO98b]. Coded [ISO83b, ISO93d, ISO91c, ISO92e, ISO98r, ISO98u, ISO98q, ISO98s, ISO98t, ISO99-32, ISO99-33, ISO99-34, ISO99-35, ISO99-36, ISO99-37, ISO99-31, ISO01c, ISO01d, ISO03c]. Codes [ISO95a, ISO97d, ISO98b, ISO99b, Kim92, ISO97d]. Coding [ISO99k, ISO99l, ISO99h, ISO99i, ISO99j, ISO00e, ISO94c, ISO95c, ISO97g, ISO99e]. Colloquium [Ano91, Ano93]. Colorado [IEE94]. colour [ISO99e]. commands [ISO78b, ISO99-29, ISOxxe]. Common [ISO98f, ISO98g, ISO98h, ISO98i, ISO05b, ISO05d, ISO06g, ISO06d, ISO06f, ISO06g]. Communication [GS87, ISO99q, Wil94, ISO00d, ISO95j]. compacts [ISO88l]. Compares [Pau95]. compilation [ISO99p]. Complete [ABM^{+}97]. Complex [ISO06a, ISO98j, ISO98k, ISO06b, ISO06b]. Compliance [ISO95c, ISOxxl]. compression [ISO94c, ISO95c, ISO97g, ISO99d, ISO99e]. Computer [Ano92, ISO82, ISO88f, ISO88h, ISO88i, ISO90e, ISO90f, ISO91h, ISO94t, ISO95h, ISO96u, ISO97p, ISO97q, ISO97r, ISO99-30, ISO00g, Eur85, ISO85e, Int88, ISO88g, ISO91d, ISO94l, ISO95d, ISO95k, ISO97o, ISO98l, ISO98p]. Computer-integrated [ISO95h]. computing [ISO89a]. concepts [ISO95i]. Conditional [ISO99p]. Conference [ITC93, IEE94]. Conformance [ISO96x, ISO00d, ISO96j, ISO96k, ISO96n, ISO96o, ISO96q, ISO96s]. conforming [Ass87a, Ass87b, Ass89]. Conformity [ISO99t, ISO90h]. constructs [ISO89c]. contacts [ISO99-29]. Content [ISO87a]. continuous [ISO94c, ISO95c, ISO97g, ISO99e]. continuous-tone [ISO94c, ISO95c, ISO97g, ISO99e]. Control [ISO92e, SP87, ISO78a, ISO78b, ISO85f, ISOxxd, ISOxxe]. controllers [Int93, Int00]. Convenience [Ano92]. Convention [ITC93]. conventions [ISO89c]. Cor [ISO92h, ISO93i, ISO93j, ISO93k, ISO94m, ISO94o, ISO94q, ISO94s, ISO95n, ISO95o, ISO96b, ISO96w, ISO99c, ISO99-41, ISO01b, ISO01e]. Core [ISO05c, ISO95g]. Correction [ISO85c]. Corrigendum [Bri03a, Bri03b, ISO01b, ISO07a]. countries [ISO97d, ISO98b, ISO99b, ISO99b]. Country [ISO97d, ISO98b]. cultural [ISO99s]. currencies [ISO95a]. current [BB91]. Cyrillic [ISO99-34].

d’application [II90]. Data [Ano91, ISO76, ISO79a, ISO85e, ISO88e, ISO95j, ISO00b, ISO08e, ISO94a, ISO96h, ISO96j, ISO96l, ISO96n, ISO97b, ISO97c, ISO97s, ISO98a, ISO98-29, ISO99w, ISOxxf, ISOxxa, ISOxxc, ISO85d, ISOxxc, ISO87d, ISO99-27, ISO99-28]. Database [ISO92g, ISO95m, ISO96x, ISO98x, ISO99-38, ISO99-39, ISO99g, ISO99-40, ISOxxg, ISO08d, ISO08e, ISO08b, ISO08c, ISO11b, ISO11e, ISO11f, ISO11d, ISO11c, ISO00h, ISOxxn]. Databases [ISO99v]. datatypes [ISO96e]. dates [ISO88e, ISO00b]. d’échange [ISO88b]. decimal [ISO09d]. declarations [ISO98j]. Dedicated [ISO04b]. Definition [DDHR94, ISO98i, ISO99n, ISO11d]. Delivery [ISO99l]. d’entités [Int91]. derived [ISO06f]. describing [Rol91]. Description [ISO95e, ISO95b, ISO95f, ISOxxh, ISO94l, ISOxxc, ISOxxm, Kom91, ISO94a, ISO97s, ISOxxf, ISO88i, ISO97p]. Designing [ITC93]. destinés [ISO88l]. developing [Hil95]. Development [ISO96f, SP87, ISO88n, ISO90d]. developments [BB91]. devices [ISO94t]. diagnostics [ITC93]. Diagram [ISO97i]. dialogues [ISO94t]. Dictionary [ISO93e]. Digest [Ano91, Ano93]. Digital [ISO94c, ISO95c, ISO97g, ISO99e, ISO12a, ISO98a]. Dimensions [ISO88i, ISO91d]. d’information [ISO88l]. direction [Kim92]. directory [ISO96q, ISO96s, ISO96p, ISO96r]. DIS [II94, ISOxxh, ISOxxi, ISOxxj, ISOxxk, ISOxxl, ISOxxm, ISOxxn, ISO05c, Rei93, ISOxxa, ISOxxb, ISOxxd, ISOxxe, ISOxxc]. Disability [Ano93]. disques [ISO88l]. Distributed [GS87, ISO99n, ISO99w]. DMIF [ISO99l]. Do [Tri95]. Document [App91, AW93, Can89, FFK92, ISO86a, ISO88k, ISO88b, ISO00d, ISO05a, ISO06e, ISO08a, ISO20a, Sch88, SP87, ATL93, BB91, Kan93, Kom91, Mol91, NKN91, Ada92, ISO83a, ISO96c, ISOxxm]. Documentation [ISO84, ISO86d, ISO87a, ISO92b, ISO93c, ISO93a, ISO96a, ISO97e, ISO98c, ISO99a, ISO07b, ISO09a, ISO12a, ISO17a, ISO20b]. documents [Mol91, Rol91]. domaines [Int91]. DP [ISO83a]. Draft [ISO83a, ISO05c, II94, ISO04a, Rei93]. DSSSL [Ada92, ISO96c]. DTR [ISOxxf].

Dublin [Ass87a]. d’utilisation [ISO88c, Int91]. DXL [ISO97i].

Early [GL91]. EAST [ISOxxc]. EC [Fre90]. échange [ISO88l]. ECMAScript [ISO98n]. Ed [Int93, Int00]. edition [ISO86b]. Electronic [Ano91, ISO99y, ISO05a, SP87, Ass87a, ISO96i, ISO96k, ISO96m, ISO96o, Mol91, Ass87a]. elementary [ISO94e, ISO98k, ISO06b, ISO06a, ISO01a]. elements [ISO78b, ISO88e, ISO00b, ISO99s]. encoded [ISO97c]. Encoding [SP87, ISO97r, ISO98l]. Endorsement [ISO79b, ISO85b]. Engineering [IEE94]. English [ISO95r, ISO97c]. Enhanced [ISO98-29]. enhancement [ISO94b, ISO00h]. enhancements [ISO98w]. Ensembles [Int91]. entity [ISO91j]. entre [II90]. Environment [ISO97a, ISO00a, ISO93m, ISO98f, ISO98g, ISO98h, ISO98i]. environments [ISO93l, II94, ISO96e, ISO96f, ISO97h, ISO99-43, ISOxxk, ISOxxl, ISO07d, ISO09c, ISO09d, ISO10b, ISO17b, ISO21, ISO23]. Equipment [ISO87c, ISO88d]. Evolving [LL93b]. exception [ISO98-28]. eXchange [ISO97i, ISO94a, ISO97s, ISO98a, ISOxxf, ISOxxa, ISO96a]. experience [GL91]. expert [ISO95i]. EXPRESS [ISO94a, ISO97s, ISOxxf]. EXPRESS-I [ISO97s, ISOxxf]. Extended [ISO91a, ISOxxk, ISO90a, ISO96t]. Extension [ISO09d, ISO94i]. Extensions [ISO96z, ISO99m, ISO07d, ISO09c, ISO10b, ISO97g]. External [ISO08e].

facilités [ISO88c, Int91]. Facilities [ISO88b, Can89, Can90, ISO87b, ISO88k, ISO88m, ISO88a, ISO88c, ISO91e, ISO91j, Int91, ISO92i, ISO98-29, ISO79a]. fichier [ISO88l]. file [ISO88l, ISO97q, ISO98a, ISO05a, ISO06f, ISO09a, ISO17a, ISOxxd, ISO97r]. files [ISO85g, MW92]. FIMS [ISO94g]. First [IEE94, ISO86b, BHK91]. fixe [ISO88l]. Floating [ISO11h, ISO20c, ISO94d, ISOxxi, ISO06b, ISO06a, ISO09d, ISO98-28]. Floating-Point [ISO11h, ISO20c, ISOxxi, ISO09d, ISO98-28]. flow [ISO85g]. Font [ISO91f, ISO91g, ISO94n, SP87, SBSG89, ISO98w, Kom93]. fonts [Kom91]. foreign [ISO95r]. Form [ISO87a, Ame94, ISO94g]. Formal [DDHR94, RAMP94, ATL93, Sev95]. Format [App91, Can89, ISO81, ISO86a, ISO88k, ISO88b, ISO96a, ISO06e, ISO98a, ISOxxd, ISO05a, ISO08a, ISO09a, ISO17a, ISO20a, MW92, ISO91g, ISO97q]. formats [ISO88e, ISO00b]. formerly [ISO99b]. FORTH [ISO97j]. Fortran [ISO04a, ISO04c, ISO10a, ISO88f, ISO90e, ISO94h, Ame97, ABM^{+}97, II91, ISO97k, ISO97l, ISO98-28, ISO98-29, ISO99p, ISO00f]. Foundation [ISO99-39, ISO11e]. Framework [ISO99l, ISO99-38, ISO11b, ISO98-27]. front [Kom93]. Full [ISO91b]. Functional [ISO97a, ISO97p, ISO88i, ISO94l, ISO98l]. functions [ISO92e, ISO94e, ISO94f, ISO98k, ISO01a, ISO04b, ISO06b, ISO06a, ISO10b, ISO09c]. Fundamental [ISO93g]. funds [ISO95a]. future [BB91, Kim92, Pri93].

general [ISO92f, ISO95g]. généralisé [ISO86b]. Generalized [Ass87b, Ass89, Can89, Hei94, ISO86e, ISO86b, ISO00d, Ass87a, Smi87]. Generic [ISO94e, ISO94f, ISO98j, ISO98k]. generics [ISO06g, ISO98d]. GKS [ISO85a, ISO88f, ISO88h, ISO88i, ISO99-30, ISO88g, Int88, ISO91d, ISO94l, ISO95k, ISO97o, ISO98p]. GKS-3D [ISO88i, ISO91d]. Glyph [ISO94n]. golden [Mee95]. Graphic [ISO98a, ISO98r, ISO98u, ISO98q, ISO98s, ISO98t, ISO99-32, ISO99-33, ISO99-34, ISO99-35, ISO99-36, ISO99-37, ISO99-31, ISO01c, ISO01d, ISO03c]. Graphical [ISO88f, ISO88h, ISO88i, ISO99-30, ISOxxb, Int88, ISO94t, ISO85a, ISO88g, ISO91d, ISO94l, ISO95k, ISO97o, ISO98p]. graphics [ISO88f, ISO88g, ISO88h, ISO88i, Int88, ISO90e, ISO90f, ISO91d, ISO91h, ISO94l, ISO94t, ISO95d, ISO95k, ISO97o, ISO97p, ISO97q, ISO97r, ISO98l, ISO98p, ISO99-30, ISO96u]. graphique [II90]. Greek [ISO03c]. Group [Ass87a]. groups [ISO85g]. GUI [ISOxxb]. GUI-R [ISOxxb]. Guide [ISO00i]. Guidelines [Int00, ISO99-43, ISO88n, ISO90h, ISO93l, ISO98z, ISO94c].

Handbook [ABM^{+}97]. handling [ISO87d, ISO98-28]. Hash [ISO04b]. Hash-functions [ISO04b]. header [ISO99d]. Hebrew [ISO99-36]. Heights [Wil94]. Hierarchical [ISO90e, ISO90f, ISO91h, ISO97p, ISO97q, ISO97r]. high [ISO98v, ISO00i]. Host [ISOxxg]. HTML [ISOxxm]. Hyper [ISOxxm]. hypermedia [Gol91, Kom94, NKN91, Pri93, ISO97f]. hypermedia/time [NKN91, ISO97f]. hypermedia/time-based [NKN91, ISO97f]. HyTime [NKN91, CBDH94, Gol91, ISO97f].

ICR [ISO95q]. Identification [ISO99-29]. identifier [ISO12a]. identifiers [ISO88a, ISO91e]. IDL [ISO98i]. IEC [ISO95r, ISO04a, ISO07a, ISO11h, Ame94, Ame95, Ame97, Ame99, Int93, Int00, ISO88m, ISO89c, ISO90e, ISO90f, ISO90d, ISO90g, ISO90h, ISO91i, ISO91a, ISO91b, ISO91d, ISO91h, ISO91c, ISO91e, ISO91f, ISO91g, ISO91j, ISO92h, ISO92c, ISO92f, ISO92d, ISO92e, ISO92g, ISO92i, ISO93f, ISO93i, ISO93j, ISO93k, ISO93d, ISO93e, ISO93g, ISO93h, ISO93l, ISO94p, ISO94r, ISO94b, ISO94m, ISO94o, ISO94q, ISO94s, II94, ISO94d, ISO94e, ISO94f, ISO94g, ISO94h, ISO94l, ISO94t, ISO94j, ISO94k, ISO94n, ISO94c, ISO94i, ISO95p, ISO95n, ISO95o, ISO95e, ISO95d, ISO95i, ISO95b, ISO95g, ISO95k, ISO95l, ISO95m, ISO95j, ISO95h, ISO95f, ISO95c, ISO96w, ISO96x, ISO96c, ISO96d, ISO96e, ISO96f, ISO96g, ISO96h, ISO96i, ISO96j, ISO96k, ISO96l, ISO96m, ISO96n, ISO96o, ISO96p, ISO96q, ISO96r, ISO96s, ISO96t, ISO96y, ISO96u, ISO96v, ISO96z, ISO97k, ISO97f, ISO97i, ISO97h, ISO97j, ISO97o, ISO97l, ISO97m, ISO97n, ISO97p, ISO97q]. IEC [ISO97r, ISO97g, ISO98w, ISO98y, ISO98v, ISO98o, ISO98x, ISO98d, ISO98e, ISO98f, ISO98g, ISO98h, ISO98i, ISO98j, ISO98k, ISO98l, ISO98n, ISO98p, ISO98r, ISO98u, ISO98q, ISO98s, ISO98t, ISO98m, ISO98z, ISO98-27, ISO98-28, ISO98-29, ISO99d, ISO99-41, ISO99f, ISO99k, ISO99l, ISO99s, ISO99-32, ISO99-33, ISO99-34, ISO99-35, ISO99-36, ISO99-37, ISO99-31, ISO99m, ISO99n, ISO99o, ISO99p, ISO99q, ISO99w, ISO99r, ISO99x, ISO99y, ISO99-29, ISO99u, ISO99z, ISO99-38, ISO99-39, ISO99-27, ISO99-28, ISO99v, ISO99t, ISO99-30, ISO99g, ISO99-40, ISO99e, ISO99h, ISO99i, ISO99j, ISO99-42, ISO99-43, ISOxxh, ISOxxi, ISOxxj, ISOxxk, ISOxxl, ISOxxm, ISOxxn, ISOxxg, ISO00g, ISO00d, ISO00h, ISO00e, ISO00f, ISO00c, ISO00i, ISO01b, ISO01e, ISO01c, ISO01d, ISO01a, ISO02, ISO03c, ISO03b, ISO03a, ISO04c, ISO04b, ISO05c, ISO05b, ISO05d, ISO06e, ISO06b, ISO06c, ISO06d, ISO06a, ISO06f, ISO06g, ISO07c, ISO07d, ISO08d, ISO08e, ISO08b, ISO08c, ISO09c, ISO09d]. IEC [ISO10b, ISO10a, ISO11g, ISO11b, ISO11e, ISO11f, ISO11d, ISO11c, ISO11a, ISO12b, ISO14b, ISO14a, ISO17c, ISO17b, ISO18, ISO20c, ISO21, ISO23, ISO24, Kar05, Kim92, SBSG89]. IEE [Ano91, Ano93]. IEEE [Ins91, ISO11h]. image [ISO94l, ISO95d, ISO97o, ISO97p, ISO97q, ISO97r, ISO98a, ISO98l, ISO98p, ISO99-30, ISO95d]. images [ISO94c, ISO95c, ISO97g, ISO99e]. imaging [BHK91]. Implementation [ISO96x, Int00, ISOxxa]. Implications [Smi86]. including [ISO98j]. incorporating [Bri03a, Bri03b]. Incorporation [ISO91i, ISO94p, ISO98y]. Independent [Kar05, Rei93, ISO94d, ISO96h, ISO96i, ISO96j, ISO96k, ISO96p, ISO96q, ISO01a, ISO06b, ISO06a, Mee95, ISO96g, ISO99-43, ISO96e, ISOxxi]. Industrial [ISO97a, ISO00a, ISO75, ISO92a, ISO95q, ISOxxb, ISO94a, ISO97s, ISOxxf, ISOxxa]. infographie [II90]. Information [Ame94, Ame95, Ame97, Ano93, App91, Can89, Can90, ISO81, ISO83b, ISO86c, ISO86b, ISO86a, ISO87c, ISO87d, ISO87b, ISO88e, ISO88f, ISO88h, ISO88d, ISO88l, ISO88k, ISO88i, ISO88b, ISO88a, ISO88c, ISO89a, ISO90e, ISO90f, ISO90d, ISO90b, ISO91a, ISO91h, ISO91c, ISO91e, ISO91f, ISO91g, ISO91j, Int91, ISO92b, ISO92f, ISO92d, ISO92e, ISO92g, ISO92i, ISO93a, ISO93d, ISO93g, ISO93h, ISO94e, ISO94t, ISO94j, ISO94k, ISO94n, ISO94c, ISO94i, ISO95i, ISO95m, ISO95j, ISO95h, ISO95f, ISO95c, ISO96a, ISO96x, ISO96u, ISO96v, ISO96z, ISO97e, ISO97k, ISO97m, ISO97n, ISO97p, ISO97q, ISO97r, ISO97g, ISO98c, ISO98o, ISO98x, ISO98r, ISO98u, ISO98q, ISO98s, ISO98t, ISO99k, ISO99l, ISO99s, ISO99-32, ISO99-33, ISO99-34, ISO99-35, ISO99-36, ISO99-37, ISO99-31, ISO99q, ISO99w, ISO99x, ISO99y, ISO99u, ISO99z, ISO99-38, ISO99-39, ISO99-27, ISO99-28, ISO99v, ISO99t, ISO99-30, ISO99g, ISO99-40, ISO99e, ISO99h, ISO99i, ISO99j, ISO99-43]. Information [ISOxxg, ISO00b, ISO00g, ISO00d, ISO00h, ISO00e, ISO00f, ISO00c, ISO00i, ISO01b, ISO01c, ISO01d, ISO02, ISO03c, ISO03b, ISO04a, ISO04c, ISO04b, ISO05b, ISO05d, ISO06c, ISO06d, ISO06a, ISO06f, ISO06g, ISO07c, ISO07b, ISO07d, ISO08d, ISO08e, ISO08b, ISO08c, ISO09a, ISO09c, ISO09d, ISO10b, ISO10a, ISO11h, ISO11g, ISO11b, ISO11e, ISO11f, ISO11d, ISO11c, ISO11a, ISO12a, ISO12b, ISO14b, ISO14a, ISO17a, ISO17c, ISO17b, ISO18, ISO20b, ISO20c, ISO21, ISO23, ISO24, II90, II91, II94, ISO97b, ISO97c, ISOxxc, Kom93, Kom94, SBSG89, ISO85g, ISO86e, ISO88g, ISO88m, ISO89c, ISO90c, ISO91b, ISO91d, ISO92c, ISO93c, ISO93e, ISO93l, ISO94d, ISO94f, ISO94g, ISO94h, ISO94l, ISO95e, ISO95d, ISO95b, ISO95g, ISO95k, ISO95l, ISO96c, ISO96d, ISO96e, ISO96f, ISO96g, ISO96h, ISO96i, ISO96j, ISO96k, ISO96l, ISO96m, ISO96n, ISO96o, ISO96p, ISO96q, ISO96r, ISO96s, ISO96t, ISO96y, ISO97f, ISO97i, ISO97h, ISO97j]. Information [ISO97o, ISO97l, ISO98d, ISO98e, ISO98f, ISO98g, ISO98h, ISO98i, ISO98j, ISO98k, ISO98l, ISO98n, ISO98p, ISO98z, ISO98-27, ISO98-28, ISO98-29, ISO99a, ISO99f, ISO99m, ISO99n, ISO99o, ISO99p, ISO99r, ISOxxh, ISOxxi, ISOxxj, ISOxxk, ISOxxl, ISOxxm, ISOxxn, ISO01a, ISO06b]. Infrastructure [ISO05b, ISO05d, ISO06d, ISO06f, ISO06g]. Initial [DDHR94]. input [ISO85f, ISO94b]. integer [ISO06b, ISO06a, ISO94d]. integrated [ISO95h, ISO99-29]. Integration [ISO99l, ISO94a, ISO97s, ISOxxf, ISOxxa, ISO00a]. integrity [ISO95r, ISO00i]. intelligence [ISO95i, ISO97n, ISO99x, ISO99z]. Interactive [ISO90e, ISO90f, ISO91h, ISO97p, ISO97q, ISO97r]. Interchange [Ano91, App91, Can89, ISO81, ISO83b, ISO86a, ISO88k, ISO88b, SP87, Gol91, ISO88e, ISO88l, ISO91f, ISO91g, ISO94n, ISO95d, ISO00b, Kom93, Kom94, MW92, Mol91, Pri93, SBSG89, ISO91c]. interchangeability [NW92]. Interconnect [ISO99q]. Interconnection [ISO96x, ISO96l, ISO98x, ISO96h, ISO96j, ISO96n, ISO93h]. Interest [Ass87a]. Interface [Ame94, ISO90b, ISO93e, ISO94g, ISO95m, ISO98i, ISO99o, ISO08d, ISO95d, ISOxxa, ISO96h, ISO96i, ISO96j, ISO96k, ISO96l, ISO96m, ISO96n, ISO96o, ISO96p, ISO96q, ISO96r, ISO96s, ISO96y, ISO99m, ISO99n, ISOxxl]. Interfaces [ISO85e, ISO99m, II90, ISO93l, II94, ISO96e, ISO96f, ISO96l, ISO96m, ISO96n, ISO96o, ISO96r, ISO96s, ISO97h, ISO99-43, ISOxxb, ISOxxk, ISOxxl, ISO07d, ISO09c, ISO09d, ISO10b, ISO17b, ISO21, ISO23, ISO00a, ISO07d]. Interfacing [ISO94t]. Interindustry [ISO99-29]. Intermediate [ISO95q]. International [Ass87b, Ass89, Dun90, Dzi96, ITC93, IEE94, ISO86d, II91, ISO92b, ISO92i, ISO93a, ISO98c, ISO04a, ISO05c, ISO07b, ISO20b, Kom91, Kom93, Kom94, Rei93, Ass87a, Ada92, BHK91, Hil95, Kim92, Pri93, ISO07a]. internationalization [ISO98-27]. Introduction [Hei94, Pri93]. IPCC [Wil94]. IR [II94]. IR-MA-1363-4 [II94]. IRDS [ISO93e]. ISBN [ISO92b]. ISLISP [ISO97h, ISO17b]. ISMN [ISO93a]. ISO [Ame94, Ame95, Ame97, Ame99, ISO72, ISO83a, ISO90a, ISO91c, ISO92i, ISO95r, ISO04a, ISO07a, ISO20b, Tri95, Ass87a, Ass87b, Ass89, ABM^{+}97, ATL93, Bri03b, Eur85, II94, Kar05, Kim92, Pau95, SBSG89]. ISO/ANSI [ABM^{+}97]. ISO/DIS [ISOxxa, ISOxxb, ISOxxd, ISOxxe, ISOxxc]. ISO/DP [ISO83a]. ISO/DTR [ISOxxf]. ISO/IEC [ISO95r, ISO04a, ISO07a, ISO88m, ISO89c, ISO90e, ISO90f, ISO90d, ISO90g, ISO90h, ISO91i, ISO91a, ISO91b, ISO91d, ISO91h, ISO91c, ISO91e, ISO91f, ISO91g, ISO91j, ISO92h, ISO92c, ISO92f, ISO92d, ISO92e, ISO92g, ISO92i, ISO93f, ISO93i, ISO93j, ISO93k, ISO93d, ISO93e, ISO93g, ISO93h, ISO93l, ISO94p, ISO94r, ISO94b, ISO94m, ISO94o, ISO94q, ISO94s, II94, ISO94d, ISO94e, ISO94f, ISO94g, ISO94h, ISO94l, ISO94t, ISO94j, ISO94k, ISO94n, ISO94c, ISO94i, ISO95p, ISO95n, ISO95o, ISO95e, ISO95d, ISO95i, ISO95b, ISO95g, ISO95k, ISO95l, ISO95m, ISO95j, ISO95h, ISO95f, ISO95c, ISO96w, ISO96x, ISO96c, ISO96d, ISO96e, ISO96f, ISO96g, ISO96h, ISO96i, ISO96j, ISO96k, ISO96l, ISO96m, ISO96n, ISO96o, ISO96p, ISO96q, ISO96r, ISO96s, ISO96t, ISO96y, ISO96u, ISO96v, ISO96z, ISO97k, ISO97f, ISO97i, ISO97h, ISO97j, ISO97o, ISO97l, ISO97m, ISO97n, ISO97p, ISO97q, ISO97r, ISO97g, ISO98w, ISO98y, ISO98v, ISO98o, ISO98x]. ISO/IEC [ISO98d, ISO98e, ISO98f, ISO98g, ISO98h, ISO98i, ISO98j, ISO98k, ISO98l, ISO98n, ISO98p, ISO98r, ISO98u, ISO98q, ISO98s, ISO98t, ISO98m, ISO98z, ISO98-27, ISO98-28, ISO98-29, ISO99d, ISO99-41, ISO99f, ISO99k, ISO99l, ISO99s, ISO99-32, ISO99-33, ISO99-34, ISO99-35, ISO99-36, ISO99-37, ISO99-31, ISO99m, ISO99n, ISO99o, ISO99p, ISO99q, ISO99w, ISO99r, ISO99x, ISO99y, ISO99-29, ISO99u, ISO99z, ISO99-38, ISO99-39, ISO99-27, ISO99-28, ISO99v, ISO99t, ISO99-30, ISO99g, ISO99-40, ISO99e, ISO99h, ISO99i, ISO99j, ISO99-42, ISO99-43, ISOxxh, ISOxxi, ISOxxj, ISOxxk, ISOxxl, ISOxxm, ISOxxn, ISOxxg, ISO00g, ISO00d, ISO00h, ISO00e, ISO00f, ISO00c, ISO00i, ISO01b, ISO01e, ISO01c, ISO01d, ISO01a, ISO03c, ISO03b, ISO03a, ISO04c, ISO05c, ISO05b, ISO05d, ISO06e, ISO06b, ISO06c, ISO06d, ISO06a, ISO06f, ISO06g, ISO07c, ISO07d, ISO08d, ISO08e, ISO08b, ISO08c, ISO09c, ISO09d, ISO10b, ISO10a, ISO11g, ISO11b, ISO11e, ISO11f, ISO11d, ISO11c, ISO11a]. ISO/IEC [ISO12b, ISO14b, ISO14a, ISO17c, ISO17b, ISO18, ISO20c, ISO21, ISO23, ISO24, Kar05, Kim92, SBSG89, ISO02, ISO04b]. ISO/IEC/IEEE [ISO11h]. ISO/R [ISO72]. ISO/TR [ISO88n, ISO93m, ISO95q, ISO97s]. ISRC [ISO86d]. ISSN [ISO98c, ISO07b, ISO20b]. IV [ISO06f].

Java [ISO08c]. join [ITC93]. JPEG [ISO99e]. JRT [ISO08c]. June [II94].

Kernel [ISO85a, ISO88f, ISO88g, ISO88h, ISO88i, ISO91d, ISO94l, ISO95k, ISO97o, ISO98p, ISO99-30, Int88]. Kommunikation [GS87]. langage [ISO86b, II90]. Language [Ass87b, Ass89, Can89, ISO82, ISO86b, ISO94t, ISO95f, ISO96d, ISO96f, ISO96h, ISO96i, ISO96j, ISO96k, ISO96p, ISO96q, ISO96y, ISO98i, ISO99m, ISOxxn, ISOxxg, ISO00d, ISO03b, ISO05b, ISO05d, ISO06d, ISO06a, ISO06f, ISO06g, ISO08b, ISO08c, Kar05, Ass87a, Ame97, Eur85, Ins91, ISO72, ISO83a, ISO85f, ISO88f, ISO88g, ISO88h, Int88, ISO88n, ISO90e, ISO90f, ISO90h, ISO91d, ISO91h, ISO93f, ISO93c, ISO93l, ISO93m, ISO94a, II94, ISO95d, ISO95k, ISO96l, ISO96m, ISO96n, ISO96o, ISO96r, ISO96s, ISO97k, ISO97h, ISO97l, ISO97b, ISO97s, ISO98v, ISO98g, ISO98h, ISO98n, ISO98z, ISO99a, ISO99t, ISOxxf, ISOxxa, ISOxxd, ISOxxc, ISOxxk, ISO00i, ISO09d, ISO17b, ISO21, ISO23, Mee95, NKN91, Smi87, Ada92, Hei94, ISO86e, ISO94d, ISO95e, ISO95b, ISO96c, ISO96e, ISO96g, ISO97f, ISO97i, ISO98l, ISO99n, ISO99-29, ISO99-43, ISOxxh, ISOxxm, ISO01a, ISO04a, ISO06b, Rei93, ISO04c, ISO10a]. language-independent [Mee95, ISO96g, ISO99-43, ISO96e]. Languages [Ame94, Ame95, Ame99, ISO99-42, ISO01b, ISO07c, Sch90, Ame97, Int93, Int00, ISO79b, ISO85d, ISO85b, ISO89b, ISO90g, ISO90c, ISO91a, ISO91b, II91, ISO92c, ISO92f, ISO92g, ISO93l, II94, ISO94e, ISO94f, ISO94g, ISO94h, ISO95e, ISO95b, ISO95g, ISO95l, ISO95m, ISO95f, ISO95r, ISO96c, ISO96d, ISO96e, ISO96f, ISO97k, ISO97h, ISO97j, ISO97l, ISO98d, ISO98e, ISO98j, ISO98k, ISO98m, ISO98-28, ISO98-29, ISO99f, ISO99o, ISO99p, ISO99r, ISO99-38, ISO99-39, ISO99t, ISO99g, ISO99-40, ISO99-43, ISOxxj, ISOxxk, ISOxxl, ISOxxm, ISOxxn, ISOxxg, ISO00f, ISO00c, ISO00i, ISO02, ISO03a, ISO04a, ISO04c, ISO06c, ISO07d, ISO08d, ISO08e, ISO08b, ISO08c, ISO09c, ISO09d, ISO10b, ISO10a, ISO11g, ISO11b, ISO11e, ISO11f, ISO11d, ISO11c, ISO11a, ISO12b, ISO14b, ISO14a, ISO17c, ISO17b, ISO18, ISO21, ISO23, ISO24, Kom91, ISO99u]. Latin [ISO84, ISO93c, ISO98r, ISO98u, ISO98q, ISO98s, ISO98t, ISO99a, ISO99-32, ISO99-33, ISO99-34, ISO99-35, ISO99-36, ISO99-37, ISO99-31, ISO01c, ISO01d, ISO03c]. Latin/Arabic [ISO99-35]. Latin/Cyrillic [ISO99-34]. Latin/Greek [ISO03c]. Latin/Hebrew [ISO99-36]. Latin/Thai [ISO01c]. learning [ISO97n]. length [ISO75, ISO94h, ISO00f]. Level [ISO95m, ISO08d, ISO98v]. LIA [Kar05, Rei93]. Libraries [ISO05d, II94]. Library [ISO09c, ISO07d, ISO10b]. Linux [ISO05c]. LIPC [ISO96g]. LISS [ISO99-43]. Local [ISO92d]. logic [ISO76]. logical [Rol91, ISO78a]. long [ISO05a]. long-term [ISO05a].

M [ISO99f, ISO99r, ISOxxj]. MA [II94]. Machine [Ano92, ISO97n]. machines [ISO78a, ISO78b, ISO85f, ISOxxd, ISOxxe]. Magnetic [ISO81]. Mail [ISO99y]. maintainability [ISO97m]. major [ISO78a]. Management [Ame94, ISO94g, ISO08e, ISO05a, ISO08a, ISO20a]. Manipulating [ISOxxb, ISO92a, ISO95q]. manipulation [ISO96h, ISO96j, ISO96l, ISO96n]. manual [ISO94a, ISO97s, ISOxxf]. Manufacturing [ISO97a, ISO00a, ISO93m, ISO95h]. manuscript [Ass87a]. MAPLE [ISO97a, ISO00a, ISO93m]. markers [ISO99e]. market [FFK92]. Markup [Ass87b, Ass89, Can89, Hei94, ISO86e, ISO86b, ISOxxm, ISO00d, Ass87a, Smi87]. Marriage [Ano92]. Maryland [ITC93]. material [Ass87b, Ass89]. mathematical [ISO09c, ISO10b, ISO09b]. Mathematics [ISO19, ISO91j]. mathématique [Int91]. matrix [ISO98j]. measurements [ISO75]. measuring [ISO96j, ISO96k, ISO96n, ISO96o, ISO96q, ISO96s, ISOxxl]. MED [ISO08e]. mémoire [ISO88l]. messaging [ISO96i, ISO96k, ISO96m, ISO96o]. metafile [ISO97o]. metalanguage [ISO96t]. Method [ISO96f]. methodologies [DDLL94]. Methods [RAMP94, ISO88n, ISO94a, ISO96j, ISO96k, ISO96n, ISO96o, ISO96q, ISO96s, ISO97s, ISOxxf, ISOxxa, ISOxxl]. MHEG [Pri93]. MHS [ISO96i, ISO96k, ISO96m, ISO96o]. MHS-based [ISO96i, ISO96m, ISO96k]. Microprocessor [ISO11h, ISO20c]. Minor [ISO98w, ISO78b]. Modeling [ISO98l]. Modula [ISO96d, ISO98d, ISO98e]. Modula-2 [ISO96d, ISO98d, ISO98e]. Module [ISO95e, ISO95f]. Modules [ISO94b, ISO99-40, ISO00c, ISO11f]. Multilingual [ISO93d]. Multimedia [ISO99l, ISO99g, Kom94]. multimedia/hypermedia [Kom94]. Multiple [ISO93d]. Multiple-Octet [ISO93d]. MUMPS [ISO99q, ISO92c]. Music [ISO93a, ISOxxh].

N38 [ISO83a]. names [ISO97d, ISO98b, ISO99b]. National [Sch90]. natural [ISO09b]. NC [ISO78a, ISO78b, ISO85f, ISOxxd, ISOxxe]. NDC [ISO97o]. Need [Tri95]. networks [ISO92d, ISO99z]. Neural [ISO99z]. No [ISO98r, ISO98u, ISO98q, ISO98s, ISO98t, ISO99-32, ISO99-33, ISO99-37, ISO99-31, ISO01d, ISO72, ISO83a]. No.103 [Ano93]. No.106 [Ano91]. normalisé [ISO07b]. Number [ISO93a, ISO98c, ISO07b, ISO20b]. Numbering [ISO92b]. numerical [ISO01a, ISO06b, ISO06a, ISO78a, ISO78b, ISO85f, ISOxxd, ISOxxe]. Numéro [ISO07b].

Object [ISO08b, ISO12a, Pri93, ISO98e, ISOxxn]. objects [ISO99k, ISO99l, ISO99h, ISO99i, ISO99j, ISO00e]. Octet [ISO93d]. October [ITC93]. ODA [App91, ATL93, Mol91, NW92]. Office [ISO86b, ISO86a, ISO94k, ISO06e, App91, Can89, ISO86e]. Ohio [Ass87a]. OLB [ISO08b, ISOxxn]. Open [ISO93h, ISO96x, ISO98x, ISO99q, ISO06e, ATL93, BB91, DDLL94, LL93a, LL93b, Mol91, FFK92, ISO96h, ISO96j, ISO96l, ISO96n, ISO99n]. OpenDocument [ISO06e]. Operating [ISO79a, ISO90b, ISO96y, ISOxxl]. operation [ISOxxb]. operations [ISO98j, ISO76]. optiques [ISO88l]. Organization [ISO99-27]. Oriented [ISO98e]. OSI [ISO96h, ISO96j, ISO96l, ISO96n, Smi87]. output [ISO78a, ISO78b, ISOxxd, ISOxxe]. Overview [Ada92, Eng93, ISO93m, Kom91]. owner [ISO88a, ISO91e].

package [ISO94e, ISO94f, ISO98k]. Packages [ISO99g, ISO98j]. Page [ISO95b]. paradigm [Hil95]. Parameter [ISO97b]. Part [Ame97, Int93, Int00, ISO76, ISO79a, ISO85e, ISO86c, ISO87c, ISO87d, ISO88f, ISO88h, ISO88d, ISO89a, ISO90e, ISO90f, ISO90d, ISO91h, ISO91f, ISO91g, ISO91j, ISO92d, ISO92i, ISO93d, ISO93g, ISO93h, ISO94t, ISO94j, ISO94k, ISO94n, ISO95i, ISO95m, ISO95j, ISO95h, ISO96x, ISO96u, ISO96v, ISO97d, ISO97k, ISO97m, ISO97n, ISO97p, ISO97q, ISO97r, ISO98b, ISO98o, ISO98x, ISO99b, ISO99k, ISO99l, ISO99w, ISO99x, ISO99y, ISO99u, ISO99z, ISO99-38, ISO99-39, ISO99-27, ISO99-28, ISO99v, ISO99-30, ISO99g, ISO99-40, ISO99h, ISO99i, ISO99j, ISOxxg, ISO00a, ISO00g, ISO00e, ISO00f, ISO00c, ISO04a, ISO04c, ISO04b, ISO05a, ISO07d, ISO08a, ISO08d, ISO08e, ISO08b, ISO08c, ISO10a, ISO11b, ISO11e, ISO11f, ISO11d, ISO11c, ISO20a, ISO85f, ISO98p, ISO85d, ISO88g, ISO91d, ISO93c, ISO94a, ISO94d, ISO94h, ISO94l, ISO95d, ISO95g, ISO95k, ISO96d, ISO96f, ISO96y, ISO97o, ISO97l, ISO97s, ISO98d, ISO98e, ISO98f, ISO98g, ISO98h, ISO98i]. Part [ISO98l, ISO98p, ISO98r, ISO98u, ISO98q, ISO98s, ISO98t, ISO99a, ISO99-32, ISO99-33, ISO99-34, ISO99-35, ISO99-36, ISO99-37, ISO99-31, ISO99p, ISO99-29, ISOxxf, ISOxxa, ISOxxl, ISOxxn, ISO01c, ISO01d, ISO01a, ISO03c, ISO06b, ISO06a, ISO09b, ISO19]. Partie [Int91]. Partition [ISO06f]. Partitions [ISO06d]. Pascal [ISO90a, Eur85, ISO82, ISO88g, ISO90c, ISO91a]. PCTE [ISO98f, ISO98g, ISO98h, ISO98i]. PDF [ISO05a, ISO08a, ISO20a, ISO05a]. PDF/A [ISO05a]. PDF/A-1 [ISO05a]. People [Ano93]. Peripheral [ISO88d]. Persian [ISO99a]. Persistent [ISO99-40, ISO11f]. PHIGS [ISO90e, ISO90f, ISO91i, ISO91h, ISO94p, ISO97p, ISO97q, ISO97r, ISO98y]. PICS [ISO96x]. Picture [ISO98p]. PL [ISO92f, ISO79b]. PL/1 [ISO92f]. PL/I [ISO79b]. Plane [ISO93d]. PLUS [ISO94p]. Point [ISO11h, ISO20c, ISO94d, ISOxxi, ISO06b, ISO06a, ISO09d, ISO98-28]. Politics [Fre90]. Portable [ISO90b, ISO08a, ISO20a, ISO96y, ISO98f, ISO98g, ISO98h, ISO98i, ISOxxl]. POSIX [GL91, ISO90b, ISO96y, ISO99m, ISOxxl, RAMP94]. Post [ISO78b, ISOxxe]. Post-processor [ISO78b]. Postal [ISO93b]. pour [ISO88c, Int91]. Practice [Ano92]. Practices [Ano91]. PREMO [DDHR94]. Preparation [ISO87d, Smi86, Ass87a, ISO90h, ISO98z, ISO99-43]. Prepress [ISO98a]. preservation [ISO05a]. primitive [ISO94f]. Procedure [ISO96g]. Procedures [ISO99s, ISO88a, ISO91e]. Proceedings [IEE94, Wil94, ITC93]. process [ISO85e, Sev95]. processes [ISO85e]. Processing [ISO83b, ISO86b, ISO86a, ISO88b, ISO99n, App91, BB91, Can89, Can90, GV92, ISO76, ISO79a, ISO85d, ISO85e, ISO85g, ISO86e, ISO86c, ISO87c, ISO87d, ISO87b, ISO88f, ISO88g, ISO88h, ISO88d, ISO88l, ISO88k, ISO88i, ISO88m, ISO88a, ISO88c, ISO89a, ISO90e, ISO92i, ISO94l, ISO95d, ISO97o, ISO97p, ISO97q, ISO97r, ISO98l, ISO98p, ISO99-30, ISOxxm, ISO00d, Kan93, ISO95b, ISO96c, ISO94j, ISO99w]. processor [ISO78a, ISO78b, ISO85f, ISOxxd, ISOxxe, ISO99t]. processors [ISO88n]. producing [Mee95]. Product [ISO94a, ISO97s, ISOxxf, ISOxxa]. products [Sch88]. Profiles [ISO05d, ISO99e]. proforma [ISO96x]. Program [ISO96h, ISO96i, ISO96j, ISO96k, ISO96l, ISO96m, ISO96n, ISO96o, ISO96p, ISO96q, ISO96r, ISO96s, ISO96y, ISO99m, ISO85f, ISO95d, ISO85g, ISO89c]. Programmable [Int93, Int00]. programme [II90]. Programmer [ISO90e, ISO90f, ISO91h, ISO97p, ISO97q, ISO97r]. Programming [Ame94, Ame95, Ame97, Ame99, Int93, ISO72, ISO82, ISO91a, ISO92f, ISO94e, ISO95f, ISO97a, ISO97k, ISO99u, ISO99t, ISO99-42, ISO99-43, ISO00a, ISO00f, ISO00c, ISO00i, ISO01b, ISO02, ISO03a, ISO04a, ISO04c, ISO06c, ISO07c, ISO07d, ISO08c, ISO09c, ISO09d, ISO10b, ISO10a, ISO11g, ISO11a, ISO12b, ISO14b, ISO14a, ISO17c, ISO17b, ISO18, ISO21, ISO23, ISO24, Sch90, Eur85, Int00, Ins91, ISO83a, ISO90h, II91, ISO93m, II94, ISO98g, ISO98h, ISO98z, ISOxxb, ISO79b, ISO85d, ISO85b, ISO88n, ISO89b, ISO90g, ISO90c, ISO91b, ISO92c, ISO93l, ISO94f, ISO94g, ISO94h, ISO95e, ISO95g, ISO95l, ISO95r, ISO96d, ISO96e, ISO96f, ISO97h, ISO97j, ISO97l, ISO98d, ISO98e, ISO98j, ISO98k, ISO98m, ISO98-28, ISO98-29, ISO99f, ISO99o, ISO99p, ISO99r, ISOxxj, ISOxxk, ISOxxl, ISO00g]. project [Eng93]. Prolog [ISO95g, ISO00c]. proposal [ISO83a]. Proposed [RAMP94, ISO83a]. Protocol [ISO96x, ISO99q]. Provisions [ISO99d]. PSM [ISO99-40, ISO11f]. Public [ISO91j, ISO88a, ISO91e]. Publications [Smi86, ISO97e, ISO07b]. Publicly [Hil95]. publiques [Int91]. Publishing [Ass87a]. purpose [ISO92f].

QL [ISO08c]. Quantities [ISO19, ISO09b]. Query [ISO99-29].

R [ISO72, ISOxxb]. Radix [ISOxxi]. Radix-independent [ISOxxi]. real [ISO96z, ISO98j]. real-time [ISO96z]. Reality [ISO98l]. Really [Tri95]. Realtime [ISO99m]. recognition [ISO99x]. record [ISO85g]. Recording [ISO86d]. records [ISO78b]. Ref [ISO72]. Reference [ABM^{+}97, ISO99k, ISO00e, ISO75, ISO85f, ISO94a, ISO97s, ISOxxf, Kim92]. References [ISO87a]. REGAUT [ISO99e]. Registration [ISO91e, ISO99e, ISO88a, ISO99d, ISO99s]. Related [ISO11c, ISO00d]. Reliability [ISO97m]. Remote [ISO96x, ISO98x, ISO00h]. report [ISO06f, ISO06g]. reports [ISO92i]. Representation [ISO88e, ISO99-28, ISO00b, Gay95, ISO89c, ISO94a, ISO95a, ISO97d, ISO97s, ISO98b, ISO99b, ISOxxf, ISOxxa, ISO94n]. Requirements [BHK91, IEE94, ISO94c, RAMP94]. Resource [ISO93e]. robots [ISO92a, ISO95q, ISOxxb, ISO95q]. Role [DDLL94, LL93b]. ROM [ISO88l]. Routines [ISO08c]. Rules [ISO97e, Mee95].

s [ISO99-29]. Safety [ISO92a]. SAMeDL [ISO95f, ISO95e]. SC5 [ISO83a]. Scaling [Wil94]. Schemas [ISO11d]. Schemata [ISO11d]. Scheme [Ins91]. science [ISO91j]. sciences [ISO09b]. Scientific [GV92, Smi86]. scientifique [Int91]. SCQL [ISO99-29]. SDIF [ISO88k, Can89, ISO86a, ISO88b]. Secretariat [ISO92i]. Security [ISO04b, ISO00h, ISO98o]. Semantic [ISO99o]. Semantics [Ada92, ISO96c]. sequential [ISO85g]. Serial [ISO98c, ISO07b, ISO20b]. série [ISO07b]. series [Ass87a]. Service [ISO99-43].

Services [ISO93e, ISO00a, ISO96p, ISO96q, ISO96r, ISO96s]. Set [ISO83b, ISO93d, ISO91c]. sets [ISO91j, ISO92e, ISO98r, ISO98u, ISO98q, ISO98s, ISO98t, ISO99-32, ISO99-33, ISO99-34, ISO99-35, ISO99-36, ISO99-37, ISO99-31, ISO01c, ISO01d, ISO03c]. Seven [Mee95]. SGML [ISO88k, ISO91e, ISO91j, ISO92i, ISO00d, Ass87a, Ass87b, Ass89, Adl92, Can89, Can90, Hay94, Hei94, ISO86e, ISO86b, ISO86a, ISO87b, ISO88m, ISO88b, ISO88a, ISO88c, Int91, Kan93, MW92, NW92, Rol91, Smi86, Smi87]. SGML-based [MW92]. shape [ISO94n]. signs [ISO09b]. Simplified [ISO93c, ISO99a]. single [ISO94b, ISO98r, ISO98u, ISO98q, ISO98s, ISO98t, ISO99-32, ISO99-33, ISO99-34, ISO99-35, ISO99-36, ISO99-37, ISO99-31, ISO01c, ISO01d, ISO03c]. single-byte [ISO98r, ISO98u, ISO98q, ISO98s, ISO98t, ISO99-32, ISO99-33, ISO99-34, ISO99-35, ISO99-36, ISO99-37, ISO99-31, ISO01c, ISO01d, ISO03c]. SMDL [ISOxxh]. SNAFU [Hay94]. software [ISO93l, II94, ISO96e, ISO96f, ISO97h, ISO99-43, ISOxxk, ISOxxl, ISO07d, ISO09c, ISO09d, ISO10b, ISO17b, ISO21, ISO23, LL93a, ISO99k, ISO00e]. son [II90]. Space [ISO97b, ISO97c, ISOxxc]. spaces [ISO99e]. Spatial [ISO99g]. SPDL [ISO95b]. Special [Ass87a, ISO09c, ISO10b, Kim92]. Specialization [ISO98x, ISO96x]. Specification [Ada92, Eur85, ISO82, ISO96c, ISO97r, ISO03b, ISO05c, RAMP94, ATL93, Hil95, ISO97b, ISO98f, ISO98l, ISO98n, ISOxxc, ISO96f, ISO99o]. Specifications [ISO11c, ISO99-43]. Speech [ISO99x]. SPIFF [ISO99d, ISO99e]. Springs [IEE94]. SQL [ISO95m, ISO95f, ISO99-38, ISO99-39, ISO99g, ISO99-40, ISOxxg, ISO00h, ISO08d, ISO08e, ISO08b, ISO08c, ISO11b, ISO11e, ISO11f, ISO11d, ISO11c, ISO92g, ISO95e, ISO96x, ISO98x, ISOxxn]. SQL/Ada [ISO95f, ISO95e]. SQL/Bindings [ISOxxg]. SQL/CLI [ISO95m, ISO08d]. SQL/Foundation [ISO99-39, ISO11e]. SQL/Framework [ISO99-38, ISO11b]. SQL/JRT [ISO08c]. SQL/MED [ISO08e]. SQL/OLB [ISO08b, ISOxxn]. SQL/PSM [ISO99-40, ISO11f]. SQL/Schemata [ISO11d]. SQL/XML [ISO11c]. Standard [Ass87b, Ass89, Bri03a, Bri03b, Can89, ISO75, ISO83b, ISO86d, ISO86b, ISO92b, ISO93a, ISO98c, ISO00d, ISO04a, ISO05c, ISO07a, Rei93, RAMP94, Sch90, SP87, Ass87a, Ada92, Adl92, ATL93, BHK91, Gol91, Ins91, ISO79b, ISO83a, ISO85b, II91, II94, ISO95r, ISOxxa, ISO07b, ISO20b, Pri93, Rol91, Smi87, SBSG89, Hei94, ISO86e, ISO95b, ISOxxh]. Standardization [Dun90, FFK92, Kom91, Kom93, Kom94]. standardizing [Kim92]. Standards [Ano91, BB91, Fre90, Gol91, Hil95, ISO92i, Sev95, AW93, DDLL94, Dzi96, ISO90h, ISO98z, LL93a, LL93b, Mee95, Sch88]. standards-CAPS [AW93]. state [BB91]. Statement [ISO96x]. Std [Ins91]. still [ISO94c, ISO95c, ISO97g, ISO99e]. Stored [ISO99-40, ISO11f]. Strategy [Hay94]. strings [ISO94h, ISO00f]. Structure [ISO87a, ISO78a, ISO88l, ISO94i, ISOxxd, Rol91]. structured [Gol91, ISO97i, ISO99-29]. Structuring [ISO97f, NKN91]. Style [Ada92, ISO96c]. subdivision [ISO98b]. subdivisions [ISO97d, ISO98b, ISO99b]. subset [ISO92f]. Support [ISO88b, Can89, Can90, ISO87b, ISO88k, ISO88m, ISO88a, ISO88c, II90, ISO91e, ISO91j, Int91, ISO92i, ISO09c, ISO09d, ISO10b]. symbols [ISO09b]. Syntactic [ISO96t]. synthesis [ISO99x]. System [ISO85a, ISO88f, ISO88h, ISO88i, ISO90e, ISO90f, ISO90d, ISO90b, ISO91h, ISO96y, ISO97p, ISO97q, ISO97r, ISO99-30, Int88, ISO93l, II94, ISO96e, ISO96f, ISO97h, ISO99-43, ISOxxk, ISOxxl, ISO07d, ISO09c, ISO09d, ISO10b, ISO12a, ISO17b, ISO21, ISO23, Ame94, ISO88g, ISO91d, ISO93e, ISO94g, ISO94l, ISO95k, ISO96y, ISO97o, ISO98p, ISO99m, ISOxxl]. Systemen [GS87]. Systemes [II90, ISO86b]. Systems [GS87, ISO86b, ISO86a, ISO96x, ISO96h, ISO96j, ISO96l, ISO96n, ISO98x, ISO11h, ISO20c, App91, Can89, DDLL94, ISO85e, ISO86e, ISO86c, ISO87c, ISO87d, ISO88f, ISO88g, ISO88h, ISO88d, ISO88i, ISO89a, ISO90e, ISO93h, ISO94a, ISO97a, ISO97b, ISO97c, ISO97s, ISOxxf, ISOxxa, ISOxxc, ISO00a, LL93b, ISO99h, ISO95i, ISO00d, ISO00i].

tabular [Ass87b, Ass89]. Tag [ISO98a]. tags [ISO99e]. Tape [ISO81]. TC97 [ISO83a]. TC97/SC5 [ISO83a]. Technical [Bri03a, Bri03b, ISO92i, ISO01b, ISO07a, Wil94, ISO85e, ISO06f, ISO06g]. Techniques [ISO87b, ISO91j, ISO92i, Can90, ISO79a, ISO88c, Int91, ISO94t, ISO94i, ISO04b, ISO88m]. Technologies [Int91]. Technology [Ame94, Ame95, ISO01b, ISO07c, Ame97, ISO89c, ISO90f, ISO90d, ISO90c, ISO90b, ISO91a, ISO91b, ISO91d, ISO91h, ISO91c, ISO91e, ISO91f, ISO91g, ISO91j, II91, Int91, ISO92c, ISO92f, ISO92d, ISO92e, ISO92g, ISO93d, ISO93e, ISO93g, ISO93h, ISO93l, II94, ISO94d, ISO94e, ISO94f, ISO94g, ISO94h, ISO94l, ISO94t, ISO94j, ISO94k, ISO94n, ISO94c, ISO94i, ISO95e, ISO95d, ISO95i, ISO95b, ISO95g, ISO95k, ISO95l, ISO95m, ISO95j, ISO95h, ISO95f, ISO95c, ISO96x, ISO96c, ISO96d, ISO96e, ISO96f, ISO96g, ISO96h, ISO96i, ISO96j, ISO96k, ISO96l, ISO96m, ISO96n, ISO96o, ISO96p, ISO96q, ISO96r, ISO96s, ISO96t, ISO96y, ISO96u, ISO96v, ISO96z, ISO97k, ISO97f, ISO97i, ISO97h, ISO97j, ISO97o, ISO97l, ISO97m, ISO97n, ISO97p, ISO97q, ISO97r, ISO97g, ISO98a, ISO98w, ISO98o, ISO98x, ISO98d, ISO98e, ISO98f, ISO98g, ISO98h, ISO98i, ISO98j, ISO98k, ISO98l, ISO98n, ISO98p, ISO98r, ISO98u, ISO98q, ISO98s, ISO98t]. technology [ISO98z, ISO98-27, ISO98-28, ISO98-29, ISO99f, ISO99k, ISO99l, ISO99s, ISO99-32, ISO99-33, ISO99-34, ISO99-35, ISO99-36, ISO99-37, ISO99-31, ISO99m, ISO99n, ISO99o, ISO99p, ISO99q, ISO99w, ISO99r, ISO99x, ISO99y, ISO99u, ISO99z, ISO99-38, ISO99-39, ISO99-27, ISO99-28, ISO99v, ISO99t, ISO99-30, ISO99g, ISO99-40, ISO99e, ISO99h, ISO99i, ISO99j, ISO99-43, ISOxxh, ISOxxi, ISOxxj, ISOxxk, ISOxxl, ISOxxm, ISOxxn, ISOxxg, ISO00g, ISO00d, ISO00h, ISO00e, ISO00f, ISO00c, ISO00i, ISO01c, ISO01d, ISO01a, ISO02, ISO03c, ISO03b, ISO04a, ISO04c, ISO04b, ISO05b, ISO05d, ISO06b, ISO06c, ISO06d, ISO06a, ISO06f, ISO06g, ISO07d, ISO08d, ISO08e, ISO08b, ISO08c, ISO09c, ISO09d, ISO09b, ISO10b, ISO10a, ISO11h, ISO11g, ISO11b, ISO11e, ISO11f, ISO11d, ISO11c, ISO11a, ISO12b, ISO14b, ISO14a, ISO17c, ISO17b, ISO18, ISO20c, ISO21, ISO23, ISO24, ISO87c]. TEI [MW92]. TEI-TERM [MW92]. temperature [ISO75]. term [ISO05a, MW92]. terminology [MW92]. terms [ISO85g, ISO93g]. Test [ITC93, ISO88n, ISO96j, ISO96k, ISO96n, ISO96o, ISO96q, ISO96s, ISOxxl]. testing [DDLL94, ITC93, ISO00d, LL93a, LL93b, ISO95c]. Text [App91, ISO86b, ISO86a, ISO94j, ISOxxm, Can89, GV92, ISO88a, ISO91e, ISO97r, ISO86e]. Thai [ISO01c]. their [ISO88n, ISO89c, ISO93l, II94, ISO96e, ISO96f, ISO97d, ISO97h, ISO98b, ISO99b, ISO99-43, ISOxxk, ISOxxl, ISO07d, ISO09c, ISO09d, ISO10b, ISO17b, ISO21, ISO23]. them [ITC93]. Theory [Ano92, ISO96v]. Three [ISO88i, ISO91d]. TIDE [Eng93]. TIDE-CAPS [Eng93]. TIFF [ISO98a]. TIFF/IT [ISO98a]. time [ISO96z]. time-based [NKN91, ISO97f]. times [ISO88e, ISO00b]. Title [ISO92g, ISO97e]. titles [ISO97e]. TM [ISO08c]. tone [ISO94c, ISO95c, ISO97g, ISO99e]. Tool [ISO98f, ISO98g, ISO98h, ISO98i]. TR [ISO88m, ISO90h, ISO91j, ISO92i, ISO93l, ISO96z, ISO98z, ISO98-27, ISO98-28, ISO98-29, ISO99-43, ISO00i, ISO05d, ISO06f, ISO06g, ISO07d, ISO09d, ISO88n, ISO93m, ISO95q, ISO97s]. TR3 [Int00]. trail [ISO99-30]. Traitement [ISO86b, ISO88l, ISO88b, ISO88c, II90]. transfer [ISO97b, ISO97c, ISOxxc]. Translating [Ano92]. Translation [Ano92]. Transliteration [ISO84, ISO93c, ISO99a, ISO93c, ISO99a]. tree [ISO97i]. tree-structured [ISO97i]. type [ISO78b, ISO98j, ISO98-29]. Types [ISO08c, ISO98j, ISO99d, ISO99e].

UCS [ISO93d]. UDC [ISO72]. Unicode [BB93, Kim92]. units [ISO09b, ISO19]. Universal [BB93, ISO93d]. USA [ITC93]. usability [Dzi96]. Use [ISO05a, CBDH94, ISO00i]. used [ISO99b, ISO09b]. user [ISOxxb]. Using [ISO08c, RAMP94, Can90, ISO87b, ISO88m, ISO88c, ISO91j, Int91, ISO92i]. UTF [ISO98l]. UTF-8 [ISO98l].

v1.0 [ISO06e]. value [ISO97b, Sev95].

Varying [ISO00f, ISO94h]. vector [ISO98j]. Version [Ass87a, Ass87b, Ass89, II94, ISO95r]. versions [ISO99d]. Verteilten [GS87]. VI [ISO06d]. Vienna [ISO96f]. Virtual [ISO98l]. virtues [Hay94]. visual [ISO99k, ISO99l, ISO99h, ISO99i, ISO99j, ISO00e, ISO99i]. Vocabulary [ISO76, ISO79a, ISO85e, ISO86c, ISO87c, ISO87d, ISO88d, ISO89a, ISO90d, ISO92d, ISO93g, ISO93h, ISO94j, ISO94k, ISO95i, ISO95j, ISO95h, ISO96u, ISO96v, ISO97m, ISO97n, ISO98o, ISO99w, ISO99x, ISO99y, ISO99u, ISO99z, ISO99-27, ISO99-28, ISO99v, ISO00g, ISO85d]. volume [ISO88l]. vulnerabilities [Hay94].

WARC [ISO09a, ISO17a]. WG6 [ISO83a]. Windowing [ISO99r]. words [ISO78a, ISO97e].

X3.23 [ISO85b]. X3.23-1985 [ISO85b]. X3.53 [ISO79b]. X3.53-1976 [ISO79b]. XML [ISO11c, ISO06f]. XML-Related [ISO11c].

References

Adams:1997:FHC

[ABM^{+}97] Jeanne C. Adams, Walter S.

Brainerd, Jeanne T. Martin, Brian T. Smith, and Jerrold L. Wagener, editors. Fortran 95 Handbook: Complete ISO/ANSI Reference. MIT Press, Cambridge, MA, USA, November 1997. ISBN 0-262-51096-0. xii + 711 pp. LCCN QA76.73.F25 F6 1997. US$50.00. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0262510960.

Adachi:1992:OIS

[Ada92] J. Adachi. Overview on international standard ‘Document Style Semantics and Specification Language (DSSSL)’. Journal of the Information Processing Society of Japan = Joho Shori, 33(2):176– 184, ???? 1992. CODEN JOSHA4. ISSN 0447-8053.

Adler:1992:BSS

[Adl92] S. Adler. The birth of a standard (SGML). Journal of the American Society for Information Science, 43(8):556–558, September 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

ANSI:1994:AII

[Ame94] American National Standards Institute. ANSI/ISO/IEC 11730-1994: Information Technology — Programming Languages — Form Interface Management System. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1994. ISBN ???? LCCN ???? US$18.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ANSI%2FISO%2FIEC+11730%
2D1994.

ANSI:1995:AII

[Ame95] American National Standards Institute. ANSI/ISO/IEC 8652-1995: Information Technology — Programming Languages — Ada. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1995. ISBN ???? ???? pp. LCCN ???? US$18.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ANSI%2FISO%2FIEC+8652%2D1995.

ANSI:1997:AII

[Ame97] American National Standards Institute. ANSI/ISO/IEC 1539-1:1997: Information technology — Programming languages — Fortran — Part 1: Base language. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1997. ISBN ???? ???? pp. LCCN ???? US$18.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ANSI%2FISO%2FIEC+1539%2D1%

3A1997; http://www.fortran.
com/fortran/iso1539.html.

ANSI:1999:AII

[Ame99] American National Standards Institute. ANSI/ISO/IEC 9899-1999: Programming Languages

— C. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1999. ISBN ???? ???? pp. LCCN ???? US$18.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ANSI%2FISO%2FIEC+9899%2D1999.

Anonymous:1991:ICS

[Ano91] Anonymous, editor. IEE Colloquium on ‘Standards and Practices in Electronic Data Interchange’ (Digest No.106). IEE, London, UK, 1991. ISBN ???? LCCN ????

Anonymous:1992:TCT

[Ano92] Anonymous, editor. Translating and the Computer 13. The Theory and Practice of Machine Translation — a Marriage of Convenience? Aslib, London, UK, ???? 1992. ISBN 0-85142-295-0. LCCN ????

Anonymous:1993:ICI

[Ano93] Anonymous, editor. IEE Colloquium on ‘Information Access for People with Disability’ (Digest No.103). IEE, London, UK, 1993. ISBN ???? LCCN ????

Appelt:1991:IPT

[App91] Wolfgang Appelt. Information processing—text and office systems—office document architecture (ODA) and interchange format. Part 10: Formal specifications Standard 8613, International Organization for Standardization, Geneva, Switzerland,

1991. URL http://www.iso.ch/
cate/d15928.html;   http://
www.iso.ch/cate/d21065.html;
http://www.iso.ch/cate/d22939.
html;   http://www.iso.ch/cate/
d22940.html; http://www.iso.
ch/cate/d22941.html;   http:
//www.iso.ch/cate/d22943.html;
http://www.iso.ch/cate/d23319.
html;   http://www.iso.ch/cate/
d23323.html; http://www.iso.
ch/cate/d23410.html;   http:
//www.iso.ch/cate/d23411.html;
http://www.iso.ch/cate/d23412.
html;   http://www.iso.ch/cate/
d25650.html; http://www.iso.
ch/cate/d26050.html;   http:
//www.iso.ch/cate/d28285.html;
http://www.iso.ch/cate/d30675.
html;   http://www.iso.ch/cate/
d30676.html; http://www.iso.
ch/cate/d30677.html;   http:
//www.iso.ch/cate/d30678.html;
http://www.iso.ch/cate/d30679.
html;   http://www.iso.ch/cate/
d30681.html; http://www.iso.
ch/cate/d30682.html;   http:
//www.iso.ch/cate/d30684.html;
http://www.iso.ch/cate/d30685.
html.

AAP:1987:AAP

[Ass87a] Association of American Publishers. Association of American Publishers electronic manuscript series standard for electronic manuscript preparation and markup: an SGML application conforming to international standard ISO 8879–standard generalized markup language. Version 2.0 Dublin, Ohio: Available from the Electronic Publishing Special Interest Group, c1987. Association of American Publishers, Dublin, OH, USA, 1987. ISBN ???? various pp. LCCN Z 286 E43 A87 1987.

AAP:1987:MTM

[Ass87b] Association of American Publishers, Electronic Manuscript and Project. Markup of tabular material: an SGML application conforming to International Standard ISO 8879, Standard Generalized Markup Language, Version 2.0. Electronic manuscript series / Association of American Publishers. Association of American Publishers, Dublin, OH, USA, 1987. ISBN ???? 19 + 4 + 1 pp. LCCN ????

AAP:1989:MTM

[Ass89] Association of American Publishers, Electronic Manuscript and Project. Markup of tabular material: an SGML application conforming to International Standard ISO 8879, Standard Generalized Markup Language, Version 2.0. Electronic manuscript series / Association of American Publishers. Association of American Publishers, Dublin, OH, USA, second edition, 1989. ISBN 1-55653-085-4. iii + 25 pp. LCCN Z 286 E43 M37 1989.

Appelt:1993:FSI

[ATL93] W. Appelt and N. Tetteh-Lartey.

The formal specification of the ISO open document architecture (ODA) standard. The Computer Journal, 36(3):269–279, June 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Ayres:1993:DS

[AW93] N. Ayres and T. Wesley. Document standards-CAPS. In Anonymous [Ano93], pages 5/1–3. ISBN ???? LCCN ????

Bormann:1991:SOD

[BB91] U. Bormann and C. Bormann. Standards for open document processing: current state and future developments. Computer Networks and ISDN Systems, 21(3): 149–163, May 1991. CODEN CNISE9. ISSN 0169-7552.

Bettels:1993:UUC

[BB93] Jörgen Bettels and F. Avery Bishop. Unicode: a universal character code. Digital Technical Journal of Digital Equipment Corporation, 5(3):21–31, Summer 1993. CODEN DTJOEL. ISSN 0898-901X. URL ftp://

                                        ftp.digital.com/pub/Digital/
                                        info/DTJ/v5n3/UNICODE_A_UNIVERSAL_
                                        CHARACTE_01jan1994DTJB02P8.
                                        ps.

Blum:1991:RFI

[BHK91] Christof Blum, Georg Rainer Hofmann, and Detlef Kromker. Requirements for the first international imaging standard. IEEE Computer Graphics and Applications, 11(2):61–70, March 1991. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

BSI:2003:CSIa

[Bri03a] British Standards Institute. The C Standard: incorporating Technical Corrigendum 1. Wiley, New York, NY, USA, 2003. ISBN 0-470-84573-2. 538 pp. LCCN QA76.73.C15C185 2003. 57.90 EUR, 85.00 SFr. BS ISO/IEC 9899:1999.

BSI:2003:CSIb

[Bri03b] British Standards Institute. The C++ Standard: incorporating Technical Corrigendum 1: BS ISO. Wiley, New York, NY, USA, second edition, 2003. ISBN 0-470-84674-7. xxxiv + 782 pp. LCCN QA76.73.C153C16 2003. 57.90 EUR, 85.00 SFr.

CSA:1989:IPT

[Can89] Canadian Standards Association and Standards Council of Canada. Information processing, text and office systems: Standard Generalized Markup Language (SGML), and SGML support facilities: SGML Document Interchange Format (SDIF). National standard of Canada CAN/CSA-Z243.210-89 [CSA standard], 0317-5669. Canadian Standards Association, Rexdale, ON, Canada, 1989. ISBN ???? viii + 164 pp. LCCN

???? URL http://www.iso.ch/
cate/d16387.html;   http://
www.iso.ch/cate/d16388.html;
http://www.iso.ch/cate/d16643.
html;   http://www.iso.ch/cate/
d28557.html; http://www.iso.
ch/cate/d30235.html.

CSA:1990:IPS

[Can90] Canadian Standards Association and Standards Council of Canada. Information processing: SGML support facilities: techniques for using SGML. National Standard of Canada CAN/CSA-Z243.210.1-89 [CSA standard] 0317-5669. Canadian Standards Association, Rexdale, ON, Canada, 1990. ISBN ???? vii + 124 pp. LCCN

                                        ???? URL http://www.iso.ch/
                                        cate/d17319.html;   http://
                                        www.iso.ch/cate/d17330.html;
                                        http://www.iso.ch/cate/d17332.
                                        html.

Carr:1994:WUH

[CBDH94] L. A. Carr, D. W. Barron, H. C. Davis, and W. Hall. Why use Hy-Time? Electronic Publishing— Origination, Dissemination, and Design, 7(3):163–178, September 1994. CODEN EPODEU. ISSN 0894-3982.

Duce:1994:PIA

[DDHR94] D. A. Duce, D. J. Duke, P. J. W. Ten Hagen, and G. J. Reynolds. PREMO — an initial approach to a formal definition. Computer Graphics Forum, 13(3): C/393–406, 1994. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Deshmukh:1994:RTM

[DDLL94] G. Deshmukh, R. Deshpande, J. Leathrum, and K. Liburdy. Role of testing methodologies in open systems standards. Proceedings — International Conference on Software Engineering, pages 233–240, 1994. CODEN PCSEDE. ISSN 0270-5257. IEEE catalog number 94CH3409-0.

Dunlop:1990:IS

[Dun90] Dominic Dunlop. International standardization. ;login: the USENIX Association newsletter, 15(3):21–24, May/June 1990. CO-DEN LOGNEM. ISSN 1044-6397.

Dzida:1996:IUS

[Dzi96] Wolfgang Dzida. International usability standards. ACM Computing Surveys, 28(1):173–175, March 1996. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/
surveys/234386.html.

Engelen:1993:OTP

[Eng93] J. J. Engelen. Overview of the TIDE-CAPS project. In Anonymous [Ano93], pages 4/1–8. ISBN ???? LCCN ????

ECS:1985:SCP

[Eur85] European Committee for Standardization. Specification for computer programming language PAS-CAL: ISO 7185-1983. British Standards Institute, London, UK, 1985. ISBN 0-580-12531-9. 73 pp.

LCCN ???? URL http://www.
iso.ch/cate/d13802.html. This

European standard was accepted by CEN on 1985-08-09 … and it … exists in three official versions (English, French, German) (BS 6192: 1982, ISO 7185-1983, EN 27 185: 1985).

Fanderl:1992:ODA

[FFK92] H. Fanderl, K. Fischer, and J. Kamper. The Open Document Architecture: From standardization to the market. IBM Systems Journal, 31(4):728–754, 1992. CO-DEN IBMSA7. ISSN 0018-8670. G321-5495.

Frenkel:1990:PSE

[Fre90] Karen A. Frenkel. The politics of standards and the EC. Communications of the ACM, 33(7):40– 51, July 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/
                                        0001-0782/79208.html.

Gaylord:1995:CR

[Gay95] H. E. Gaylord. Character representation. Computers and the humanities, 29(1):51–73, 1995. CO-DEN COHUAD. ISSN 0010-4817.

Gallmeister:1991:EEP

[GL91] Bill O. Gallmeister and Chris Lanier. Early experience with POSIX 1003.4 and POSIX 1003.4 A. Proceedings — Real-Time Systems Symposium, pages 190–198 (of ix + 307), 1991. CODEN PRSYEA. ISBN 0-8186-2450-7. IEEE catalog number 91CH3090-8.

Goldfarb:1991:SHS

[Gol91] Charles F. Goldfarb. Standards: HyTime: a standard for structured hypermedia interchange. Computer, 24(8):81–84, August 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gerner:1987:KVS

[GS87] N. Gerner and O. Spaniol, editors. Kommunikation in Verteilten Systemen (Communication in Distributed Systems). Springer-Verlag Inc., New York, NY, USA, 1987. ISBN 3-540-17213-0. LCCN TK5105.5 .G561 1987.

Goossens:1992:STP

[GV92] M. Goossens and E. Van Herwijnen. Scientific text processing. International Journal of Modern Physics C [Physics and Computers], 3(3):479–546, June 1992. CO-DEN IJMPEO. ISSN 0129-1831 (print), 1793-6586 (electronic).

Hayhoe:1994:SSV

[Hay94] George F. Hayhoe. Strategy or SNAFU? the virtues and vulnerabilities of SGML. In Wilkes [Wil94], pages 378–379. ISBN 0-7803-1936-2. LCCN T10.5.I54 1994. IEEE cat no. 94CH3442-1.

Heimburger:1994:ISG

[Hei94] Anneli Heimburger. Introduction to Standard Generalized Markup Language (SGML). Microcomputers for Information Management, 11(4):239–260, December 1994. CODEN MIIMEW. ISSN 0742-2342.

Hill:1995:SPA

[Hil95] J. L. Hill. Standards: Publicly available specification. A new paradigm for developing international standards. Computer, 28 (10):97–98, October 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

IEEE:1994:PFI

[IEE94] IEEE, editor. Proceedings: the First International Conference on Requirements Engineering, April 18–22, 1994, Colorado Springs, Colorado, International Conference on Requirements Engineering 1994; 1st. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5480-5, 0-8186-5481-3. LCCN QA76.758.I5717 1994.

ISO:1990:IPS

[II90] International Organization for Standardization and International Electrotechnical Commission. Systemes de traitement de l’information–infographie–interfaces langage entre un programme d’application et son support graphique. International standard; ISO/IEC 9593-1. International Organization for Standardization, Geneva, Switzerland, 1990. ISBN ???? v + 213 pp. LCCN

???? URL http://www.iso.ch/
cate/d17366.html;   http://
www.iso.ch/cate/d17368.html;
http://www.iso.ch/cate/d17369.
html;   http://www.iso.ch/cate/
d1911.html; http://www.iso.
ch/cate/d21152.html;   http:
//www.iso.ch/cate/d21312.html;
http://www.iso.ch/cate/d22441.
html;   http://www.iso.ch/cate/
d22442.html; http://www.iso.
ch/cate/d22878.html;   http:
//www.iso.ch/cate/d22879.html;
http://www.iso.ch/cate/d22880.
html;   http://www.iso.ch/cate/
d27266.html.

ISO:1991:ISI

[II91] International Organization for Standardization and International Electrotechnical Commission. International standard: information, technology, programming languages, Fortran. International Organization for Standardization, Geneva, Switzerland, second edition, 1991. ISBN ???? xvii + 369 pp. LCCN

???? URL http://www.iso.ch/
cate/d26933.html;   http://
www.iso.ch/cate/d26934.html;
http://www.iso.ch/cate/d29926.
html.

ISO:1994:IID

[II94] International Organization for Standardization and International Electrotechnical Commission. ISO/IEC DIS 8652: information technology — programming languages — their environments and system software interfaces, programming language Ada, language and standard libraries, draft, version 5.0, 1 June 1994, IR-MA-1363-4. Draft international standard. Intermetrics, Inc., Cambridge, MA, USA, 1994. ISBN ???? various pp. LCCN ????

URL http://www.iso.ch/cate/
d22983.html.

IEEE:1991:ISS

[Ins91] Institute of Electrical and Electronics Engineers and IEEE Computer Society. Microprocessor and Microcomputer Standards and Subcommittee. IEEE standard for the Scheme programming language: IEEE Std 1178-1990. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 1-55937-125-0. various pp. LCCN ???? Approved December 10, 1990, IEEE Standards Board.

ISO:1988:IPSa

[Int88] International Organization for Standardization.Technical and Information Processing Systems Committee ISO/TC 97. Information processing systems: computer graphics, graphical kernel system (GKS) language bindings. International standard; ISO 8651-1. International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? iii + 116 pp. LCCN ????

ISO:1991:ITS

[Int91] International Organization for Standardization and International Electrotechnical Commission. Information technology–SGML support facilities–techniques for using SGML. technologies de l’information: facilités de support pour SGML: techniques d’utilisation du SGML: Partie 13, ensembles d’entités publiques pour les domaines mathématique et scientifique. Technical report ISO/IEC TR 9573-13, International Organization for Standardization, Geneva, Switzerland, 1991. iii + 82

pp. URL http://www.iso.ch/
cate/d17319.html;   http://
www.iso.ch/cate/d17330.html;
http://www.iso.ch/cate/d17332.
html.

IEC:1993:IEP

[Int93] International Electrotechnical Commission. IEC 61131-3 Ed. 1.0 en:1993: Programmable controllers — Part 3: Programming languages. International Electrotechnical Commission, Geneva, Switzerland, 1993. ISBN ???? ???? pp. LCCN ???? US$141.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=IEC+61131%2D3+Ed%2E+1%2E0+
en%3A1993.

IEC:2000:ITE

[Int00] International Electrotechnical Commission. IEC 61131-8 TR3 Ed. 1.0 b:2000: Programmable controllers — Part 8: Guidelines for the application and implementation of programming languages. International Electrotechnical Commission, Geneva, Switzerland, 2000. ISBN ???? ???? pp. LCCN ???? US$133.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=IEC+61131%2D8+TR3+Ed%2E+
1%2E0+b%3A2000.

ISO:1972:PLA

[ISO72] ISO. Programming language: AL-GOL: UDC 681.3.04. Ref. no. ISO/R 1538-1972 (E), volume R 1538 of ISO recommendation. International Organization for Standardization, Geneva, Switzerland, 1972. ISBN ???? 82 pp. LCCN ????

ISO:1975:ISR

[ISO75] ISO. ISO 1:1975: Standard reference temperature for industrial length measurements. International Organization for Standardization, Geneva, Switzerland, 1975. ISBN ???? 1 pp. LCCN ???? CHF 33. URL http://www.

                                        iso.ch/cate/d3561.html.

ISO:1976:IDP

[ISO76] ISO. ISO 2382-2:1976: Data processing — Vocabulary — Part 2: Arithmetic and logic operations. International Organization for Standardization, Geneva, Switzerland, 1976. ISBN ???? 31 pp. LCCN ???? CHF 110.

                                        URL http://www.iso.ch/cate/
                                        d7230.html.

ISO:1978:INCa

[ISO78a] ISO. ISO 3592:1978: Numerical control of machines — NC processor output — Logical structure (and major words). International Organization for Standardization, Geneva, Switzerland, 1978. ISBN ???? 30 pp. LCCN ???? CHF

                                        110. URL http://www.iso.ch/
                                        cate/d9003.html.

ISO:1978:INCb

[ISO78b] ISO. ISO 4343:1978: Numerical control of machines — NC processor output — Minor elements of 2000-type records (Postprocessor commands). International Organization for Standardization, Geneva, Switzerland, 1978. ISBN ???? 92 pp. LCCN ???? CHF 176. URL http://www.

                                        iso.ch/cate/d10221.html.

ISO:1979:IDP

[ISO79a] ISO. ISO 2382-10:1979: Data processing — Vocabulary — Part 10: Operating techniques and facilities. International Organization for Standardization, Geneva, Switzerland, 1979. ISBN ???? 16 pp. LCCN ???? CHF 74.

URL http://www.iso.ch/cate/
d7247.html.

ISO:1979:IPL

[ISO79b] ISO. ISO 6160:1979: Programming languages — PL/I (Endorsement of ANSI standard X3.53-1976). International Organization for Standardization, Geneva, Switzerland, 1979. ISBN ???? LCCN ???? CHF 13; US$10.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO+6160%3A1979;   http:
//www.iso.ch/cate/d12406.html.

ISO:1981:FBI

[ISO81] ISO. Format for Bibliographic Information Interchange on Magnetic Tape: ISO 2709-1981. International Organization for Standardization, Geneva, Switzerland, 1981. ISBN ???? LCCN ????

URL http://www.iso.ch/cate/
d7675.html.

ISO:1982:SCP

[ISO82] ISO. Specification for Computer Programming Language Pascal, ISO 7185-1982. International Organization for Standardization, Geneva, Switzerland, 1982. ISBN ???? LCCN ???? URL http://

www.iso.ch/cate/d13802.html.

ISO:1983:DPS

[ISO83a] ISO. Draft proposed standard programming language APL: Document no. ISO TC97/SC5 and WG6 N38. ISO draft proposal ISO/DP 8485, volume 14(2) of APL quote quad. ACM Press, New York, NY 10036, USA, June 30, 1983. ISBN 0-89791-149-0. viii + 308 pp. LCCN ???? Fifth working draft standard for programming language APL, International Standards Organization Document number ISO TC 97/SC5 WG6 N38, June 30, 1983.

ISO:1983:ISB

[ISO83b] ISO. ISO Standard 646, 7-Bit Coded Character Set for Information Processing Interchange. International Organization for Standardization, Geneva, Switzerland, second edition, 1983. ISBN ????

                                        LCCN ???? URL http://www.
                                        iso.ch/cate/d4777.html. Also

available as ECMA-6.

ISO:1984:IDT

[ISO84] ISO. ISO 233:1984: Documentation — Transliteration of Arabic characters into Latin characters. International Organization for Standardization, Geneva, Switzerland, 1984. ISBN ???? 7 pp. LCCN ???? CHF 50.

                                        URL http://www.iso.ch/cate/
                                        d4117.html.

ISO:1985:GKS

[ISO85a] ISO. The Graphical Kernel System (GKS): ISO 7942. International Organization for Standardization, Geneva, Switzerland, 1985. ISBN ???? LCCN ???? URL http://

                                        www.iso.ch/cate/d14915.html.

ISO:1985:IPL

[ISO85b] ISO. ISO 1989:1985: Programming languages — COBOL (Endorsement of ANSI standard X3.23-1985). International Organization for Standardization, Geneva, Switzerland, 1985. ISBN ???? 1 pp. LCCN ???? CHF 13; US$10.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO+1989%3A1985;   http:
//www.iso.ch/cate/d6724.html.

ISO:1985:IAC

[ISO85c] ISO. ISO 1989:1985/Amd 2:1994: Correction and clarification amendment for COBOL. International Organization for Standardization, Geneva, Switzerland, 1985. ISBN ???? 1 pp. LCCN ????

CHF 13. URL http://www.iso.
ch/cate/d21818.html. Available

in English only.

ISO:1985:IDPa

[ISO85d] ISO. ISO 2382-15:1985: Data processing — Vocabulary — Part 15: Programming languages. International Organization for Standardization, Geneva, Switzerland, 1985. ISBN ???? 46 pp. LCCN ???? CHF 136. URL http://www.

iso.ch/cate/d7256.html.

ISO:1985:IDPb

[ISO85e] ISO. ISO 2382-21:1985: Data processing — Vocabulary — Part 21: Interfaces between process computer systems and technical processes. International Organization for Standardization, Geneva, Switzerland, 1985. ISBN ????

16 pp. LCCN ???? CHF 74.

                                        URL http://www.iso.ch/cate/
                                        d7265.html.

ISO:1985:INC

[ISO85f] ISO. ISO 4342:1985: Numerical control of machines — NC processor input — Basic part program reference language. International Organization for Standardization, Geneva, Switzerland, 1985. ISBN ???? 278 pp. LCCN ???? CHF

                                        252. URL http://www.iso.ch/
                                        cate/d10220.html.

ISO:1985:IIP

[ISO85g] ISO. ISO 6593:1985: Information processing — Program flow for processing sequential files in terms of record groups. International Organization for Standardization, Geneva, Switzerland, 1985. ISBN ???? 6 pp. LCCN ???? CHF

                                        44. URL http://www.iso.ch/
                                        cate/d13001.html.

ISO:1986:IPTb

[ISO86a] ISO. Information Processing: Text and Office Systems: (SGML) Document Interchange Format (SDIF). International Organization for Standardization, Geneva, Switzerland, 1986. ISBN ???? 4 pp. LCCN QA76.73.S44 I56 1986. Draft International Standard ISO/DIS 9069.

ISO:1986:IPTa

[ISO86b] ISO. Information Processing, Text and Office Systems, Standard Generalized Markup Language (SGML) = Traitement de l’information, systèmes bureautiques, langage standard généralisé de balisage (SGML). First edition, 1986-10-15. International Organization for Standardization, Geneva, Switzerland, 1986. ISBN ???? xi + 155 pp. LCCN QA76.9.T48 I57 1986.

URL http://www.iso.ch/cate/

d16387.html. International Standard ISO 8879-1986. Federal information processing standard; FIPS PUB 152.

ISO:1986:IIPb

[ISO86c] ISO. ISO 2382-22:1986: Information processing systems — Vocabulary — Part 22: Calculators. International Organization for Standardization, Geneva, Switzerland, 1986. ISBN ???? 19 pp. LCCN ???? CHF 86. URL http://www.

iso.ch/cate/d7266.html.

ISO:1986:IDI

[ISO86d] ISO. ISO 3901:1986 Documentation — International Standard Recording Code (ISRC). International Organization for Standardization, Geneva, Switzerland, 1986. ISBN ???? 4 pp. LCCN ???? CHF 38.00. URL http://

www.iso.ch/cate/d9515.html.

ISO:1986:IIPa

[ISO86e] ISO. ISO 8879:1986: Information processing — Text and office systems — Standard Generalized Markup Language (SGML). International Organization for Standardization, Geneva, Switzerland, August 17, 1986. ISBN ???? 155 pp. LCCN ???? CHF 212.

URL http://www.iso.ch/cate/
d16387.html.

ISO:1987:DBR

[ISO87a] ISO. Documentation — Bibliographic References — Content, Form and Structure (ISO 690). International Organization for Standardization, Geneva, Switzerland, 1987. ???? pp.

ISO:1987:IPS

[ISO87b] ISO. Information processing: SGML support facilities: Techniques for using SGML. International Organization for Standardization, Geneva, Switzerland, 1987. ISBN ???? vi + 96 pp. LCCN ????

ISO:1987:IIPa

[ISO87c] ISO. ISO 2382-3:1987: Information processing systems — Vocabulary — Part 3: Equipment technology. International Organization for Standardization, Geneva, Switzerland, 1987. ISBN ???? 13 pp. LCCN ???? CHF 68.

                                        URL http://www.iso.ch/cate/
                                        d7232.html.

ISO:1987:IIPb

[ISO87d] ISO. ISO 2382-6:1987: Information processing systems — Vocabulary — Part 6: Preparation and handling of data. International Organization for Standardization, Geneva, Switzerland, 1987. ISBN ???? 15 pp. LCCN ???? CHF

                                        74. URL http://www.iso.ch/
                                        cate/d7238.html.

ISO:1988:IPSc

[ISO88a] ISO. Information processing, SGML support facilities: registration procedures for public text owner identifiers. International standard ISO 9070. International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? iii + 5 pp. LCCN ????

URL http://www.iso.ch/cate/
d16645.html.

ISO:1988:IPSb

[ISO88b] ISO. Information Processing, SGML Support Facilities: SGML Document Interchange Format (SDIF) = Traitement de l’information, bureaucratique: format d’échange de document SGML (SDIF). International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? iii + 6 pp. LCCN QA76.73.S44 I55 1988. International Standard ISO/DIS 9069.

ISO:1988:IPSd

[ISO88c] ISO. Information processing, SGML support facilities: Traitement de l’information, facilités de support pour SGML: techniques for using SGML = techniques d’utilisation du SGML. Technical report ISO/IEC TR 9573, International Organization for Standardization, Geneva, Switzerland, 1988. vii + 124

pp. URL http://www.iso.ch/
cate/d17319.html;   http://
www.iso.ch/cate/d17330.html;
http://www.iso.ch/cate/d17332.
html.

ISO:1988:IIPd

[ISO88d] ISO. ISO 2382-12:1988: Information processing systems — Vocabulary — Part 12: Peripheral equipment. International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? 46 pp. LCCN ???? CHF 136.

                                        URL http://www.iso.ch/cate/
                                        d7251.html.

ISO:1988:IDE

[ISO88e] ISO. ISO 8601:1988. Data elements and interchange formats — Information interchange — Representation of dates and times. International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? 14 pp. LCCN ???? CHF 68. URL http://

                                        www.iso.ch/cate/d26780.html.

See also 1-page correction, ISO 8601:1988/Cor 1:1991.

ISO:1988:IIPa

[ISO88f] ISO. ISO 8651-1:1988: Information processing systems — Computer graphics — Graphical Kernel System (GKS) language bindings — Part 1: FORTRAN. International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? 116 pp. LCCN ???? CHF 188. URL http://

                                        www.ansi.org/;   http://www.
                                        iso.ch/cate/d16024.html.

ISO:1988:IIPb

[ISO88g] ISO. ISO 8651-2:1988: Information processing systems — Computer graphics — Graphical Kernel System (GKS) language bindings — Part 2: Pascal. International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? 168 pp. LCCN ???? CHF 212. URL http://www.

                                        iso.ch/cate/d16025.html.

ISO:1988:IIPc

[ISO88h] ISO. ISO 8651-3:1988: Information processing systems — Computer graphics — Graphical Kernel System (GKS) language bindings — Part 3: Ada. International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? 184 pp. LCCN ???? CHF

224. URL http://www.iso.ch/
cate/d16026.html.

ISO:1988:IIPg

[ISO88i] ISO. ISO 8805:1988: Information processing systems — Computer graphics — Graphical Kernel System for Three Dimensions (GKS-3D) functional description. International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? 379 pp. LCCN ???? CHF 306. URL http://www.

iso.ch/cate/d16253.html.

ISO:1988:IA

[ISO88j] ISO. ISO 8879:1986/Amd 1:1988.

International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? 15 pp. LCCN ???? CHF 74.

URL http://www.iso.ch/cate/
d16388.html.

ISO:1988:IIPf

[ISO88k] ISO. ISO 9069:1988: Information processing — SGML support facilities — SGML Document Interchange Format (SDIF). International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? 8 pp. LCCN ???? CHF 50. URL http://www.

iso.ch/cate/d16643.html.

ISO:1988:IIPe

[ISO88l] ISO. ISO 9660:1988: Information processing: volume and file structure of CD-ROM for information interchange = Traitement de l’information: structure de volume et de fichier des disques optiques compacts a mémoire fixe (CD-ROM) destinés a l’échange d’information. International Organization for Standardization, Geneva, Switzerland, corrected and reprinted edition, 1988. ISBN ???? x + 31 pp. LCCN ???? CHF 110. URL http://www.

                                        iso.ch/cate/d17505.html.

ISO:1988:IIT

[ISO88m] ISO. ISO/IEC TR 9573:1988: Information processing — SGML support facilities — Techniques for using SGML. International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? 124 pp. LCCN ???? CHF

                                        200. URL http://www.iso.ch/
                                        cate/d17319.html.

ISO:1988:ITP

[ISO88n] ISO. ISO/TR 9547:1988: Programming language processors — Test methods — Guidelines for their development and acceptability. International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? 6 pp. LCCN ???? CHF 44.

                                        URL http://www.iso.ch/cate/
                                        d17292.html.

ISO:1989:IIP

[ISO89a] ISO. ISO 2382-19:1989: Information processing systems — Vocabulary — Part 19: Analog computing. International Organization for Standardization, Geneva, Switzerland, 1989. ISBN ???? 10 pp. LCCN ???? CHF 56. URL http:/

/www.iso.ch/cate/d7263.html.

ISO:1989:IPL

[ISO89b] ISO. ISO 8485:1989: Programming languages — APL. International Organization for Standardization, Geneva, Switzerland, 1989. ISBN ???? 259 pp. LCCN ???? CHF 252; US$170.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO+8485%3A1989;   http:
//www.iso.ch/cate/d15692.html.

ISO:1989:III

[ISO89c] ISO. ISO/IEC 8631:1989: Information technology — Program constructs and conventions for their representation. International Organization for Standardization, Geneva, Switzerland, 1989. ISBN ???? 7 pp. LCCN ???? CHF

50. URL http://www.iso.ch/
cate/d15968.html.

ISO:1990:EPI

[ISO90a] ISO. Extended Pascal ISO 10206:1990. International Organization for Standardization, Geneva, Switzerland, 1990. ISBN ???? xii + 218 pp. LCCN

???? URL http://www.iso.ch/
cate/d18237.html. Available in

PostScript for personal use only at http://pascal.miningco.com/msub1.htm.

ISO:1990:ITP

[ISO90b] ISO. Information technology– Portable Operating System Interface (POSIX). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 1-55937-061-0 (v. 1). various pp. LCCN QA76.76.O63 I58 1990. International standard ISO/IEC 9945. IEEE Std 1003.1-1990 (revision of IEEE Std 1003.1-1988). Part 1. System application program interface (API) [C language].

ISO:1990:IITa

[ISO90c] ISO. ISO 7185:1990: Information technology — Programming languages — Pascal. International Organization for Standardization, Geneva, Switzerland, 1990. ISBN ???? 78 pp. LCCN ???? CHF 164; US$112.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO+7185%3A1990;   http:
//www.iso.ch/cate/d13802.html.

ISO:1990:IIIc

[ISO90d] ISO. ISO/IEC 2382-20:1990: Information technology — Vocabulary — Part 20: System development. International Organization for Standardization, Geneva, Switzerland, 1990. ISBN ???? 15 pp. LCCN ???? CHF 74.

URL http://www.iso.ch/cate/
d7264.html.

ISO:1990:IIIa

[ISO90e] ISO. ISO/IEC 9593-1:1990: Information processing systems — Computer graphics — Programmer’s Hierarchical Interactive Graphics System (PHIGS) language bindings — Part 1:

FORTRAN. International Organization for Standardization, Geneva, Switzerland, 1990. ISBN ???? 213 pp. LCCN ????

CHF 236. URL http://www.
ansi.org/;   http://www.iso.
ch/cate/d17366.html. Available

in English only. ISO/IEC JTC 1/SC 24/WG 4.

ISO:1990:IIIb

[ISO90f] ISO. ISO/IEC 9593-3:1990: Information technology — Computer graphics — Programmer’s Hierarchical Interactive Graphics System (PHIGS) language bindings — Part 3: ADA. International Organization for Standardization, Geneva, Switzerland, 1990. ISBN ???? 285 pp. LCCN ???? CHF

270. URL http://www.iso.ch/
cate/d17368.html. Available in

English only.

ISO:1990:IIP

[ISO90g] ISO. ISO/IEC 9899:1990: Programming languages — C. International Organization for Standardization, Geneva, Switzerland, 1990. ISBN ???? ???? pp. LCCN

???? URL http://www.iso.ch/
cate/d17782.html.

ISO:1990:IITb

[ISO90h] ISO. ISO/IEC TR 10034:1990:

Guidelines for the preparation of conformity clauses in programming language standards. International Organization for Standardization, Geneva, Switzerland, 1990. ISBN ???? 4 pp. LCCN ???? CHF 38. URL http://www.

iso.ch/cate/d17996.html.

ISO:1991:IIIa

[ISO91a] ISO. ISO/IEC 10206:1991 Information technology — Programming languages — Extended Pascal. International Organization for Standardization, Geneva, Switzerland, 1991. ISBN ???? 214 pp. LCCN ???? CHF 236.

                                        URL http://www.iso.ch/cate/

d18237.html. Available in English only.

ISO:1991:IIIb

[ISO91b] ISO. ISO/IEC 10279:1991: Information technology — Programming languages — Full BASIC. International Organization for Standardization, Geneva, Switzerland, 1991. ISBN ???? ii + 14 pp. LCCN ???? CHF 68; US$46.00.

                                        URL http://webstore.ansi.
                                        org/ansidocstore/product.asp?
                                        sku=ISO%2FIEC+10279%3A1991;
                                        http://www.iso.ch/cate/d18321.
                                        html.

ISO:1991:IIIe

[ISO91c] ISO. ISO/IEC 646:1991 Information technology — ISO 7-bit coded character set for information interchange. International Organization for Standardization, Geneva, Switzerland, 1991. ISBN ???? 15 pp. LCCN ???? CHF 74.

                                        URL http://www.iso.ch/cate/
                                        d4777.html.

ISO:1991:IIIc

[ISO91d] ISO. ISO/IEC 8806-4:1991: Information technology — Computer graphics — Graphical Kernel System for Three Dimensions (GKS-3D) language bindings — Part 4:

C. International Organization for Standardization, Geneva, Switzerland, 1991. ISBN ???? 264 pp. LCCN ???? CHF 252.

URL http://www.iso.ch/cate/
d16257.html.

ISO:1991:IIIf

[ISO91e] ISO. ISO/IEC 9070:1991 Information technology — SGML support facilities — Registration procedures for public text owner identifiers. International Organization for Standardization, Geneva, Switzerland, 1991. ISBN ???? 12 pp. LCCN ???? CHF 62.

URL http://www.iso.ch/cate/

d16645.html. Available in English only.

ISO:1991:IIIg

[ISO91f] ISO. ISO/IEC 9541-1:1991: Information technology — Font information interchange — Part 1: Architecture. International Organization for Standardization, Geneva, Switzerland, 1991. ISBN ???? 81 pp. LCCN ???? CHF

176. URL http://www.iso.ch/
cate/d17277.html.

ISO:1991:IIIh

[ISO91g] ISO. ISO/IEC 9541-2:1991: Information technology — Font information interchange — Part 2: Interchange Format. International Organization for Standardization, Geneva, Switzerland, 1991. ISBN ???? 28 pp. LCCN ???? CHF

104. URL http://www.iso.ch/
cate/d17278.html.

ISO:1991:IIId

[ISO91h] ISO. ISO/IEC 9593-4:1991: Information technology — Computer graphics — Programmer’s Hierarchical Interactive Graphics System (PHIGS) language bindings — Part 4: C. International Organization for Standardization, Geneva, Switzerland, 1991. ISBN ???? 309 pp. LCCN ???? CHF

                                        270. URL http://www.iso.ch/
                                        cate/d17369.html. Available in

English only.

ISO:1991:IIA

[ISO91i] ISO. ISO/IEC 9593-4:1991/Amd 2:1998: Incorporation of PHIGS amendments. International Organization for Standardization, Geneva, Switzerland, 1991. ISBN ???? 355 pp. LCCN ???? CHF

                                        288. URL http://www.iso.ch/
                                        cate/d27266.html. Available in

English only.

ISO:1991:IIT

[ISO91j] ISO. ISO/IEC TR 9573-13:1991: Information technology — SGML support facilities — Techniques for using SGML — Part 13: Public entity sets for mathematics and science. International Organization for Standardization, Geneva, Switzerland, 1991. ISBN ???? 82 pp. LCCN ???? CHF 176.

                                        URL http://www.iso.ch/cate/
                                        d17332.html.

ISO:1992:IMI

[ISO92a] ISO. ISO 10218:1992: Manipulating industrial robots — Safety. International Organization for Standardization, Geneva, Switzerland, 1992. ISBN ???? 10 pp. LCCN ???? CHF 44. URL http://www.

iso.ch/cate/d18252.html.

ISO:1992:IID

[ISO92b] ISO. ISO 2108:1992: Information and documentation — International Standard Book Numbering (ISBN). International Organization for Standardization, Geneva, Switzerland, 1992. ISBN ???? 3 pp. LCCN ???? CHF 38.00. URL http://

www.bowker.com/standards/;
http://www.isbn.org/; http:/
/www.iso.ch/cate/d6898.html.

ISO:1992:IIIa

[ISO92c] ISO. ISO/IEC 11756:1992: Information technology — Programming languages — MUMPS. International Organization for Standardization, Geneva, Switzerland, 1992. ISBN ???? ???? pp. LCCN

???? URL http://www.iso.ch/
cate/d19978.html.

ISO:1992:IIIc

[ISO92d] ISO. ISO/IEC 2382-25:1992: Information technology — Vocabulary — Part 25: Local area networks. International Organization for Standardization, Geneva, Switzerland, 1992. ISBN ???? 12 pp. LCCN ???? CHF 98.

URL http://www.iso.ch/cate/
d7269.html.

ISO:1992:IIId

[ISO92e] ISO. ISO/IEC 6429:1992: Information technology — Control functions for coded character sets. International Organization for Standardization, Geneva, Switzerland, third edition, 1992. ISBN ???? 89 pp. LCCN ????

                                        CHF 176. URL http://www.
                                        iso.ch/cate/d12782.html. Su-

percedes ANSI X3.64.

ISO:1992:IIIb

[ISO92f] ISO. ISO/IEC 6522:1992: Information technology — Programming languages — PL/1 general purpose subset. International Organization for Standardization, Geneva, Switzerland, 1992. ISBN ???? 443 pp. LCCN ???? CHF 320; US$218.00.

                                        URL http://webstore.ansi.
                                        org/ansidocstore/product.asp?
                                        sku=ISO%2FIEC+6522%3A1992; http:
                                        //www.iso.ch/cate/d12907.html.

Available in English only.

ISO:1992:IITa

[ISO92g] ISO. ISO/IEC 9075:1992: Title: Information technology — Database languages — SQL. International Organization for Standardization, Geneva, Switzerland, 1992. ISBN ???? 587 pp. LCCN ???? CHF 348.

                                        URL http://www.iso.ch/cate/

d16663.html. Available in English only.

ISO:1992:IIC

[ISO92h] ISO. ISO/IEC 9541-1:1991/Cor 1:1992. International Organization for Standardization, Geneva, Switzerland, 1992. ISBN ???? 1 pp. LCCN ???? URL http://

                                        www.iso.ch/cate/d20932.html.

Available in English only.

ISO:1992:IITb

[ISO92i] ISO. ISO/IEC TR 9573-11:1992 Information processing — SGML support facilities — Techniques for using SGML — Part 11: Application at ISO Central Secretariat for International Standards and Technical reports. International Organization for Standardization, Geneva, Switzerland, 1992. ISBN ???? 73 pp. LCCN ???? CHF

164. URL http://www.iso.ch/
cate/d17330.html.

ISO:1993:IIDb

[ISO93a] ISO. ISO 10957:1993: Information and documentation — International Standard Music Number (ISMN). International Organization for Standardization, Geneva, Switzerland, 1993. ISBN ???? 4 pp. LCCN ???? CHF 38.00.

URL http://www.iso.ch/cate/
d18931.html.

ISO:1993:IPA

[ISO93b] ISO. ISO 11180:1993: Postal addressing. International Organization for Standardization, Geneva, Switzerland, 1993. ISBN ???? 8 pp. LCCN ???? URL http://

www.iso.ch/cate/d19185.html.

ISO:1993:IIDa

[ISO93c] ISO. ISO 233-2:1993: Information and documentation — Transliteration of Arabic characters into Latin characters — Part 2: Arabic language — Simplified transliteration. International Organization for Standardization, Geneva, Switzerland, 1993. ISBN ???? 6 pp. LCCN ???? CHF 44.

                                        URL http://www.iso.ch/cate/
                                        d4118.html.

ISO:1993:IIIa

[ISO93d] ISO. ISO/IEC 10646-1:1993, Information technology. Universal Multiple-Octet Coded Character Set (UCS). Part 1: Architecture and Basic Multilingual Plane. International Organization for Standardization, Geneva, Switzerland, 1993. ISBN ???? 754 pp. LCCN ???? CHF 360.

                                        URL http://www.iso.ch/cate/

d18741.html. About two dozen corrections and amendments to this Standard have been published.

ISO:1993:IIIb

[ISO93e] ISO. ISO/IEC 10728:1993: Information technology — Information Resource Dictionary System (IRDS) Services Interface. International Organization for Standardization, Geneva, Switzerland, 1993. ISBN ???? 108 pp. LCCN ???? CHF 188. URL http://www.

                                        iso.ch/cate/d18821.html.

ISO:1993:IIA

[ISO93f] ISO. ISO/IEC 10728:1993/Amd 2:1996: Ada language binding. International Organization for Standardization, Geneva, Switzerland, 1993. ISBN ???? 20 pp. LCCN ???? CHF 13. URL http://

                                        www.iso.ch/cate/d24342.html.

Available in English only.

ISO:1993:IIIc

[ISO93g] ISO. ISO/IEC 2382-1:1993 Information technology — Vocabulary — Part 1: Fundamental terms. International Organization for Standardization, Geneva, Switzerland, 1993. ISBN ???? 32 pp. LCCN ???? CHF 110. URL http://www.

iso.ch/cate/d7229.html.

ISO:1993:IIId

[ISO93h] ISO. ISO/IEC 2382-26:1993: Information technology — Vocabulary — Part 26: Open systems interconnection. International Organization for Standardization, Geneva, Switzerland, 1993. ISBN ???? 24 pp. LCCN ???? CHF

98. URL http://www.iso.ch/
cate/d7270.html.

ISO:1993:IICa

[ISO93i] ISO. ISO/IEC 9541-2:1991/Cor 1:1993. International Organization for Standardization, Geneva, Switzerland, 1993. ISBN ???? 4 pp. LCCN ???? URL http://

www.iso.ch/cate/d20933.html.

ISO:1993:IICb

[ISO93j] ISO. ISO/IEC 9593-1:1990/Cor 1:1993. International Organization for Standardization, Geneva, Switzerland, 1993. ISBN ???? 1 pp. LCCN ???? URL http://

www.iso.ch/cate/d22441.html.

Available in English only.

ISO:1993:IICc

[ISO93k] ISO. ISO/IEC 9593-3:1990/Cor 1:1993. International Organization for Standardization, Geneva, Switzerland, 1993. ISBN ???? 1 pp. LCCN ???? URL http://

www.iso.ch/cate/d22442.html.

Available in English only.

ISO:1993:IIT

[ISO93l] ISO. ISO/IEC TR 10182:1993: Information technology — Programming languages, their environments and system software interfaces — Guidelines for language bindings. International Organization for Standardization, Geneva, Switzerland, 1993. ISBN ???? 44 pp. LCCN ???? CHF 128; US$88.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+TR+10182%3A1993;
http://www.iso.ch/cate/d18201.
html.

ISO:1993:ITM

[ISO93m] ISO. ISO/TR 12186:1993: Manufacturing automation programming language environment overview (MAPLE). International Organization for Standardization, Geneva, Switzerland, 1993. ISBN ???? 27 pp. LCCN ???? CHF

104. URL http://www.iso.ch/
cate/d2284.html.

ISO:1994:IIAa

[ISO94a] ISO. ISO 10303-11:1994: Industrial automation systems and integration — Product data representation and exchange — Part 11: Description methods: The EX-PRESS language reference manual. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 208 pp. LCCN ???? CHF 224.

URL http://www.iso.ch/cate/
d18348.html.

ISO:1994:IIAd

[ISO94b] ISO. ISO/IEC 10279:1991/Amd 1:1994: Modules and single character input enhancement. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 3 pp. LCCN ???? CHF 13. URL http://www.

iso.ch/cate/d20742.html.

ISO:1994:IIIk

[ISO94c] ISO. ISO/IEC 10918-1:1994:

Information technology — Digital compression and coding of continuous-tone still images: Requirements and guidelines. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 182 pp. LCCN ???? CHF 224. URL http://www.

iso.ch/cate/d18902.html.

ISO:1994:IIIa

[ISO94d] ISO. ISO/IEC 10967-1 (1994-12):

Information technology — Language independent arithmetic — Part 1: Integer and floating point arithmetic. International Electrotechnical Commission, Geneva, Switzerland, December 15, 1994. viii + 92 pp. CHF 176, US$136.00.

URL http://standards.iso.
org/ittf/PubliclyAvailableStandards/
c018939_ISO_IEC_10967-1_1994(
E).zip;   http://www.iec.ch/
cgi-bin/procgi.pl/www/iecwww.
p?wwwlang=E&wwwprog=cat-det.
p&wartnum=019729; http://www.
iso.ch/cate/d18939.html.

ISO:1994:IIIb

[ISO94e] ISO. ISO/IEC 11430:1994: Information technology — Programming languages — Generic package of elementary functions for ADA. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 50 pp. LCCN ???? CHF 136; US$92.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+11430%3A1994;
http://www.iso.ch/cate/d19370.

html. Available in English only.

ISO:1994:IIIc

[ISO94f] ISO. ISO/IEC 11729:1994: Information technology — Programming languages — Generic package of primitive functions for Ada. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 19 pp. LCCN ???? CHF 86; US$58.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+11729%3A1994;
http://www.iso.ch/cate/d20879.

html. Available in English only.

ISO:1994:IIId

[ISO94g] ISO. ISO/IEC 11730:1994: Information technology — Programming languages — Form Interface Management System (FIMS). International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 494 pp. LCCN ???? CHF 334; US$225.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+11730%3A1994;
http://www.iso.ch/cate/d20881.
html.

ISO:1994:IIIe

[ISO94h] ISO. ISO/IEC 1539-2:1994: Information technology — Programming languages — FORTRAN — Part 2: Varying length character strings. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 66 pp. LCCN ????

CHF 152. URL http://www.
ansi.org/;   http://www.iso.
ch/cate/d6129.html. Available

in English only.

ISO:1994:IIIl

[ISO94i] ISO. ISO/IEC 2022:1994: Information technology — Character code structure and extension techniques. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 47 pp. LCCN ???? CHF 136.

URL http://www.iso.ch/cate/

d22747.html. Supercedes ANSI X3.41.

ISO:1994:IIIh

[ISO94j] ISO. ISO/IEC 2382-23:1994: Information technology — Vocabulary — Part 23: Text processing. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 24 pp. LCCN ???? CHF 98. URL http:/

/www.iso.ch/cate/d7267.html.

ISO:1994:IIIi

[ISO94k] ISO. ISO/IEC 2382-27:1994: Information technology — Vocabulary — Part 27: Office automation. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ????

16 pp. LCCN ???? CHF 74.

                                        URL http://www.iso.ch/cate/
                                        d7271.html.

ISO:1994:IIIf

[ISO94l] ISO. ISO/IEC 7942-1:1994: Information technology — Computer graphics and image processing — Graphical Kernel System (GKS)

— Part 1: Functional description. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 164 pp. LCCN ???? CHF

                                        212. URL http://www.iso.ch/
                                        cate/d14915.html.

ISO:1994:IICa

[ISO94m] ISO. ISO/IEC 9541-1:1991/Cor 2:1994. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 3 pp. LCCN ???? URL http://

                                        www.iso.ch/cate/d23758.html.

ISO:1994:IIIj

[ISO94n] ISO. ISO/IEC 9541-3:1994: Information technology — Font information interchange — Part 3: Glyph shape representation. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 60 pp. LCCN ???? CHF 144. URL http://www.

                                        iso.ch/cate/d17280.html.

ISO:1994:IICb

[ISO94o] ISO. ISO/IEC 9593-1:1990/Cor 2:1994. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 2 pp. LCCN ???? URL http://

                                        www.iso.ch/cate/d22878.html.

ISO:1994:IIAb

[ISO94p] ISO. ISO/IEC 9593-3:1990/Amd 1:1994: Incorporation of PHIGS PLUS. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 333 pp. LCCN ???? CHF

288. URL http://www.iso.ch/
cate/d21312.html. Available in

English only.

ISO:1994:IICc

[ISO94q] ISO. ISO/IEC 9593-3:1990/Cor 2:1994. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 3 pp. LCCN ???? URL http://

www.iso.ch/cate/d22879.html.

ISO:1994:IIAc

[ISO94r] ISO. ISO/IEC 9593-4:1991/Amd 1:1994. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 157 pp. LCCN ???? CHF

212. URL http://www.iso.ch/
cate/d21152.html. Available in

English only.

ISO:1994:IICd

[ISO94s] ISO. ISO/IEC 9593-4:1991/Cor 1:1994. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 4 pp. LCCN ???? URL http://

www.iso.ch/cate/d22880.html.

Available in English only.

ISO:1994:IIIg

[ISO94t] ISO. ISO/IEC 9638-3:1994: Information technology — Computer graphics — Interfacing techniques for dialogues with graphical devices (CGI) — Language bindings — Part 3: Ada. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 361 pp. LCCN ???? CHF

306. URL http://www.iso.ch/
cate/d17461.html. Available in

English only.

ISO:1995:ICR

[ISO95a] ISO. ISO 4217:1995: Codes for the representation of currencies and funds. International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 32 pp. LCCN ???? CHF 110.

URL http://www.iso.ch/cate/
d23132.html.

ISO:1995:IIId

[ISO95b] ISO. ISO/IEC 10180:1995: Information technology — Processing languages — Standard Page Description Language (SPDL). International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 464 pp. LCCN ???? CHF 320. URL http://www.

iso.ch/cate/d18197.html.

ISO:1995:IIIl

[ISO95c] ISO. ISO/IEC 10918-2:1995: Information technology — Digital compression and coding of continuous-tone still images: Compliance testing. International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 181 pp. LCCN ???? ????

URL http://www.iso.ch/cate/
d20689.html.

ISO:1995:IIIb

[ISO95d] ISO. ISO/IEC 12088-4:1995:

Information technology — Computer graphics and image processing — Image processing and interchange — Application program interface language bindings — Part 4: C. International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 613 pp. LCCN ???? CHF

348. URL http://www.iso.ch/
cate/d23287.html.

ISO:1995:IIIa

[ISO95e] ISO. ISO/IEC 12227:1995: Information technology — Programming languages — SQL/ Ada Module Description Language (SAMeDL). International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 120 pp. LCCN ????

CHF 188. URL http://www.
ansi.org/;   http://www.iso.
ch/cate/d2345.html.

ISO:1995:IIIk

[ISO95f] ISO. ISO/IEC 12227:1995: Information technology — Programming languages — SQL/ Ada Module Description Language (SAMeDL). International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 120 pp. LCCN ???? CHF 188; US$128.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+12227%3A1995;
http://www.iso.ch/cate/d2345.

html. Available in English only.

ISO:1995:IIIe

[ISO95g] ISO. ISO/IEC 13211-1:1995: Information technology — Programming languages — Prolog — Part 1: General core. International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 199 pp. LCCN ???? CHF 224; US$152.00.

                                        URL http://webstore.ansi.
                                        org/ansidocstore/product.asp?
                                        sku=ISO%2FIEC+13211%2D1%3A1995;
                                        http://www.iso.ch/cate/d21413.
                                        html.

ISO:1995:IIIj

[ISO95h] ISO. ISO/IEC 2382-24:1995: Information technology — Vocabulary — Part 24: Computerintegrated manufacturing. International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 13 pp. LCCN ???? CHF 68. URL http://www.

                                        iso.ch/cate/d7268.html.

ISO:1995:IIIc

[ISO95i] ISO. ISO/IEC 2382-28:1995: Information technology — Vocabulary — Part 28: Artificial intelligence — Basic concepts and expert systems. International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 25 pp. LCCN ???? CHF 98.

                                        URL http://www.iso.ch/cate/
                                        d7272.html.

ISO:1995:IIIi

[ISO95j] ISO. ISO/IEC 2382-9:1995: Information technology — Vocabulary — Part 9: Data communication. International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 43 pp. LCCN ???? CHF 128.

URL http://www.iso.ch/cate/
d20929.html.

ISO:1995:IIIf

[ISO95k] ISO. ISO/IEC 8651-4:1995: Information technology — Computer graphics — Graphical Kernel System (GKS) language bindings — Part 4: C. International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 339 pp. LCCN ???? CHF

288. URL http://www.iso.ch/
cate/d23070.html.

ISO:1995:IIIg

[ISO95l] ISO. ISO/IEC 8652:1995: Information technology — Programming languages — Ada. International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 511 pp. LCCN ???? CHF 334.

URL http://www.iso.ch/cate/

d22983.html. Available in English only.

ISO:1995:IIIh

[ISO95m] ISO. ISO/IEC 9075-3:1995: Information technology — Database languages — SQL — Part 3: Call-Level Interface (SQL/CLI). International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 401 pp. LCCN ???? CHF 306. URL http://www.

iso.ch/cate/d30609.html.

ISO:1995:IICa

[ISO95n] ISO. ISO/IEC 9541-1:1991/Cor 3:1995. International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 1 pp. LCCN ???? URL http://

                                        www.iso.ch/cate/d25788.html.

ISO:1995:IICb

[ISO95o] ISO. ISO/IEC 9541-2:1991/Cor 2:1995. International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 6 pp. LCCN ???? URL http://

                                        www.iso.ch/cate/d25789.html.

ISO:1995:IIA

[ISO95p] ISO. ISO/IEC 9593-1:1990/Amd 1:1995. International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 149 pp. LCCN ???? CHF

                                        200. URL http://www.iso.ch/
                                        cate/d1911.html.

ISO:1995:ITM

[ISO95q] ISO. ISO/TR 10562:1995: Manipulating industrial robots — Intermediate Code for Robots (ICR). International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 120 pp. LCCN ???? CHF 188.

                                        URL http://www.iso.ch/cate/
                                        d18647.html.

ISO:1995:PLC

[ISO95r] ISO. Programming languages — C — Amendment 1: C integrity (ISO/IEC 9899:1990/AMD 1:1995); English version EN 29899:1993/A1:1996 (foreign standard). International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? ???? pp. LCCN ????

                                        US$121. URL http://www.ansi.
                                        org/.

ISO:1996:IID

[ISO96a] ISO. ISO 2709:1996: Information and documentation — Format for Information Exchange. International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 6 pp. LCCN ???? CHF 44. URL http://www.

iso.ch/cate/d7675.html.

ISO:1996:IC

[ISO96b] ISO. ISO 8879:1986/Cor 1:1996.

International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 2 pp. LCCN ???? URL http://

www.iso.ch/cate/d28557.html.

Available in English only.

ISO:1996:IIIb

[ISO96c] ISO. ISO/IEC 10179:1996: Information technology — Processing languages — Document Style Semantics and Specification Language (DSSSL). International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 292 pp. LCCN ???? CHF

270. URL http://www.iso.ch/
cate/d18196.html.

ISO:1996:IIIc

[ISO96d] ISO. ISO/IEC 10514-1:1996:

Information technology — Programming languages — Part 1: Modula-2, Base Language. International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 707 pp. LCCN ???? CHF 360; US$245.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+10514%2D1%3A1996;

http://www.iso.ch/cate/d18583.
html.

ISO:1996:IIId

[ISO96e] ISO. ISO/IEC 11404:1996: Information technology — Programming languages, their environments and system software interfaces — Language-independent datatypes. International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 96 pp. LCCN ???? CHF 176; US$120.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+11404%3A1996;
http://www.iso.ch/cate/d19346.
html.

ISO:1996:IIIe

[ISO96f] ISO. ISO/IEC 13817-1:1996: Information technology — Programming languages, their environments and system software interfaces — Vienna Development Method — Specification Language — Part 1: Base language. International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 399 pp. LCCN ???? CHF 306; US$208.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+13817%2D1%3A1996;
http://www.iso.ch/cate/d22988.
html.

ISO:1996:IIIf

[ISO96g] ISO. ISO/IEC 13886:1996: Information technology — Language-Independent Procedure Calling (LIPC). International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 69 pp. LCCN ???? CHF 152.

URL http://www.iso.ch/cate/
d23183.html.

ISO:1996:IIIg

[ISO96h] ISO. ISO/IEC 14360:1996: Information technology — Open Systems Interconnection (OSI) abstract data manipulation — Application Program Interface (API) (Language independent). International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 77 pp. LCCN ???? CHF 164. URL http://www.

iso.ch/cate/d24130.html.

ISO:1996:IIIh

[ISO96i] ISO. ISO/IEC 14361:1996: Information technology — MHSbased electronic messaging — Application Program Interface (API) (Language independent). International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 190 pp. LCCN ???? CHF 224. URL http://www.

iso.ch/cate/d24131.html.

ISO:1996:IIIi

[ISO96j] ISO. ISO/IEC 14362:1996: Information technology — Test methods for measuring conformance to Open Systems Interconnection (OSI) abstract data manipulation — Application Program Interface (API) (Language independent). International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 80 pp. LCCN ???? CHF 164.

URL http://www.iso.ch/cate/
d24132.html.

ISO:1996:IIIj

[ISO96k] ISO. ISO/IEC 14363:1996: Information technology — Test methods for measuring conformance to MHS-based electronic messaging — Application Program Interface (API) (Language independent). International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 583 pp. LCCN ???? CHF

                                        348. URL http://www.iso.ch/
                                        cate/d24133.html.

ISO:1996:IIIk

[ISO96l] ISO. ISO/IEC 14364:1996: Information technology — Open Systems Interconnection (OSI) abstract data manipulation C language interfaces — Binding for Application Program Interface (API). International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 63 pp. LCCN ???? CHF 152.

                                        URL http://www.iso.ch/cate/
                                        d24134.html.

ISO:1996:IIIl

[ISO96m] ISO. ISO/IEC 14365:1996: Information technology — MHS-based electronic messaging C language interfaces — Binding for Application Program Interface (API). International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 52 pp. LCCN ???? CHF 144. URL http://www.

                                        iso.ch/cate/d24135.html.

ISO:1996:IIIm

[ISO96n] ISO. ISO/IEC 14366:1996: Information technology — Test methods for measuring conformance to Open Systems Interconnection (OSI) abstract data manipulation C language interfaces — Binding for Application Program Interface (API). International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 62 pp. LCCN ???? CHF 152.

URL http://www.iso.ch/cate/
d24136.html.

ISO:1996:IIIn

[ISO96o] ISO. ISO/IEC 14367:1996: Information technology — Test methods for measuring conformance to MHS based electronic messaging C language interfaces — Binding for Application Program Interface (API). International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 55 pp. LCCN ???? CHF 144.

URL http://www.iso.ch/cate/
d24137.html.

ISO:1996:IIIo

[ISO96p] ISO. ISO/IEC 14392:1996: Information technology — Directory services — Application Program Interface (API) (Language independent). International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 154 pp. LCCN ???? CHF

212. URL http://www.iso.ch/
cate/d24192.html.

ISO:1996:IIIp

[ISO96q] ISO. ISO/IEC 14393:1996: Information technology — Test methods for measuring conformance to directory services — Application Program Interface (API) (Language independent). International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 229 pp. LCCN ???? CHF

                                        236. URL http://www.iso.ch/
                                        cate/d24193.html.

ISO:1996:IIIq

[ISO96r] ISO. ISO/IEC 14394:1996: Information technology — Directory services C language interfaces — Binding for Application Program Interface (API). International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 91 pp. LCCN ???? CHF

                                        176. URL http://www.iso.ch/
                                        cate/d24194.html.

ISO:1996:IIIr

[ISO96s] ISO. ISO/IEC 14395:1996: Information technology — Test methods for measuring conformance to directory services C language interfaces — Binding for Application Program Interface (API). International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 61 pp. LCCN ???? CHF 152. URL http://www.

                                        iso.ch/cate/d24195.html.

ISO:1996:IIIs

[ISO96t] ISO. ISO/IEC 14977:1996: Information technology — Syntactic metalanguage — Extended BNF. International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 12 pp. LCCN ???? CHF 62.

                                        URL http://www.iso.ch/cate/
                                        d26153.html.

ISO:1996:IIIu

[ISO96u] ISO. ISO/IEC 2382-13:1996: Information technology — Vocabulary — Part 13: Computer graphics. International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 42 pp. LCCN ???? CHF 128.

URL http://www.iso.ch/cate/
d7253.html.

ISO:1996:IIIv

[ISO96v] ISO. ISO/IEC 2382-16:1996: Information technology — Vocabulary — Part 16: Information theory. International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 19 pp. LCCN ???? CHF 86.

URL http://www.iso.ch/cate/
d7259.html.

ISO:1996:IIC

[ISO96w] ISO. ISO/IEC 9075:1992/Cor 1:1996. International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 80 pp. LCCN ???? URL http://

www.iso.ch/cate/d27302.html.

Available in English only.

ISO:1996:IIIa

[ISO96x] ISO. ISO/IEC 9579-3:1996: Information technology — Open Systems Interconnection — Remote Database Access — Part 3: SQL specialization Protocol Implementation Conformance Statement (PICS) proforma. International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? ???? pp. LCCN ????

US$125. URL http://www.ansi.
org/.

ISO:1996:IIIt

[ISO96y] ISO. ISO/IEC 9945-1:1996: Information technology — Portable Operating System Interface (POSIX)

— Part 1: System Application Program Interface (API) [C Language]. International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 743 pp. LCCN ???? CHF 360.

URL http://www.iso.ch/cate/
d24426.html.

ISO:1996:IIT

[ISO96z] ISO. ISO/IEC TR 11735:1996: Information technology — Extensions for real-time Ada. International Organization for Standardization, Geneva, Switzerland, 1996. ISBN ???? 337 pp. LCCN ???? CHF 288. URL http://www.

iso.ch/cate/d22986.html.

ISO:1997:IIA

[ISO97a] ISO. ISO 13281:1997: Industrial automation systems — Manufacturing Automation Programming Environment (MAPLE) — Functional architecture. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 30 pp. LCCN ???? CHF

110. URL http://www.iso.ch/
cate/d21513.html.

ISO:1997:ISDa

[ISO97b] ISO. ISO 14961:1997: Space data and information transfer systems — Parameter value language specification. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 50 pp. LCCN ???? CHF 136.

URL http://www.iso.ch/cate/
d26095.html.

ISO:1997:ISDb

[ISO97c] ISO. ISO 14962:1997: Space data and information transfer systems — ASCII encoded English. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 4 pp. LCCN ???? CHF 38. URL http://www.

iso.ch/cate/d26096.html.

ISO:1997:ICR

[ISO97d] ISO. ISO 3166-1:1997: Codes for the representation of names of countries and their subdivisions — Part 1: Country codes. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 58 pp. LCCN ???? CHF 144.00. URL http://

www.din.de/gremien/nas/nabd/
iso3166ma/codlstp1/index.html;
http://www.iso.ch/cate/d24591.
html.

ISO:1997:IID

[ISO97e] ISO. ISO 4:1997: Information and documentation — Rules for the abbreviation of title words and titles of publications. International Organization for Standardization, Geneva, Switzerland, third edition, December 1997. 11

pp. URL https://www.iso.org/
standard/3569.html.

ISO:1997:IIIb

[ISO97f] ISO. ISO/IEC 10744:1997: Information technology — Hypermedia/Time-based Structuring Language (HyTime). International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 468 pp. LCCN ???? CHF

320. URL http://www.iso.ch/
cate/d29303.html.

ISO:1997:IIIm

[ISO97g] ISO. ISO/IEC 10918-3:1997: Information technology — Digital compression and coding of continuous-tone still images: Extensions. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 80 pp. LCCN ???? CHF 164.

URL http://www.iso.ch/cate/
d25037.html.

ISO:1997:IIId

[ISO97h] ISO. ISO/IEC 13816:1997: Information technology — Programming languages, their environments and system software interfaces — Programming language ISLISP. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 126 pp. LCCN ???? CHF 200; US$136.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+13816%3A1997;
http://www.iso.ch/cate/d22987.
html.

ISO:1997:IIIc

[ISO97i] ISO. ISO/IEC 14568:1997: Information technology — DXL: Diagram eXchange Language for treestructured charts. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 13 pp. LCCN ???? CHF

68. URL http://www.iso.ch/
cate/d24813.html.

ISO:1997:IIIe

[ISO97j] ISO. ISO/IEC 15145:1997: Information technology — Programming languages — FORTH. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 210 pp. LCCN ???? CHF 224; US$152.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+15145%3A1997;
http://www.iso.ch/cate/d26479.
html.

ISO:1997:IIIa

[ISO97k] ISO. ISO/IEC 1539-1:1997: Information technology — Programming languages — Fortran — Part 1: Base language. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? ???? pp. LCCN ???? US$195.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+1539%2D1%3A1997.

ISO:1997:IIIg

[ISO97l] ISO. ISO/IEC 1539-1:1997: Information technology — Programming languages — Fortran — Part 1: Base language. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 346 pp. LCCN ???? CHF

288. URL http://www.ansi.org/
;   http://www.fortran.com/
fortran/Books/f95std.html;
http://www.iso.ch/cate/d26933.

html. ISO/IEC JTC 1/SC 22/WG 5. This is the Fortran 95 Standard. Available in English only.

ISO:1997:IIIh

[ISO97m] ISO. ISO/IEC 2382-14:1997: Information technology — Vocabulary — Part 14: Reliability, maintainability and availability. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 18 pp. LCCN ???? CHF 80. URL http://www.

                                        iso.ch/cate/d7255.html.

ISO:1997:IIIi

[ISO97n] ISO. ISO/IEC 2382-31:1997: Information technology — Vocabulary — Part 31: Artificial intelligence — Machine learning. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 16 pp. LCCN ???? CHF 74. URL http://www.

                                        iso.ch/cate/d21845.html.

ISO:1997:IIIf

[ISO97o] ISO. ISO/IEC 7942-2:1997: Information technology — Computer graphics and image processing — Graphical Kernel System (GKS)

— Part 2: NDC metafile. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 54 pp. LCCN ???? CHF 144. URL http://www.

                                        iso.ch/cate/d24746.html.

ISO:1997:IIIj

[ISO97p] ISO. ISO/IEC 9592-1:1997: Information technology — Computer graphics and image processing — Programmer’s Hierarchical Interactive Graphics System (PHIGS)

— Part 1: Functional description. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 510 pp. LCCN ???? CHF

334. URL http://www.iso.ch/
cate/d29417.html. Available in

English only.

ISO:1997:IIIk

[ISO97q] ISO. ISO/IEC 9592-2:1997: Information technology — Computer graphics and image processing — Programmer’s Hierarchical Interactive Graphics System (PHIGS)

— Part 2: Archive file format. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 6 pp. LCCN ????

CHF 44. URL http://www.iso.
ch/cate/d29419.html. Available

in English only.

ISO:1997:IIIl

[ISO97r] ISO. ISO/IEC 9592-3:1997: Information technology — Computer graphics and image processing — Programmer’s Hierarchical Interactive Graphics System (PHIGS) — Part 3: Specification for clear-text encoding of archive file. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 43 pp. LCCN ???? CHF 128.

URL http://www.iso.ch/cate/

d29421.html. Available in English only.

ISO:1997:ITI

[ISO97s] ISO. ISO/TR 10303-12:1997: Industrial automation systems and integration — Product data representation and exchange — Part 12: Description methods: The EXPRESS-I language reference manual. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 111 pp. LCCN ???? CHF 188.

URL http://www.iso.ch/cate/
d25071.html.

ISO:1998:IGT

[ISO98a] ISO. ISO 12639:1998: Graphic technology — Prepress digital data exchange — Tag image file format for image technology (TIFF/IT). International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 43 pp. LCCN ???? CHF 128.

URL http://www.iso.ch/cate/
d2181.html.

ISO:1998:ICR

[ISO98b] ISO. ISO 3166-2:1998 Codes for the representation of names of countries and their subdivisions — Part 2: Country subdivision code. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 105 pp. LCCN ???? CHF 188.00.

URL http://www.iso.ch/cate/
d8349.html.

ISO:1998:IID

[ISO98c] ISO. ISO 3297:1998: Information and documentation — International Standard Serial Number (ISSN). International Organization for Standardization, Geneva, Switzerland, November 12, 1998. ISBN ???? 7 pp. LCCN ???? CHF

50.00. URL http://www.iso.ch/
cate/d23400.html;   http://

www.issn.org/brochure/issnbroc.
htm.

ISO:1998:IIIc

[ISO98d] ISO. ISO/IEC 10514-2:1998:

Information technology — Programming languages — Part 2: Generics Modula-2. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 45 pp. LCCN ???? CHF 128; US$88.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+10514%2D2%3A1998;
http://www.iso.ch/cate/d20792.
html.

ISO:1998:IIId

[ISO98e] ISO. ISO/IEC 10514-3:1998: Information technology — Programming languages — Part 3: Object Oriented Modula-2. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 48 pp. LCCN ???? CHF 136; US$92.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+10514%2D3%3A1998;
http://www.iso.ch/cate/d20793.
html.

ISO:1998:IIIe

[ISO98f] ISO. ISO/IEC 13719-1:1998: Information technology — Portable Common Tool Environment (PCTE)

— Part 1: Abstract specification. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 460 pp. LCCN ???? CHF 320.

URL http://www.iso.ch/cate/
d30520.html.

ISO:1998:IIIf

[ISO98g] ISO. ISO/IEC 13719-2:1998: Information technology — Portable Common Tool Environment (PCTE)

— Part 2: C programming language binding. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 145 pp. LCCN ???? CHF

                                        200. URL http://www.iso.ch/
                                        cate/d30521.html.

ISO:1998:IIIg

[ISO98h] ISO. ISO/IEC 13719-3:1998: Information technology — Portable Common Tool Environment (PCTE)

— Part 3: Ada programming language binding. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 161 pp. LCCN ???? CHF

                                        212. URL http://www.iso.ch/
                                        cate/d30522.html. Available in

English only.

ISO:1998:IIIh

[ISO98i] ISO. ISO/IEC 13719-4:1998: Information technology — Portable Common Tool Environment (PCTE)

— Part 4: IDL binding (Interface Definition Language). International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 125 pp. LCCN ???? CHF 200. URL http://www.

                                        iso.ch/cate/d28940.html.

ISO:1998:IIIi

[ISO98j] ISO. ISO/IEC 13813:1998: Information technology — Programming languages — Generic packages of real and complex type declarations and basic operations for Ada (including vector and matrix types). International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 62 pp. LCCN ???? CHF 152; US$105.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+13813%3A1998;
http://www.iso.ch/cate/d28992.

html. Available in English only.

ISO:1998:IIIj

[ISO98k] ISO. ISO/IEC 13814:1998: Information technology — Programming languages — Generic package of complex elementary functions for Ada. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 36 pp. LCCN ???? CHF 122; US$84.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+13814%3A1998;
http://www.iso.ch/cate/d22985.

html. Available in English only.

ISO:1998:IIIk

[ISO98l] ISO. ISO/IEC 14772-1:1998: Information technology — Computer graphics and image processing — The Virtual Reality Modeling Language — Part 1: Functional specification and UTF-8 encoding. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 236 pp. LCCN ???? CHF 252. URL http://www.

iso.ch/cate/d25508.html.

ISO:1998:IIP

[ISO98m] ISO. ISO/IEC 14882:1998: Programming languages — C++. International Organization for Standardization, Geneva, Switzerland, September 1, 1998. ISBN ???? 732 pp. LCCN ???? CHF 351, US$18 (electronic), US$252 (print); US$245.00. URL http://

                                        webstore.ansi.org/ansidocstore/
                                        product.asp?sku=ISO%2FIEC+14882%
                                        2D1998;   http://webstore.ansi.
                                        org/ansidocstore/product.asp?
                                        sku=ISO%2FIEC+14882%3A1998;
                                         http://www.iso.ch/cate/d25845.
                                        html; https://webstore.ansi.

org/. Available in electronic form for online purchase at http://

                                        webstore.ansi.org/ and http:
                                        //www.cssinfo.com/.

ISO:1998:IIIl

[ISO98n] ISO. ISO/IEC 16262:1998: Information technology — ECMAScript language specification. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 98 pp. LCCN ???? CHF 176. URL http://www.

                                        iso.ch/cate/d29696.html.

ISO:1998:IIIa

[ISO98o] ISO. ISO/IEC 2382-8:1998: Information technology — Vocabulary — Part 8: Security. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 38 pp. LCCN ???? CHF 122. URL http://www.

                                        iso.ch/cate/d7243.html.

ISO:1998:IIIm

[ISO98p] ISO. ISO/IEC 7942-4:1998: Information technology — Computer graphics and image processing — Graphical Kernel System (GKS) — Part 4: Picture part archive. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 10 pp. LCCN ???? CHF 56.

URL http://www.iso.ch/cate/
d24748.html.

ISO:1998:IIIp

[ISO98q] ISO. ISO/IEC 8859-10:1998:

Information technology — 8-bit single-byte coded graphic character sets — Part 4: Latin alphabet No. 6. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 10 pp. LCCN ???? 59 Sfr (print), 59 Sfr (PDF).

URL http://www.iso.org/iso/
en/CatalogueDetailPage.CatalogueDetail?
CSNUMBER=28254.

ISO:1998:IIIn

[ISO98r] ISO. ISO/IEC 8859-1:1998:

Information technology — 8-bit single-byte coded graphic character sets — Part 1: Latin alphabet No. 1. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 10 pp. LCCN ???? 59

Sfr. URL http://www.iso.org/
iso/en/CatalogueDetailPage.
CatalogueDetail?CSNUMBER=28245.

ISO:1998:IIIq

[ISO98s] ISO. ISO/IEC 8859-13:1998:

Information technology — 8-bit single-byte coded graphic character sets — Part 13: Latin alphabet No. 7. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 10 pp. LCCN ????

59 Sfr (print), 59 Sfr (PDF).

                                        URL http://www.iso.org/iso/
                                        en/CatalogueDetailPage.CatalogueDetail?
                                        CSNUMBER=28265.

ISO:1998:IIIr

[ISO98t] ISO. ISO/IEC 8859-14:1998: Information technology — 8-bit single-byte coded graphic character sets — Part 14: Latin alphabet No. 8 (Celtic). International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 9 pp. LCCN ???? 59 Sfr (print), 59 Sfr (PDF).

                                        URL http://www.iso.org/iso/
                                        en/CatalogueDetailPage.CatalogueDetail?
                                        CSNUMBER=29630.

ISO:1998:IIIo

[ISO98u] ISO. ISO/IEC 8859-4:1998: Information technology — 8-bit single-byte coded graphic character sets — Part 4: Latin alphabet No. 4. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 10 pp. LCCN ???? 59 Sfr (print), 59 Sfr (PDF).

                                        URL http://www.iso.org/iso/
                                        en/CatalogueDetailPage.CatalogueDetail?
                                        CSNUMBER=28248.

ISO:1998:IIC

[ISO98v] ISO. ISO/IEC 9496:1998: CCITT high level language (CHILL). International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 240 pp. LCCN ???? CHF 236. URL http://www.

                                        iso.ch/cate/d30537.html.

ISO:1998:IIAa

[ISO98w] ISO. ISO/IEC 9541-1:1991/Amd 2:1998: Minor enhancements to the architecture to address font technology advances. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 4 pp. LCCN ???? CHF

13. URL http://www.iso.ch/
cate/d24741.html. Available in

English only.

ISO:1998:IIIb

[ISO98x] ISO. ISO/IEC 9579-2:1998: Information technology — Open Systems Interconnection — Remote Database Access — Part 2: SQL Specialization. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? ???? pp. LCCN ????

US$105. URL http://www.ansi.
org/.

ISO:1998:IIAb

[ISO98y] ISO. ISO/IEC 9593-4:1991/Amd 2:1998: Incorporation of PHIGS amendments. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 355 pp. LCCN ???? CHF

288. URL http://www.iso.ch.

ISO:1998:IITa

[ISO98z] ISO. ISO/IEC TR 10176:1998:

Information technology — Guidelines for the preparation of programming language standards. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 33 pp. LCCN ???? CHF 116. URL http://www.

iso.ch/cate/d28707.html.

ISO:1998:IITb

[ISO98-27] ISO. ISO/IEC TR 11017:1998: Information technology — Framework for internationalization. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 47 pp. LCCN ???? CHF 136. URL http://www.

                                        iso.ch/cate/d19001.html.

ISO:1998:IITc

[ISO98-28] ISO. ISO/IEC TR 15580:1998: Information technology — Programming languages — Fortran — Floating-point exception handling. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 27 pp. LCCN ???? CHF 104; US$72.00.

                                        URL http://webstore.ansi.
                                        org/ansidocstore/product.asp?
                                        sku=ISO%2FIEC+TR+15580%3A1998;
                                        http://www.iso.ch/cate/d28230.

html. Available in English only.

ISO:1998:IITd

[ISO98-29] ISO. ISO/IEC TR 15581:1998: Information technology — Programming languages — Fortran — Enhanced data type facilities. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 13 pp. LCCN ???? CHF 68; US$46.00.

                                        URL http://webstore.ansi.
                                        org/ansidocstore/product.asp?
                                        sku=ISO%2FIEC+TR+15581%3A1998;
                                        http://www.iso.ch/cate/d28231.

html. Available in English only.

ISO:1999:IID

[ISO99a] ISO. ISO 233-3:1999: Information and documentation — Transliteration of Arabic characters into Latin characters — Part 3: Persian language — Simplified transliteration. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 11 pp. LCCN ???? CHF

62. URL http://www.iso.ch/
cate/d2398.html.

ISO:1999:ICR

[ISO99b] ISO. ISO 3166-3:1999 Codes for the representation of names of countries and their subdivisions — Part 3: Code for formerly used names of countries. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 11 pp. LCCN ???? CHF

62.00. URL http://www.iso.ch/
cate/d2130.html.

ISO:1999:IC

[ISO99c] ISO. ISO 8879:1986/Cor 2:1999.

International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 18 pp. LCCN ???? URL http://

www.iso.ch/cate/d30235.html.

Available in English only.

ISO:1999:IIA

[ISO99d] ISO. ISO/IEC 10918-3:1997/ Amd 1:1999: Provisions to allow registration of new compression types and versions in the SPIFF header. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 1 pp. LCCN ???? CHF 13.

URL http://www.iso.ch/cate/
d30961.html.

ISO:1999:IIIw

[ISO99e] ISO. ISO/IEC 10918-4:1999: Information technology — Digital compression and coding of continuous-tone still images: Registration of JPEG profiles, SPIFF profiles, SPIFF tags, SPIFF colour spaces, APPn markers, SPIFF compression types and Registration Authorities (REGAUT). International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 29 pp. LCCN ???? CHF 104.

URL http://www.iso.ch/cate/

d25431.html. Available in English only.

ISO:1999:IIIa

[ISO99f] ISO. ISO/IEC 11756:1999: Information technology — Programming languages — M. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 115 pp. LCCN ???? CHF 188; US$128.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+11756%3A1999;
http://www.iso.ch/cate/d19978.
html.

ISO:1999:IIIu

[ISO99g] ISO. ISO/IEC 13249-3:1999: Information technology — Database languages — SQL Multimedia and Application Packages — Part 3: Spatial. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 332 pp. LCCN ???? CHF

288. URL http://www.iso.ch/
cate/d28054.html.

ISO:1999:IIIx

[ISO99h] ISO. ISO/IEC 14496-1:1999: Information technology — Coding of audio-visual objects — Part 1: Systems. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 281 pp. LCCN ???? ????

URL http://www.iso.ch/cate/

d24462.html. Available in English only.

ISO:1999:IIIy

[ISO99i] ISO. ISO/IEC 14496-2:1999: Information technology — Coding of audio-visual objects — Part 2: Visual. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 330 pp. LCCN ???? CHF

288. URL http://www.iso.ch/
cate/d25034.html. Available in

English only.

ISO:1999:IIIz

[ISO99j] ISO. ISO/IEC 14496-3:1999: Information technology — Coding of audio-visual objects — Part 3: Audio. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 644 pp. LCCN ???? CHF

348. URL http://www.iso.ch/
cate/d25035.html. Available in

English only.

ISO:1999:IIIaa

[ISO99k] ISO. ISO/IEC 14496-5:2000: Information technology — Coding of audio-visual objects — Part 5: Reference software. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 7 pp. LCCN ???? ????

URL http://www.iso.ch/cate/

d30554.html. Available in English only.

ISO:1999:IIIab

[ISO99l] ISO. ISO/IEC 14496-6:1999: Information technology — Coding of audio-visual objects — Part 6: Delivery Multimedia Integration Framework (DMIF). International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 101 pp. LCCN ???? CHF 188. URL http://www.

iso.ch/cate/d30556.html.

ISO:1999:IIIb

[ISO99m] ISO. ISO/IEC 14519:1999: Information technology — POSIX Ada Language Interfaces — Binding for System Application Program Interface (API) — Realtime Extensions. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 529 pp. LCCN ???? CHF

334. URL http://www.iso.ch/
cate/d29522.html. Available in

English only.

ISO:1999:IIIc

[ISO99n] ISO. ISO/IEC 14750:1999: Information technology — Open Distributed Processing — Interface Definition Language. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 31 pp. LCCN ???? CHF

110. URL http://www.iso.ch/
cate/d25486.html.

ISO:1999:IIId

[ISO99o] ISO. ISO/IEC 15291:1999: Information technology — Programming languages — Ada Semantic Interface Specification (ASIS). International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 283 pp. LCCN ???? CHF 270; US$185.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+15291%3A1999;
http://www.iso.ch/cate/d27169.

html. Available in English only.

ISO:1999:IIIe

[ISO99p] ISO. ISO/IEC 1539-3:1999: Information technology — Programming languages — Fortran — Part 3: Conditional compilation. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 21 pp. LCCN ???? CHF 92; US$62.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+1539%2D3%3A1999;
http://www.iso.ch/cate/d29926.

html. Available in English only.

ISO:1999:IIIf

[ISO99q] ISO. ISO/IEC 15851:1999: Information technology — Communication protocol — Open MUMPS Interconnect. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 25 pp. LCCN ???? CHF 98.

URL http://www.iso.ch/cate/
d29269.html.

ISO:1999:IIIh

[ISO99r] ISO. ISO/IEC 15852:1999: Information technology — Programming languages — M Windowing API. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 85 pp. LCCN ???? CHF 176; US$120.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+15852%3A1999;
http://www.iso.ch/cate/d29270.
html.

ISO:1999:IIIac

[ISO99s] ISO. ISO/IEC 15897:1999: Information technology — Procedures for registration of cultural elements. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 39 pp. LCCN ???? CHF 122.

URL http://anubis.dkuug.dk/
cultreg/;   http://www.iso.
ch/cate/d29452.html. Available

in English only.

ISO:1999:IIIs

[ISO99t] ISO. ISO/IEC 18009:1999: Information technology — Programming languages — Ada: Conformity assessment of a language processor. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 25 pp. LCCN ???? CHF 98; US$68.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+18009%3A1999;
http://www.iso.ch/cate/d31051.

html. Available in English only.

ISO:1999:IIIl

[ISO99u] ISO. ISO/IEC 2382-15:1999:

Information technology — Vocabulary — Part 15: Programming languages. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 46 pp. LCCN ???? CHF 136; US$92.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+2382%2D15%3A1999;
http://www.iso.ch/cate/d7257.
html.

ISO:1999:IIIr

[ISO99v] ISO. ISO/IEC 2382-17:1999: Information technology — Vocabulary — Part 17: Databases. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 26 pp. LCCN ???? CHF 98. URL http://www.

iso.ch/cate/d30853.html.

ISO:1999:IIIg

[ISO99w] ISO. ISO/IEC 2382-18:1999: Information technology — Vocabulary — Part 18: Distributed data processing. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? ca. 20 pp. LCCN ????

CHF 86. URL http://www.iso.
ch/cate/d26734.html.

ISO:1999:IIIi

[ISO99x] ISO. ISO/IEC 2382-29:1999: Information technology — Vocabulary — Part 29: Artificial intelligence — Speech recognition and synthesis. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? ca. 21 pp. LCCN ???? CHF

                                        92. URL http://www.iso.ch/
                                        cate/d7273.html.

ISO:1999:IIIj

[ISO99y] ISO. ISO/IEC 2382-32:1999: Information technology — Vocabulary — Part 32: Electronic Mail. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? ca. 41 pp. LCCN ???? CHF

                                        128. URL http://www.iso.ch/
                                        cate/d21846.html.

ISO:1999:IIIm

[ISO99z] ISO. ISO/IEC 2382-34:1999: Information technology — Vocabulary — Part 34: Artificial intelligence — Neural networks. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? ca. 20 pp. LCCN ???? CHF 86. URL http://www.

                                        iso.ch/cate/d21848.html.

ISO:1999:IIIp

[ISO99-27] ISO. ISO/IEC 2382-4:1999: Information technology — Vocabulary — Part 4: Organization of data. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 24 pp. LCCN ???? CHF 98.

                                        URL http://www.iso.ch/cate/
                                        d27922.html.

ISO:1999:IIIq

[ISO99-28] ISO. ISO/IEC 2382-5:1999: Information technology — Vocabulary — Part 5: Representation of data. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 22 pp. LCCN ???? CHF 92.

URL http://www.iso.ch/cate/
d30851.html.

ISO:1999:IIIk

[ISO99-29] ISO. ISO/IEC 7816-7:1999: Identification cards — Integrated circuit(s) cards with contacts — Part 7: Interindustry commands for Structured Card Query Language (SCQL). International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 36 pp. LCCN ???? CHF 122.

URL http://www.iso.ch/cate/
d28869.html.

ISO:1999:IIIt

[ISO99-30] ISO. ISO/IEC 7942-3:1999 Information technology — Computer graphics and image processing — Graphical Kernel System (GKS)

— Part 3: Audit trail. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 38 pp. LCCN ???? CHF 122. URL http://www.

iso.ch/cate/d24747.html.

ISO:1999:IIIaj

[ISO99-31] ISO. ISO/IEC 8859-15:1999:

Information technology — 8-bit single-byte coded graphic character sets — Part 15: Latin alphabet No. 9. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 9 pp. LCCN ???? 59 Sfr (print), 44 Sfr (PDF).

URL http://www.iso.org/iso/
en/CatalogueDetailPage.CatalogueDetail?
CSNUMBER=29505.

ISO:1999:IIIad

[ISO99-32] ISO. ISO/IEC 8859-2:1999: Information technology — 8-bit single-byte coded graphic character sets — Part 2: Latin alphabet No. 2. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 10 pp. LCCN ???? 59 Sfr (print), 44 Sfr (PDF).

                                        URL http://www.iso.org/iso/
                                        en/CatalogueDetailPage.CatalogueDetail?
                                        CSNUMBER=28246.

ISO:1999:IIIae

[ISO99-33] ISO. ISO/IEC 8859-3:1999: Information technology — 8-bit single-byte coded graphic character sets — Part 3: Latin alphabet No. 3. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 10 pp. LCCN ???? 59 Sfr (print), 44 Sfr (PDF).

                                        URL http://www.iso.org/iso/
                                        en/CatalogueDetailPage.CatalogueDetail?
                                        CSNUMBER=28247.

ISO:1999:IIIaf

[ISO99-34] ISO. ISO/IEC 8859-5:1999: Information technology — 8-bit single-byte coded graphic character sets — Part 5: Latin/ Cyrillic alphabet. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 10 pp. LCCN ???? 59 Sfr (print), 44 Sfr (PDF).

                                        URL http://www.iso.org/iso/
                                        en/CatalogueDetailPage.CatalogueDetail?
                                        CSNUMBER=28249.

ISO:1999:IIIag

[ISO99-35] ISO. ISO/IEC 8859-6:1999:

Information technology — 8-bit single-byte coded graphic character sets — Part 6: Latin/ Arabic alphabet. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 11 pp. LCCN ???? 59 Sfr (print), 44 Sfr (PDF).

URL http://www.iso.org/iso/
en/CatalogueDetailPage.CatalogueDetail?
CSNUMBER=28250.

ISO:1999:IIIah

[ISO99-36] ISO. ISO/IEC 8859-8:1999:

Information technology — 8-bit single-byte coded graphic character sets — Part 8: Latin/ Hebrew alphabet. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 12 pp. LCCN ???? 59 Sfr (print), 44 Sfr (PDF).

URL http://www.iso.org/iso/
en/CatalogueDetailPage.CatalogueDetail?
CSNUMBER=28252.

ISO:1999:IIIai

[ISO99-37] ISO. ISO/IEC 8859-9:1999:

Information technology — 8-bit single-byte coded graphic character sets — Part 9: Latin alphabet No. 5. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 10 pp. LCCN ???? 59 Sfr (print), 44 Sfr (PDF).

URL http://www.iso.org/iso/
en/CatalogueDetailPage.CatalogueDetail?
CSNUMBER=28253.

ISO:1999:IIIn

[ISO99-38] ISO. ISO/IEC 9075-1:1999: Information technology — Database languages — SQL — Part 1: Framework (SQL/Framework). International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 67 pp. LCCN ???? CHF 144. URL http://www.

                                        iso.ch/cate/d26196.html.

ISO:1999:IIIo

[ISO99-39] ISO. ISO/IEC 9075-2:1999: Information technology — Database languages — SQL — Part 2: Foundation (SQL/Foundation). International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 1121 pp. LCCN ???? CHF 402.

                                        URL http://www.iso.ch/cate/
                                        d26197.html.

ISO:1999:IIIv

[ISO99-40] ISO. ISO/IEC 9075-4:1999: Information technology — Database languages — SQL — Part 4: Persistent Stored Modules (SQL/ PSM). International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 152 pp. LCCN ???? CHF

                                        200. URL http://www.iso.ch/
                                        cate/d29864.html.

ISO:1999:IIC

[ISO99-41] ISO. ISO/IEC 9075:1992/Cor 3:1999. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 166 pp. LCCN ???? URL http://

                                        www.iso.ch/cate/d30673.html.

ISO:1999:IIP

[ISO99-42] ISO. ISO/IEC 9899:1999: Programming Languages — C. International Organization for Standardization, Geneva, Switzerland, December 16, 1999. ISBN ???? 538 pp. LCCN ???? US$18 (electronic), US$225 (print).

URL http://anubis.dkuug.dk/
JTC1/SC22/open/n2620/n2620.
pdf;   http://anubis.dkuug.
dk/JTC1/SC22/WG14/www/docs/
n897.pdf;   http://webstore.
ansi.org/ansidocstore/product.
asp?sku=ISO%2FIEC+9899%3A1999;
 http://www.iso.ch/cate/d29237.

html. Available in electronic form for online purchase at http://

webstore.ansi.org/ and http:
//www.cssinfo.com/.

ISO:1999:IIT

[ISO99-43] ISO. ISO/IEC TR 14369:1999:

“Information technology — Programming languages, their environments and system software interfaces — Guidelines for the preparation of Language-Independent Service Specifications (LISS)”. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? ???? pp. LCCN ???? US$105.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+TR+14369%3A1999.

ISO:19xx:IDIb

[ISOxxa] ISO. ISO/DIS 10303-23: Industrial automation systems and integration — Product data representation and exchange — Part 23: Implementation methods: C++ language binding to the standard data access interface. International Organization for Standardization, Geneva, Switzerland, 19xx. ISBN ???? ???? pp. LCCN

                                        ???? URL http://www.iso.ch/
                                        cate/d25098.html.

ISO:19xx:IDM

[ISOxxb] ISO. ISO/DIS 15187: Manipulating industrial robots — Graphical user interfaces for programming and operation of robots (GUI-R). International Organization for Standardization, Geneva, Switzerland, 19xx. ISBN ???? ???? pp.

                                        LCCN ???? URL http://www.
                                        iso.ch/cate/d26697.html.

ISO:19xx:IDS

[ISOxxc] ISO. ISO/DIS 15889: Space data and information transfer systems — Data description language — EAST specification. International Organization for Standardization, Geneva, Switzerland, 19xx. ISBN ???? ???? pp. LCCN ????

                                        URL http://www.iso.ch/cate/
                                        d29442.html.

ISO:19xx:IDNa

[ISOxxd] ISO. ISO/DIS 3592: Numerical control of machines — NC processor output — File structure and language format. International Organization for Standardization, Geneva, Switzerland, 19xx. ISBN ???? ???? pp. LCCN ????

                                        URL http://www.iso.ch/cate/

d2553.html. Available in English only.

ISO:19xx:IDNb

[ISOxxe] ISO. ISO/DIS 4343: Numerical control of machines — NC processor output — Post processor commands. International Organization for Standardization, Geneva, Switzerland, 19xx. ISBN ???? ???? pp. LCCN ???? URL http:/

/www.iso.ch/cate/d2537.html.

ISO:19xx:IDIa

[ISOxxf] ISO. ISO/DTR 10303-12: Industrial automation systems and integration — Product data representation and exchange — Part 12: Description methods: The EXPRESS-I language reference manual. International Organization for Standardization, Geneva, Switzerland, 19xx. ISBN ???? 115 pp. LCCN ???? CHF 188.

URL http://www.iso.ch/cate/
d32107.html.

ISO:19xx:IIDk

[ISOxxg] ISO. ISO/IEC 9075-5:1999: Information technology —- Database languages — SQL — Part 5: Host Language Bindings (SQL/ Bindings). International Organization for Standardization, Geneva, Switzerland, 19xx. ISBN ???? 243 pp. LCCN ???? CHF

236. URL http://www.iso.ch/
cate/d26198.html.

ISO:19xx:IIDa

[ISOxxh] ISO. ISO/IEC DIS 10743: Information technology — Standard Music Description Language (SMDL). International Organization for Standardization, Geneva, Switzerland, 19xx. ISBN ???? ???? pp. LCCN ????

                                        URL http://www.iso.ch/cate/
                                        d18833.html.

ISO:19xx:IIDb

[ISOxxi] ISO. ISO/IEC DIS 10858: Information technology — Radixindependent floating-point arithmetic. International Organization for Standardization, Geneva, Switzerland, 19xx. ISBN ????

????  pp.  LCCN ????

                                        URL http://www.iso.ch/cate/
                                        d18890.html.

ISO:19xx:IIDe

[ISOxxj] ISO. ISO/IEC DIS 11756: Information technology — Programming languages — M. International Organization for Standardization, Geneva, Switzerland, 19xx. ISBN ???? ???? pp. LCCN

                                        ???? URL http://www.iso.ch/
                                        cate/d29268.html.

ISO:19xx:IIDg

[ISOxxk] ISO. ISO/IEC DIS 13751: Information technology — Programming languages, their environments and system software interfaces — Programming language APL, extended. International Organization for Standardization, Geneva, Switzerland, 19xx. ISBN ???? ???? pp. LCCN ????

                                        URL http://www.iso.ch/cate/
                                        d22817.html.

ISO:19xx:IIDh

[ISOxxl] ISO. ISO/IEC DIS 14515-1: Information technology — Programming languages, their environments and system software interfaces — Portable Operating System Interface (POSIX) — Test methods for measuring compliance to POSIX — Part 1: System interfaces. International Organization for Standardization, Geneva, Switzerland, 19xx. ISBN ???? ???? pp. LCCN ????

URL http://www.iso.ch/cate/
d24575.html.

ISO:19xx:IIDi

[ISOxxm] ISO. ISO/IEC DIS 15445: Information technology — Document description and processing languages — Hyper Text Markup Language (HTML). International Organization for Standardization, Geneva, Switzerland, 19xx. ISBN ???? ???? pp. LCCN ????

URL http://www.iso.ch/cate/
d27688.html.

ISO:19xx:IIDj

[ISOxxn] ISO. ISO/IEC DIS 9075-10: Information technology — Database languages — SQL — Part 10: Object Language Bindings (SQL/OLB). International Organization for Standardization, Geneva, Switzerland, 19xx. ISBN ???? ???? pp. LCCN ????

URL http://www.iso.ch/cate/
d30613.html.

ISO:2000:IIA

[ISO00a] ISO. ISO 13281-2:2000: Industrial automation systems and integration — Manufacturing Automation Programming Environment (MAPLE) — Part 2: Services and interfaces. International Organization for Standardization, Geneva, Switzerland, 2000. ISBN ???? 46 pp. LCCN ???? CHF

                                        46. URL http://www.iso.ch/
                                        cate/d27722.html.

ISO:2000:IDE

[ISO00b] ISO. ISO 8601:2000. Data elements and interchange formats — Information interchange — Representation of dates and times. International Organization for Standardization, Geneva, Switzerland, 2000. ISBN ???? 29 pp. LCCN ???? CHF 108. URL http://www.

                                        iso.ch/cate/d26780.html.

ISO:2000:IIIf

[ISO00c] ISO. ISO/IEC 13211-2:2000: Information technology — Programming languages — Prolog — Part 2: Modules. International Organization for Standardization, Geneva, Switzerland, 2000. ISBN ???? ???? pp. LCCN ???? US$62.00.

                                        URL http://webstore.ansi.
                                        org/ansidocstore/product.asp?
                                        sku=ISO%2FIEC+13211%2D2%3A2000.

ISO:2000:IIIb

[ISO00d] ISO. ISO/IEC 13673:2000: Information technology — Document processing and related communication — Conformance testing for Standard Generalized Markup Language (SGML) systems. International Organization for Standardization, Geneva, Switzerland, 2000. ISBN ???? 33 pp. LCCN ???? CHF 110. URL http://www.

                                        iso.ch/cate/d22553.html.

ISO:2000:IIId

[ISO00e] ISO. ISO/IEC 14496-5:2000: Information technology — Coding of audio-visual objects — Part 5: Reference software. International Organization for Standardization, Geneva, Switzerland, 2000. ISBN ???? 7 pp. LCCN ???? CHF

56. URL http://www.iso.ch/
cate/d30554.html. Available in

English only.

ISO:2000:IIIe

[ISO00f] ISO. ISO/IEC 1539-2:2000: Information technology — Programming languages — Fortran — Part 2: Varying length character strings. International Organization for Standardization, Geneva, Switzerland, 2000. ISBN ???? ???? pp. LCCN ???? US$58.00.

URL http://webstore.ansi.
org/ansidocstore/product.asp?
sku=ISO%2FIEC+1539%2D2%3A2000.

ISO:2000:IIIa

[ISO00g] ISO. ISO/IEC 2382-7:2000: Information technology — Vocabulary — Part 7: Computer programming. International Organization for Standardization, Geneva, Switzerland, 2000. ISBN ???? 69 pp. LCCN ???? CHF

164. URL http://www.iso.ch/
cate/d7241.html.

ISO:2000:IIIc

[ISO00h] ISO. ISO/IEC 9579:2000: Information technology — Remote database access for SQL with security enhancement. International Organization for Standardization, Geneva, Switzerland, 2000. ISBN ???? 115 pp. LCCN ???? CHF

188. URL http://www.iso.ch/
cate/d30615.html.

ISO:2000:IIT

[ISO00i] ISO. ISO/IEC TR 15942:2000: Information technology — Programming languages — Guide for the use of the Ada programming language in high integrity systems. International Organization for Standardization, Geneva, Switzerland, 2000. ISBN ???? 48 pp. LCCN ???? CHF 136; US$92.00.

                                        URL http://webstore.ansi.
                                        org/ansidocstore/product.asp?
                                        sku=ISO%2FIEC+TR+15942%3A2000;
                                        http://www.iso.ch/cate/d29575.

html. Available in English only.

ISO:2001:IIIc

[ISO01a] ISO. ISO/IEC 10967-2: Information technology — Language independent arithmetic — Part 2: Elementary numerical functions. International Organization for Standardization, Geneva, Switzerland, August 15, 2001. ISBN ???? x + 177 pp. LCCN ????

                                        URL http://standards.iso.
                                        org/ittf/PubliclyAvailableStandards/
                                        c024427_ISO_IEC_10967-2_2001(
                                        E).zip; http://www.iso.ch/cate/
                                        d24427.html.

ISO:2001:IICa

[ISO01b] ISO. ISO/IEC 8652:1995/Cor 1:2001, Technical Corrigendum to Information Technology — Programming Languages — Ada. International Organization for Standardization, Geneva, Switzerland, 2001. ISBN ???? ???? pp. LCCN ????

ISO:2001:IIIa

[ISO01c] ISO. ISO/IEC 8859-11:1998:

Information technology — 8-bit single-byte coded graphic character sets — Part 11: Latin/ Thai alphabet. International Organization for Standardization, Geneva, Switzerland, 2001. ISBN ???? 10 pp. LCCN ???? 59 Sfr (print), 44 Sfr (PDF).

URL http://www.iso.org/iso/
en/CatalogueDetailPage.CatalogueDetail?
CSNUMBER=28263.

ISO:2001:IIIb

[ISO01d] ISO. ISO/IEC 8859-16:1998:

Information technology — 8-bit single-byte coded graphic character sets — Part 16: Latin alphabet No. 10. International Organization for Standardization, Geneva, Switzerland, 2001. ISBN ???? 9 pp. LCCN ???? 59 Sfr (print), 4 Sfr (PDF).

URL http://www.iso.org/iso/
en/CatalogueDetailPage.CatalogueDetail?
CSNUMBER=28263.

ISO:2001:IICb

[ISO01e] ISO. ISO/IEC 9899:1999/ Cor 1:2001. International Organization for Standardization, Geneva, Switzerland, 2001. ISBN

????  6 pp.  LCCN

???? URL http://www.iso.org/
iso/en/CatalogueDetailPage.
CatalogueDetail?CSNUMBER=35952.

ISO:2002:III

[ISO02] ISO. ISO/IEC 1989:2002: Information technology — Programming languages — COBOL. International Organization for Standardization, Geneva, Switzerland, 2002. ISBN ???? 859

pp.  LCCN ????  352

SFr (print), 44 SFr (PDF).

                                        URL http://www.iso.org/iso/
                                        en/CatalogueDetailPage.CatalogueDetail?
                                        CSNUMBER=28805.

ISO:2003:IIP

[ISO03a] ISO. ISO/IEC 14882:2003: Programming languages — C++. International Organization for Standardization, Geneva, Switzerland, 2003. ISBN ???? 757 pp. LCCN ???? 352 Sfr (print or electronic).

                                        URL http://www.iso.org/iso/
                                        en/CatalogueDetailPage.CatalogueDetail?
                                        CSNUMBER=38110.

ISO:2003:IIIb

[ISO03b] ISO. ISO/IEC 23270:2003: Information technology — C# Language Specification. International Organization for Standardization, Geneva, Switzerland, 2003. ISBN ???? xiii + 471 pp. LCCN ???? 328 Sfr (print).

                                        URL http://standards.iso.
                                        org/ittf/PubliclyAvailableStandards/
                                        c036768_ISO_IEC_23270_2003(
                                        E).zip;   http://www.iso.org/
                                        iso/en/CatalogueDetailPage.
                                        CatalogueDetail?CSNUMBER=36768.

ISO:2003:IIIa

[ISO03c] ISO. ISO/IEC 8859-7:2003: Information technology — 8-bit single-byte coded graphic character sets — Part 7: Latin/ Greek alphabet. International Organization for Standardization, Geneva, Switzerland, 2003. ISBN ???? 9 pp. LCCN ???? 59 Sfr (print), 44 Sfr (PDF).

                                        URL http://www.iso.org/iso/
                                        en/CatalogueDetailPage.CatalogueDetail?
                                        CSNUMBER=38580.

ISO:2004:DIS

[ISO04a] ISO. Draft International Standard ISO/IEC 1539-1:2004(E): Information technology — Programming languages — Fortran Part 1: Base Language. International Organization for Standardization, Geneva, Switzerland, May 2004. xiv + 569 pp. URL ftp://

ftp.nag.co.uk/sc22wg5/N1601-
N1650/N1601.pdf.gz.

ISO:2004:IIIb

[ISO04b] ISO. ISO/IEC 10118-3:2004:

Information technology — Security techniques — Hash-functions — Part 3: Dedicated hashfunctions. International Organization for Standardization, Geneva, Switzerland, February 2004. 94

pp. URL http://www.iso.org/
iso/en/CatalogueDetailPage.
CatalogueDetail?CSNUMBER=39876.

ISO:2004:IIIa

[ISO04c] ISO. ISO/IEC 1539-1:2004 Information technology — Programming languages — Fortran – Part 1: Base language. International Organization for Standardization, Geneva, Switzerland, 2004. xiv +

569 pp. URL http://www.dkuug.
dk/jtc1/sc22/open/n3661.pdf.

ISO:2005:IDM

[ISO05a] ISO. ISO 19005-1:2005, Document management—Electronic document file format for longterm preservation—Part 1: Use of PDF 1.4 (PDF/A-1). International Organization for Standardization, Geneva, Switzerland,

2005. URL http://www.aiim.
org/documents/standards/ISO_
19005-1_(E).doc;   http://www.
aiim.org/pdf_a/.

ISO:2005:IIIa

[ISO05b] ISO. ISO/IEC 23271:2003: Information technology — Common Language Infrastructure. International Organization for Standardization, Geneva, Switzerland, 2005. ISBN ???? xi + 99 (Part. I), ix + 164 (Part. II), vi + 125 (Part. III), iii + 16 (Part. IV), iv + 79 (Part. V) pp. LCCN ???? 328 Sfr (print).

URL http://standards.iso.
org/ittf/PubliclyAvailableStandards/
c036769_ISO_IEC_23271_2003(
E).zip;   http://www.iso.org/
iso/en/CatalogueDetailPage.
CatalogueDetail?CSNUMBER=36769.

ISO:2005:IID

[ISO05c] ISO. ISO/IEC DIS 23360: Draft International Standard: Linux Standard Base Core Specification 2.0.1. International Organization for Standardization, Geneva, Switzerland, 2005. ISBN

????  1214 pp.  LCCN

???? URL http://www.iso.org/
iso/en/CatalogueDetailPage.
CatalogueDetail?CSNUMBER=41481&
scopelist=PROGRAMME; http:/
/www.linuxbase.org/LSBWiki/
IsoBallot.

ISO:2005:IIIb

[ISO05d] ISO. ISO/IEC TR 23272: Information technology — Common Language Infrastructure — Profiles and Libraries. International Organization for Standardization, Geneva, Switzerland, 2005. ISBN ???? 6 pp. LCCN ???? 328 Sfr (print).

URL http://standards.iso.
org/ittf/PubliclyAvailableStandards/
c036770_ISO_IEC_TR_23272_2003(
E).zip;   http://www.iso.org/
iso/en/CatalogueDetailPage.
CatalogueDetail?CSNUMBER=36770.

ISO:2006:IIId

[ISO06a] ISO. ISO/IEC 10967-3: Information technology — Language independent arithmetic — Part 3: Complex integer and floating point arithmetic and complex elementary numerical functions. International Organization for Standardization, Geneva, Switzerland, 2006. viii + 150 pp.

URL https://standards.iso.
org/ittf/PubliclyAvailableStandards/
c037994_ISO_IEC_10967-3_2006(
E).zip.

ISO:2006:IIIa

[ISO06b] ISO. ISO/IEC 10967-2: Information technology — Language independent arithmetic — Part 3: Complex integer and floating point arithmetic and complex elementary numerical functions. International Organization for Standardization, Geneva, Switzerland, August 15, 2006. ISBN ???? viii + 149 pp. LCCN ????

URL http://standards.iso.
org/ittf/PubliclyAvailableStandards/
c024427_ISO_IEC_10967-2_2001(
E).zip; http://www.iso.ch/cate/
d24427.html.

ISO:2006:IIIb

[ISO06c] ISO. ISO/IEC 23270:2006: Information technology — Programming languages — C#. Technical report. International Organization for Standardization, Geneva, Switzerland, 2006. LCCN ????

                                        URL http://www.iso.org/iso/
                                        iso_catalogue/catalogue_tc/
                                        catalogue_detail.htm?csnumber=
                                        42926.

ISO:2006:IIIc

[ISO06d] ISO. ISO/IEC 23271:2006: Information technology: Common Language Infrastructure (CLI) Partitions I to VI. International standard. International Organization for Standardization, Geneva, Switzerland, second edition, 2006. LCCN ????

ISO:2006:IDO

[ISO06e] ISO. ISO/IEC 26300: Open Document Format for Office Applications (OpenDocument) v1.0. International Organization for Standardization, Geneva, Switzerland, 2006. ISBN ???? 706 pp. LCCN

                                        ???? URL http://www.iso.org/
                                        iso/en/CatalogueDetailPage.
                                        CatalogueDetail?CSNUMBER=43485.

ISO:2006:ITCa

[ISO06f] ISO. ISO/IEC TR 23272:2006: Information technology: Common Language Infrastructure (CLI): technical report on information derived from Partition IV XML file. Technical report. International Organization for Standardization, Geneva, Switzerland, second edition, 2006. ???? pp. LCCN ????

ISO:2006:ITCb

[ISO06g] ISO. ISO/IEC TR 25438:2006:

Information technology: Common Language Infrastructure (CLI): technical report: common generics. Technical report. International Organization for Standardization, Geneva, Switzerland, 2006. LCCN ????

ISO:2007:ISI

[ISO07a] ISO. International Standard ISO/ IEC 9899:1999: Technical Corrigendum 3. International Organization for Standardization, Geneva, Switzerland, November 15, 2007. ISBN ???? 10 pp. LCCN

???? URL http://www.iso.org/
iso/en/CatalogueDetailPage.
CatalogueDetail?CSNUMBER=43485;
http://www.open-std.org/jtc1/
sc22/wg14/www/docs/n1256.pdf.

ISO:2007:IID

[ISO07b] ISO. ISO 3297: Information and documentation — International standard serial number (ISSN): Information et documentation — Numéro international normalisé des publications en série (ISSN). International Organization for Standardization, Geneva, Switzerland, fourth edition, September 1, 2007. ISBN 0-580-54132-0. v + 20 pp. LCCN

???? URL https://www.iso.
org/standard/39601.html.

ISO:2007:IIA

[ISO07c] ISO. ISO/IEC 8652:1995/Amd 1:2007, Amendment to Information Technology — Programming Languages — Ada. International Organization for Standardization, Geneva, Switzerland, 2007. ???? pp. LCCN ???? URL http:/

/www.iso.org/iso/home/store/
catalogue_ics/catalogue_detail_
ics.htm?csnumber=45001.

ISO:2007:IIT

[ISO07d] ISO. ISO/IEC TR 24731-1:2007 Information technology — Programming languages, their environments and system software interfaces — Extensions to the C library — Part 1: Bounds-checking interfaces. International Organization for Standardization, Geneva, Switzerland, November 15, 2007. ISBN ???? 81 pp. LCCN ????

URL http://www.iso.org/iso/
iso_catalogue/catalogue_tc/
catalogue_detail.htm?csnumber=
38841.

ISO:2008:IDM

[ISO08a] ISO. ISO 32000-1:2008. Document management — Portable document format — Part 1: PDF 1.7. International Organization for Standardization, Geneva, Switzerland, 2008. ISBN ???? 747 pp. LCCN ???? 238 Sfr (PDF or CD).

URL http://www.iso.org/iso/
iso_catalogue/catalogue_tc/
catalogue_detail.htm?csnumber=
51502.

ISO:2008:IIIc

[ISO08b] ISO. ISO/IEC 9075-10:2008 Information technology — Database languages — SQL — Part 10: Object Language Bindings (SQL/OLB). International Organization for Standardization, Geneva, Switzerland, July 17, 2008. ISBN ???? 376 pp. LCCN ????

URL http://www.iso.org/iso/
iso_catalogue/catalogue_tc/
catalogue_detail.htm?csnumber=
38644.

ISO:2008:IIId

[ISO08c] ISO. ISO/IEC 9075-13:2008 Information technology — Database languages — SQL — Part 13: QL Routines and Types Using the Java TM Programming Language (SQL/JRT). International Organization for Standardization, Geneva, Switzerland, July 17, 2008. ISBN ???? 186 pp. LCCN ????

URL http://www.iso.org/iso/
iso_catalogue/catalogue_tc/
catalogue_detail.htm?csnumber=
38646.

ISO:2008:IIIa

[ISO08d] ISO. ISO/IEC 9075-3:2008 Information technology — Database languages — SQL — Part 3: Call-Level Interface (SQL/CLI). International Organization for Standardization, Geneva, Switzerland, July 17, 2008. ISBN ???? 378 pp. LCCN ????

URL http://www.iso.org/iso/
iso_catalogue/catalogue_tc/
catalogue_detail.htm?csnumber=
38641.

ISO:2008:IIIb

[ISO08e] ISO. ISO/IEC 9075-9:2008 Information technology — Database languages — SQL — Part 9: Management of External Data (SQL/MED). International Organization for Standardization, Geneva, Switzerland, July 17, 2008. ISBN ???? 451 pp. LCCN ????

                                        URL http://www.iso.org/iso/
                                        iso_catalogue/catalogue_tc/
                                        catalogue_detail.htm?csnumber=
                                        38643.

ISO:2009:IID

[ISO09a] ISO. ISO 28500:2009: Information and documentation — WARC file format. International Organization for Standardization, Geneva, Switzerland, May 2009.

                                        ?? pp. URL https://www.iso.
                                        org/standard/68004.html.

ISO:2009:IQU

[ISO09b] ISO. ISO 80000-2: Quantities and units — Part 2: Mathematical signs and symbols to be used in the natural sciences and technology. International Organization for Standardization, Geneva, Switzerland, December 2009. 40

                                        pp. URL https://www.iso.org/
                                        standard/31887.html.

ISO:2009:III

[ISO09c] ISO. ISO/IEC 24747:2009 Information technology — Programming languages, their environments and system software interfaces — Extensions to the C Library to support mathematical special functions. Technical report. International Organization for Standardization, Geneva, Switzerland, 2009. 25 pp. LCCN ????

                                        URL http://www.iso.org/iso/
                                        iso_catalogue/catalogue_tc/

catalogue_detail.htm?csnumber=
38857.

ISO:2009:IIT

[ISO09d] ISO. ISO/IEC TR 24732:2009 Information technology — Programming languages, their environments and system software interfaces — Extension for the programming language C to support decimal floatingpoint arithmetic. Technical report. International Organization for Standardization, Geneva, Switzerland, 2009. LCCN ????

URL http://www.iso.org/iso/
iso_catalogue/catalogue_tc/
catalogue_detail.htm?csnumber=
38842.

ISO:2010:IIIb

[ISO10a] ISO. ISO/IEC 1539-1:2010 Information technology — Programming languages — Fortran — Part 1: Base language. International Organization for Standardization, Geneva, Switzerland, June 2010. xviii + 603 pp. URL ftp://

ftp.nag.co.uk/sc22wg5/N1801-
N1850/N1830.pdf.

ISO:2010:IIIa

[ISO10b] ISO. ISO/IEC 29124:2010: Information technology — Programming languages, their environments and system software interfaces — Extensions to the C++ Library to support mathematical special functions. Technical report. International Organization for Standardization, Geneva, Switzerland, 2010. LCCN ????

URL http://www.iso.org/iso/
iso_catalogue/catalogue_tc/

                                        catalogue_detail.htm?csnumber=
                                        50511.

ISO:2011:IIIh

[ISO11a] ISO. ISO/IEC 14882:2011 Information technology — Programming languages — C++. International Organization for Standardization, Geneva, Switzerland, third edition, September 1, 2011. ISBN ???? ???? pp. LCCN ????

                                        URL http://www.iso.org/iso/
                                        iso_catalogue/catalogue_tc/
                                        catalogue_detail.htm?csnumber=
                                        50372.

ISO:2011:IIIc

[ISO11b] ISO. ISO/IEC 9075-1:2011 Information technology — Database languages — SQL — Part 1: Framework (SQL/Framework). International Organization for Standardization, Geneva, Switzerland, December 15, 2011. ISBN ???? ???? pp. LCCN ????

                                        URL http://www.iso.org/iso/
                                        iso_catalogue/catalogue_tc/
                                        catalogue_detail.htm?csnumber=
                                        53681.

ISO:2011:IIIg

[ISO11c] ISO. ISO/IEC 9075-14:2011 Information technology — Database languages — SQL — Part 14: XML-Related Specifications (SQL/XML). International Organization for Standardization, Geneva, Switzerland, December 15, 2011. ISBN ???? 422 pp. LCCN ????

                                        URL http://www.iso.org/iso/
                                        iso_catalogue/catalogue_tc/
                                        catalogue_detail.htm?csnumber=
                                        53686.

ISO:2011:IIIf

[ISO11d] ISO. ISO/IEC 9075-2:2011 Information technology — Database languages — SQL — Part 11: Information and Definition Schemas (SQL/Schemata). International Organization for Standardization, Geneva, Switzerland, December 15, 2011. ISBN ???? 282 pp. LCCN ????

URL http://www.iso.org/iso/
iso_catalogue/catalogue_tc/
catalogue_detail.htm?csnumber=
53685.

ISO:2011:IIId

[ISO11e] ISO. ISO/IEC 9075-2:2011 Information technology — Database languages — SQL — Part 2: Foundation (SQL/Foundation). International Organization for Standardization, Geneva, Switzerland, December 15, 2011. ISBN ???? 1434 pp. LCCN ????

ISO:2011:IIIe

[ISO11f] ISO. ISO/IEC 9075-4:2011 Information technology — Database languages — SQL — Part 4: Persistent Stored Modules (SQL/PSM). International Organization for Standardization, Geneva, Switzerland, December 15, 2011. ISBN ???? 164

pp.  LCCN ????  URL

http://www.iso.org/iso/iso_
catalogue/catalogue_tc/catalogue_
detail.htm?csnumber=53684.

ISO:2011:IIIb

[ISO11g] ISO. ISO/IEC 9899:2011 Information technology — Programming languages — C. International Organization for Standardization, Geneva, Switzerland, December 8, 2011. ISBN ???? 683 (est.) pp. LCCN ????

URL http://www.iso.org/iso/
iso_catalogue/catalogue_tc/
catalogue_detail.htm?csnumber=
57853.

ISO:2011:IIIa

[ISO11h] ISO. ISO/IEC/IEEE 60559:2011 Information technology — Microprocessor Systems — Floating-Point arithmetic. International Organization for Standardization, Geneva, Switzerland, 2011. 58 pp.

URL http://www.iso.org/iso/
iso_catalogue/catalogue_tc/
catalogue_detail.htm?csnumber=
57469.

ISO:2012:IID

[ISO12a] ISO. ISO 26324:2012: Information and documentation — Digital object identifier system. International Organization for Standardization, Geneva, Switzerland, 2012. 17 pp. URL http:/

/www.iso.org/iso/home/store/
catalogue_tc/catalogue_detail.
htm?csnumber=43506.

ISO:2012:IIIb

[ISO12b] ISO. ISO/IEC 8652:2012 Information technology — Programming languages — Ada. International Organization for Standardization, Geneva, Switzerland, 2012. ISBN ???? 832 (est.) pp. LCCN ???? URL http:/

/www.ada-auth.org/standards/
ada12.html;   http://www.iso.
org/iso/home/store/catalogue_

ics/catalogue_detail_ics.htm?
csnumber=61507.

ISO:2014:IIIb

[ISO14a] ISO. ISO/IEC 14882:2014 Information technology — Programming languages — C++. International Organization for Standardization, Geneva, Switzerland, fourth edition, December 15, 2014. ISBN ???? ???? pp. LCCN ???? URL http:/

/www.iso.org/iso/home/store/
catalogue_ics/catalogue_detail_
ics.htm?csnumber=64029.

ISO:2014:IIIa

[ISO14b] ISO. ISO/IEC 1989:2014: Information technology — Programming languages — COBOL. International Organization for Standardization, Geneva, Switzerland, second edition, 2014. ISBN ???? 927 pp. LCCN ???? CHF 198 (PDF ISO); $232 (PDF, ANSI). URL https://

webstore.ansi.org/Standards/
ISO/ISOIEC19892014;   https:
//www.iso.org/standard/51416.
html.

ISO:2017:IID

[ISO17a] ISO. ISO 28500:2017: Information and documentation — WARC file format. International Organization for Standardization, Geneva, Switzerland, second edition, August 2017. 26

pp. URL https://www.iso.org/
standard/68004.html.

ISO:2017:IIIb

[ISO17b] ISO. ISO/IEC 13816:2007: Information technology — Programming languages, their environments and system software interfaces — Programming language ISLISP. International Organization for Standardization, Geneva, Switzerland, November 2017. 127 pp. URL http:

//islisp.info/Documents/PDF/
islisp-2007-03-17-pd-v23.pdf;
 https://www.iso.org/standard/
44338.html.

ISO:2017:IIIa

[ISO17c] ISO. ISO/IEC 14882:2017 Information technology — Programming languages — C++. International Organization for Standardization, Geneva, Switzerland, fifth edition, December 2017. ISBN

????  1605 pp.  LCCN

???? URL https://www.iso.
org/standard/68564.html.

ISO:2018:III

[ISO18] ISO. ISO/IEC 9899:2018 Information technology — Programming languages — C. International Organization for Standardization, Geneva, Switzerland, fourth edition, June 2018. ISBN ???? 520 pp. LCCN

???? URL https://www.iso.
org/standard/68564.html.

ISO:2019:IQU

[ISO19] ISO. ISO 80000-2:2019: Quantities and units. Part 2: Mathematics. International Organization for Standardization, Geneva, Switzerland, 2019. 36 pp. URL https:

//en.wikipedia.org/wiki/ISO/
IEC_80000;   https://www.iso.
org/standard/64973.html.

ISO:2020:IDM

[ISO20a] ISO. ISO 32000-2:2010. Document management — Portable document format — Part 2: PDF 3.0. International Organization for Standardization, Geneva, Switzerland, 2020. ISBN ???? 986 pp. LCCN ???? 198 Sfr (PDF

or CD). URL https://www.iso.
org/standard/75839.html.

ISO:2020:IID

[ISO20b] ISO. ISO 3297:2020: Information and documentation — International standard serial number (ISSN). International Organization for Standardization, Geneva, Switzerland, sixth edition, 2020.

26 pp. URL https://www.iso.
org/standard/73846.html.

ISO:2020:III

[ISO20c] ISO. ISO/IEC 60559:2020 Information technology — Microprocessor Systems — Floating-Point arithmetic. International Organization for Standardization, Geneva, Switzerland, 2020. 74

pp. URL https://www.iso.org/
standard/80985.html.

ISO:2021:IIC

[ISO21] ISO. ISO/IEC CD 1989.2: Information technology — Programming languages, their environments and system software interfaces — Programming language COBOL. International Organization for Standardization, Geneva, Switzerland, third edition, 2021. ISBN ???? ???? pp. LCCN

???? URL https://www.iso.
org/standard/74527.html.

ISO:2023:III

[ISO23] ISO. ISO/IEC 1989:2023: Information technology — Programming languages, their environments and system software interfaces — Programming language COBOL. International Organization for Standardization, Geneva, Switzerland, third edition, January 2023. ISBN ???? 1229 pp.

                                        LCCN ???? URL https://www.
                                        iso.org/standard/74527.html.

ISO:2024:III

[ISO24] ISO. ISO/IEC 9899:2024 Information technology — Programming languages — C. International Organization for Standardization, Geneva, Switzerland, fifth edition, October 2024. ISBN ???? 758 pp. LCCN ???? URL https:/

                                        /www.iso.org/standard/82075.
                                        html.

IEEE:1993:DTD

[ITC93] Designing, testing, and diagnostics — join them: International Test Conference 1993 proceedings: October 17–21, 1993, Convention Center, Baltimore, Maryland, USA, Proceedings of the International Test Conference. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-1429-8. LCCN TK7874.I474 1993. IEEE catalog number 93CH3356-3.

Kanazy:1993:SDP

[Kan93] Y. Kanazy. SGML document processing. Fujitsu, 44(6):521–526, ???? 1993. CODEN FUJTAR. ISSN 0016-2515.

Karlsson:2005:IIL

[Kar05] Kent Karlsson. ISO/IEC 10967, Language Independent Arithmetic (LIA). Wikipedia article, October 31, 2005. URL http://en.

wikipedia.org/wiki/ISO/IEC_
10967;   http://www.open-std.
org/JTC1/SC22/WG11/docs/n364.
pdf;   http://www.open-std.
org/JTC1/SC22/WG11/docs/n462.
pdf;   http://www.open-std.
org/jtc1/sc22/wg11/docs/n490.
pdf.

Kim:1992:FDS

[Kim92] Kyongsok Kim. A future direction in standardizing international character codes — with a special reference to ISO/IEC 10646 and Unicode. Computer Standards and Interfaces, 14(3):209– 221, May 1992. CODEN CSTIEZ. ISSN 0920-5489.

Komachi:1991:ISO

[Kom91] Y. Komachi. International standardization overview on document description languages and fonts. Journal of the Information Processing Society of Japan = Joho Shori, 32(10):1110–1117, ???? 1991. CODEN JOSHA4. ISSN 0447-8053.

Komachi:1993:ISF

[Kom93] Y. Komachi. International standardization for front (??font??) information interchange. Journal of the Information Processing Society of Japan = Joho Shori, 34(7):915– 921, July 1993. CODEN JOSHA4. ISSN 0447-8053.

Komachi:1994:ISM

[Kom94] Y. Komachi. International standardization for multimedia/ hypermedia information interchange. Journal of the Information Processing Society of Japan = Joho Shori, 35(7):632–641, July 1994. CODEN JOSHA4. ISSN 0447-8053.

Leathrum:1993:ATO

[LL93a] J. F. Leathrum and K. A. Liburdy. Automated testing of open software standards. In ITC’93 [ITC93], pages 854–858. ISBN 0-7803-1429-8. LCCN TK7874.I474 1993. IEEE catalog number 93CH3356-3.

Leathrum:1993:ERT

[LL93b] J. F. Leathrum and K. A. Liburdy. Evolving role of testing in open systems standards. In ITC’93 [ITC93], pages 273–274. ISBN 0-7803-1429-8. IEEE catalog number 93CH3356-3.

Meek:1995:SGR

[Mee95] Brian L. Meek. Seven golden rules for producing languageindependent standards. Proceedings of the IEEE International Software Engineering Standards Symposium, pages 250–256, 1995. ISSN 1082-3670.

Molesworth:1991:OIE

[Mol91] R. Molesworth. The open interchange of electronic documents (open document architecture (ODA)). In Anonymous [Ano91], pages 4/1–6. ISBN ???? LCCN ????

Melby:1992:TSI

[MW92] A. K. Melby and S. E. Wright.

TEI-TERM: an SGML-based interchange format for terminology files. In Anonymous [Ano92], pages 19–37. ISBN 0-85142-295-0. LCCN ????

Newcomb:1991:HHD

[NKN91] Steven R. Newcomb, Neill A.

Kipp, and Victoria T. Newcomb. The “HyTime”: hypermedia/ time-based document structuring language. Communications of the ACM, 34(11):67–83, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.
org/pubs/toc/Abstracts/0001-
0782/125495.html.

Nicholas:1992:ISO

[NW92] C. K. Nicholas and L. A. Welsch.

On the interchangeability of SGML and ODA. Electronic Publishing—Origination, Dissemination, and Design, 5(3):105– 130, September 1992. CODEN EPODEU. ISSN 0894-3982.

Paulk:1995:HIC

[Pau95] Mark C. Paulk. How ISO 9001 compares with the CMM. IEEE Software, 12(1):74–83, January 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Price:1993:MIF

[Pri93] Roger Price. MHEG: An introduction to the future international standard for hypermedia object interchange. In Proceedings of the 1st ACM International Conference on Multimedia Aug 1–6, 1993, Anaheim, CA, USA, pages 121–128. ACM Press, New York, NY 10036, USA, 1993. US$.

Reizer:1994:UFM

[RAMP94] Neal R. Reizer, Gregory D. Abowd, B. Craig Meyers, and Patrick R. H. Place. Using formal methods for requirements specification of a proposed POSIX standard. In IEEE [IEE94], pages 118– 125. ISBN 0-8186-5480-5, 0-8186-5481-3. LCCN QA76.758.I5717 1994.

Reid:1993:LIA

[Rei93] John Reid. Language Independent Arithmetic (LIA) — a Draft International Standard (DIS). ACM SIGNUM Newsletter, 28(1):2–7, January 1993. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Role:1991:SSD

[Rol91] F. Role. The SGML standard: describing the logical structure of documents. Documentaliste – Sciences de l’Information, 28(4-5): 187–192, July–October 1991. CO-DEN DSINE6. ISSN 0012-4508.

Smura:1989:FII

[SBSG89] Edwin J. Smura, Barbara Beeton, Karla Savage, and Alan Griffee. Font information interchange standard ISO/IEC 9541. Computer Communications, 12 (2):93–96, April 1989. CODEN COCOD7. ISSN 0140-3664.

Scheller:1988:DSA

[Sch88] A. Scheller. Document standards:

availability and products. Computer Networks and ISDN Systems, 16(1-2):138–142, September 1988. CODEN CNISE9. ISSN 0169-7552.

Schildt:1990:AAC

[Sch90] Herbert Schildt. The Annotated ANSI C Standard: American National Standard for Programming Languages: C. Osborne/Mc-Graw-Hill, Berkeley, CA, USA, 1990. ISBN 0-07-881952-0. xvi + 219 pp. LCCN QA76.73.C15S356 1990. US$39.95.

Severance:1995:SVF

[Sev95] C. Severance. Standards: The value of the formal standards process. Computer, 28(8):82–83, August 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:1986:ISP

[Smi86] Joan M. Smith. Implications of SGML for the preparation of scientific publications. The Computer Journal, 29(3):193–200, June 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Smith:1987:SGM

[Smi87] J. M. Smith. The standard generalized markup language (SGML): an application for OSI. In Gerner and Spaniol [GS87], pages 254– 259. ISBN 3-540-17213-0. LCCN TK5105.5 .G561 1987.

Smura:1987:TNB

[SP87] Edwin J. Smura and Archie D. Provan. Toward a new beginning: the development of a standard for font and character encoding to control electronic document interchange. IEEE Transactions on Professional Communication, PC-30(4):259–264, December 1987. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Tristram:1995:DYR

[Tri95] Claire Tristram. Do You Really Need ISO 9000? Open Computing, 12(5):25, May 1995. CODEN OPCOEB. ISSN 1078-2370.

Wilkes:1994:IPS

[Wil94] T. Wilkes, editor. IPCC 94 Proceedings. Scaling New Heights in Technical Communication. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-1936-2. LCCN T10.5.I54 1994. IEEE Catalog No. 94CH3442-1.